### v1

# Assortment Optimization for Patient-Provider Matching

Abstract
Rising provider turnover forces healthcare administrators to frequently rematch patients to available providers, which can be cumbersome and labor-intensive. To reduce the burden of rematching, we study algorithms for matching patients and providers through assortment optimization. We develop a patient-provider matching model in which we simultaneously offer each patient a menu of providers, and patients subsequently respond and select providers. By offering assortments upfront, administrators can balance logistical ease and patient autonomy. We study policies for assortment optimization and characterize their performance under different problem settings. We demonstrate that the selection of assortment policy is highly dependent on problem specifics and, in particular, on a patient’s willingness to match and the ratio between patients and providers . On real-world data, we show that our best policy can improve match quality by 13% over a greedy solution by tailoring assortment sizes based on patient characteristics. We conclude with recommendations for running a real-world patient-provider matching system inspired by our results.
1 Introduction
Primary care providers are essential to the healthcare ecosystem because they are the first point of contact for many patients (Pearson and Raeke, 2000; Wu et al., 2022). Patients rely on primary care providers for routine checkups and referrals to specialists. Moreover, care continuity can instill trust and improve medication uptake rates and patient health (Wu et al., 2022).
Unfortunately, high provider turnover rates frequently lead to patients without an assigned provider (Reddy et al., 2015). Provider turnover disrupts patient care and leads to worse care (Reddy et al., 2015). In principle, healthcare administrators reassign unmatched patients to other providers; however, in practice, the process takes months due to provider scarcity and the logistical burden of rematching and coordinating patient matches (Hedden et al., 2021). While many patients find their new provider quickly, others have to wait years to find a new provider due to large numbers of patients, high turnover rates, and provider scarcity (Hedden et al., 2021; Shanafelt et al., 2012).
Algorithms that automatically match patients and providers can reduce logistical hassle but require balancing patient autonomy and system-wide utility. For example, while automatically assigning each patient to a provider would decrease wait times, it also reduces patient autonomy because patients cannot select their provider (Entwistle et al., 2010; Gaynor et al., 2016). In contrast, offering patients full autonomy could lead to suboptimal matches; for example, patients who rematch quickly might prevent better matches for late-responding patients. In addition, granting patients full independence can be overwhelming and delay decisions (Bate and Robert, 2005). We propose to match patients and providers through an assortment framework (Shi, 2016; Rios and Torrico, 2023; Davis et al., 2013). Assortments are menus of options given to customers who select an option from such a menu (Bertsimas and Mišic, 2015). Assortments are used in e-commerce settings where platforms give customers a suggested list of products and have them select from this list (Tian et al., 2024). In an assortment setting, each patient receives a menu of providers and selects a provider from such a menu. These curated offer sets allow administrators to balance patient autonomy and system-wide match utility.
We devise policies for offering assortments in a patient-provider matching scenario. To do so, we first develop a model of patient-provider matching then construct a set of policies for offering assortments. We characterize how policy performance depends on problem characteristics such as patient-to-provider ratio, and we empirically detail the regimes under which each of our policies performs best. We then evaluate our policies in a real-world scenario constructed on top of a Medicare dataset to assess the performance of our policies in a real-world example.  [^1] We include all code at https://github.com/naveenr414/patient-provider
Our key question is the following: How should healthcare administrators design provider menus for patients to optimize system-wide match rates and quality?
1.1 Contributions
We develop a model of patient-provider matching using assortment offerings and analyze its impact on match rate and match quality. We characterize well-performing policies when varying properties such as the patient/provider ratio. We then empirically analyze these policies in a synthetic and real-world dataset built using a Medicare dataset. We use the insights from these studies to propose recommendations for designing patient-provider matching systems in practice.
Developing a Model of Patient-Provider Matching
We develop a model of patient-provider matching where administrators offer assortments upfront, then patients sequentially respond in a random order and select providers one by one. We do so because it captures two critical elements: 1) patients have autonomy in selecting providers and 2) administrators have some control over the matching process by varying assortments offered to patients. We model patient decisions through a choice model that captures how the set of providers offered impacts the final provider chosen. The selection of choice model is key to the fidelity of our model, so we analyze various choice models, including the uniform choice model and the multinomial logit choice model. Finally, we capture patient heterogeneity through a match quality matrix, which captures differences in match quality between pairs of patients and providers. Match quality can encompass geographical proximity, provider specialties, and concordance along language, race, and gender (Greenwood et al., 2020, 2018; Manson, 1988). Patients tend to match with higher quality providers, though the exact match probabilities are dictated by their choice model.
Constructing Assortment Policies
We characterize the optimal policy in the single-provider scenario and use these insights to construct a set of policies for assortment optimization. We first construct a baseline policy, greedy, which offers every provider to every patient. We demonstrate that the greedy policy can do arbitrarily poorly, so we build upon this by developing three new policies that adaptively vary assortment sizes: pairwise, group-based, and gradient descent. The pairwise policy offers each patient at most one provider, and we demonstrate an instance-dependent approximation guarantee for this policy. We then develop the group-based policy by augmenting the pairwise policy so we offer assortments to groups of patients. We demonstrate that such group-based maintains the match rate of the pairwise policy while potentially improving the match quality. For the gradient descent policy, we construct an objective function that provably lower bounds the match quality, and we use gradient descent to optimize this quantity.
Characterizing Policy Performance
We analyze the performance of our policies in a synthetic setting and empirically characterize the conditions under which each policy performs best. We demonstrate that the best-performing policy depends on two factors: the match probability and the ratio of patients to providers. When patients outnumber providers, as is common in many real-world health systems, the gradient descent policy performs best because the objective function is closer to the lower bound. When patient and provider counts are balanced, we show that the greedy policy achieves a high match quality when patients have low match probabilities, while the group-based policy achieves a high match quality when patients have high match probabilities.
We evaluate policies in a real-world scenario constructed via Medicare data (Medicare & Medicaid, 2025) and we show here that the gradient descent policy performs best. We extend this scenario to real-world considerations such as fairness and cognitive overload.
Contribution Summary
Overall, we make three contributions to the patient-provider matching and assortment optimization literature: we i) develop a model of patient-provider matching under assortments, ii) characterize the performance of assortment policies, and detail how the best-performing policy depends on problem characteristics such as patient match probability and patient/provider ratio, and iii) make recommendations for the implementation of a real-world patient-provider matching system based on the results of our theoretical and empirical analysis.
Figure 1: A) In this work, we match patients with providers B) by offering a set of assortments to patients, where each assortment consists of a set of providers. C) Patients then respond sequentially and either select a provider from their assortments or abstain. D) This results in matches between patients and providers.
1.2 Related Work
Matching Patients and Providers Care continuity is critical for patient health, as it allows for better patient communication, lowers operating costs, and improves provider teamwork (Plomondon et al., 2007). Despite this, provider turnover rates are high, with the average healthcare system reporting a 7% turnover rate year-to-year (Plomondon et al., 2007). Moreover, such issues are exacerbated recently due to high rates of provider burnout (Shanafelt et al., 2012) and the Covid-19 pandemic (Shanafelt et al., 2022). Once a provider drops out, it is difficult for patients to find a new one; while 54% of patients find their new provider within 12 months, 6% fail to find a new provider even after 36 months (Hedden et al., 2021). High levels of provider turnover can worsen primary care because patients lack access to routine checkups or reminders to take medication (Reddy et al., 2015; Levine et al., 2019; Nguyen et al., 2020; Alsan et al., 2019).
While provider turnover impacts patient satisfaction, quantifying “match quality” is hard. Patients care mainly about geographic proximity (Dunlea and Lenert, 2015), with concordance along race (Greenwood et al., 2020), gender (Greenwood et al., 2018), and language (Manson, 1988) playing a secondary role. Communication style alignment can improve match quality (Street et al., 2008). These factors build trust and lead patients to more truthfully report symptoms (Persky et al., 2013).
Algorithms for Patient-Provider Matching Existing work in patient-provider matching uses either a one-shot matching framework or a genetic programming-based approach. Within the one-shot framework, linear programs can manage provider panel sizes (Harrington et al., 2021), the deferred acceptance framework can perform two-stage matching (Chen et al., 2020), and scheduling algorithms are used to model optimal rates of on-demand scheduling (Qu et al., 2007).
A second approach to patient-provider matching builds matches through genetic programming. Genetic programming algorithms learn matches over time through a fitness function. Within this framework, one line of work uses genetic programming to find fair matches between patients and providers (Chen et al., 2020), while another approach uses genetic programming to balance workloads between providers (Zhu et al., 2023). We build on these lines of work by incorporating patient autonomy into the process, which motivates the need for an assortment-based approach.
Assortment Optimization Our work analyzes patient-provider matching through assortment optimization, where retailers offer a set of menus to customers. In assortment optimization, retailers construct assortments and customers make decisions based on offered options (Bertsimas and Mišic, 2015). Retailers first construct models of customer decisions, such as the multinomial logit model (Davis et al., 2013) and nested logit models (Alfandari et al., 2021), both of which use logit probabilities to model customer decisions. Assortments are employed in domains including school choice (Shi, 2016), dating markets (Rios and Torrico, 2023), and e-commerce (Tian et al., 2024).
Assortment optimization algorithms come in offline and online variants. In offline assortment optimization, assortments are offered one-shot and customers make decisions simultaneously and independently (Bertsimas and Mišic, 2015). In online assortment optimization, customers arrive sequentially, and options are offered for each (Aouad and Saban, 2023). Our setting resides between offline and online, as we offer assortments offline, but patients respond online.
Matching under Random Ordering Patient-provider matching is challenging from an optimization perspective because patients make decisions in a random order. Randomized response orders are found in variants of online selection problems such as secretary problems (Freeman, 1983) and online learning (Orabona, 2019). One key difference is the lack of flexibility in our scenario; changing assortments online for the patient-provider matching setting would involve considerable logistical hassle, so we offer assortments in a one-shot manner.
2 Problem Formulation
2.1 Formal Model of Patient-Provider Matching
We focus on matching $N$ patients to $M$ providers. We present each patient $i$ with an assortment $\mathbf{X}_{i}\in\{0,1\}^{M}$ upfront. Here, $X_{i,j}$ denotes whether patient $i$ has provider $j$ in their assortment. Patients then select providers from their assortment sequentially by responding in an order $\sigma$, and can abstain from provider selection. Here, $\sigma_{1}\in[N]$ represents the first patient in the order. We assume patient response times are i.i.d. and exponential; prior work demonstrates this can model real-world wait times (Newell, 2013). Under this assumption, $\sigma$ is a random permutation of $[N]$.
Each patient selects a provider from their assortment according to a choice model $f_{i}:\{0,1\}^{M}\rightarrow\{0,1\}^{M}$. The choice model takes as input a 0-1 vector $\mathbf{z}$ and outputs a random 0-1 vector representing the provider selected by patient $i$. Formally, let $\mathbf{y}^{(t)}$ represent the set of unselected providers at time $t$. Then, patient $\sigma_{t}$ selects providers according to $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})$, where $\odot$ is the element-wise product ($\mathbf{a}\odot\mathbf{b}=\mathbf{c}\rightarrow a_{k}b_{k}=c_{k}\forall k$). $\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}$ represents the set of providers that are on patient $\sigma_{t}$’s assortment and unselected by previous patients. If $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})=\mathbf{0}$, then no provider is selected, while if $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})_{j}=1$, then provider $j$ is selected. We note that a patient can match with at most one provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\leq 1$
Patients are more likely to select providers with a higher match quality, $\theta_{i,j}\in[0,1]$, where $\theta_{i,j}$ is analogous to a reward for patient $i$ when selecting provider $j$. Match quality can encompass various factors, including demographic concordance (Greenwood et al., 2020, 2018; Manson, 1988), physical proximity, and patient needs. We assume administrators know $\theta_{i,j}$. Prior work demonstrates providers are able to successfully identify new high-match quality providers for patients (Lohr et al., 2013), which implies that we can predict match quality from patient data.
We present examples of choice models below:
  1. 1.
    Uniform - Patients select their most preferred provider with probability $p$ and otherwise abstain. Let $\mathbf{e}_{k}$ denote the $k$-th standard basis vector, then:
    $f_{i}(\mathbf{z})=\left\{\begin{array}[]{ll}\mathbf{e}_{\operatorname*{arg\,max}_{j}(\theta_{i}\odot\mathbf{z})_{j}}&\text{with prb. p}\\ \mathbf{0}&\text{otherwise}\end{array}\right.$ (1)
  2. 2.
    Threshold - Patients follow the uniform choice model but only select providers with match quality above some threshold $\alpha$. If match quality corresponds to geographic proximity, $\alpha$ corresponds to maximum travel distance. Formally:
    $f_{i}(\mathbf{z})=\left\{\begin{array}[]{ll}\mathbf{e}_{\operatorname*{arg\,max}_{j}(\theta_{i}\odot\mathbf{z})_{j}}&\text{with prb. p if }\max\theta_{i}\odot\mathbf{z}\geq\alpha\\ \mathbf{0}&\text{otherwise}\end{array}\right.$ (2)
  3. 3.
    Multinomial Logit (MNL) - Patients select providers according to $\theta$, and have an exit option $\gamma$ which represents selecting no provider. Under the MNL choice model (Train, 2009), this is:
    $f_{i}(\mathbf{z})=\left\{\begin{array}[]{ll}\mathbf{e}_{j}&\text{with prb. }\frac{z_{j}\exp(\theta_{i,j})}{\exp(\gamma)+\sum_{j^{\prime}}z_{j^{\prime}}\exp(\theta_{i,j^{\prime}})}\\ \\ \mathbf{0}&\text{with prb. }\frac{\exp(\gamma)}{\exp(\gamma)+\sum_{j^{\prime}}z_{j^{\prime}}\exp(\theta_{i,j^{\prime}})}\end{array}\right.$ (3)
We aim to find a policy $\pi(\theta)$ that constructs an assortment $\mathbf{X}$, given the match quality matrix $\theta$. We evaluate policies through two metrics: match quality and match rate.
  1. 1.
    Match Quality measures $\theta$ across selected patient-provider pairs. Formally, it aggregates $\theta_{\sigma_{t}}$ across provider selections $f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})$ for all $N$ patients (= $N$ time periods):
    $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})\cdot\theta_{\sigma_{t}}]$ (4)
  2. 2.
    Match Rate is the proportion of patients who are matched. Formally, it computes if patient $\sigma_{t}$ matches with a provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\geq 0$:
    $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})\rVert_{1}]$ (5)
Example 1.
We quantify the example in Figure 1. Suppose $N=M=3$, and let $\theta$ and $\mathbf{X}$ be
$\theta=\begin{bmatrix}0.5&0.6&0.9\\ 0.7&0.5&0.4\\ 0.6&0.2&0.8\end{bmatrix},\hskip 14.22636pt\mathbf{X}=\begin{bmatrix}0&0&1\\ 1&1&0\\ 1&0&1\end{bmatrix}$ (6)
Suppose $\sigma=[2,1,3]$. Then one trial could be the following: patient $\sigma_{1}=2$ abstains from selecting a provider, patient $1$ selects provider $3$ and patient $3$ selects provider $1$: $f_{2}(\mathbf{y}^{(1)}\odot\mathbf{X}_{2})=\mathbf{0}$, $f_{1}(\mathbf{y}^{(2)}\odot\mathbf{X}_{1})=\mathbf{e}_{3}$, and $f_{3}(\mathbf{y}^{(3)}\odot\mathbf{X}_{3})=\mathbf{e}_{1}$. Here, $\mathbf{y}^{(1)}=[1,1,1]$, $\mathbf{y}^{(2)}=[1,1,1]$, and $\mathbf{y}^{(3)}=[1,1,0]$. The resulting match quality is $\frac{1}{3}(\theta_{1,3}+\theta_{3,1})=0.5$ and match rate is $\frac{2}{3}$
2.2 Model Fidelity
We discuss our model fidelity for real-world patient-provider matching:
  1. 1.
    Patient autonomy is a critical part of our model, as assortments allow patients to select between providers while allowing administrators some control. Prior work demonstrated that people are willing to trade autonomy for efficiency, including in online markets (Diederich et al., 2025) and domestic life (Munro et al., 2014). We find similar ideas during our conversations with our healthcare partners, where patients might prefer a few high-quality choices to an overwhelmingly large number of options (Chernev et al., 2015).
  2. 2.
    Match quality and match rate underlie assortment computation. We focus on match quality because it reflects the benefit patients get from matching, while match rate reflects the fraction of patients with access to providers. For example, if patients primarily care about proximity, match quality should be based on the distance between patients and providers. We discuss different notions of match quality in Section 5.3.1.
  3. 3.
    Choice models reflect how patients select providers. Choice models are critical to our setting, so we analyze various choice models in Section 3 and experiment with different selections in Section 5.2. We focus on choice models that are clinically motivated or ubiquitous in prior work. For example, the uniform choice model reflects low-effort patient decisions (Salisbury, 1989), while the MNL choice model is commonly seen in prior work (Davis et al., 2013).
2.3 Analysis in the Single Provider Scenario
To gain insight into the optimal match quality policy structure, we analyze the $M=1$ scenario under the uniform choice model. We first explicitly detail the optimal policy: {theorem}[] Let $f_{i}$ be the uniform choice model with probability $p$. Suppose $M=1$, and let $u_{1},u_{2},\ldots,u_{N}$ be a set of coefficients such that $\theta_{u_{1}}\geq\theta_{u_{2}}\cdots\theta_{u_{N}}$. Let $s$ be defined as follows:
$s=\operatorname*{arg\,max}_{s}(1-(1-p)^{s})\frac{\sum_{i=1}^{N}\theta_{u_{i},1}}{s}$ (7)
Then the policy which maximizes match quality is $\mathbf{X}_{u_{1},1}=\mathbf{X}_{u_{2},1}\cdots\mathbf{X}_{u_{s},1}=1$, where $\mathbf{X}_{i,1}$ is 0 otherwise. When $M=1$, we can decompose the match quality into the probability any patient matches and the average match quality given matching. If we offer the single provider to $s$ patients, then the match probability is $(1-(1-p)^{s})$. For fixed $s$, we offer provider $j=1$ to the top-$s$ patients by match quality. Note that the optimal match rate is achieved with $s=N$, while the optimal match quality might arise from $s<N$ (as we’ll show in Section 3.5). We include all full proofs in Appendix B.
3 Constructing Matching Policies
We introduce policies $\pi(\theta)$ to construct assortments. We first introduce a baseline greedy policy and discuss why such a policy can perform poorly. We then propose three new policies: pairwise, which assigns each provider a single patient, group-based, which improves upon the weaknesses of the pairwise policy, and gradient descent, which optimizes for a lower bound on the match quality.
3.1 Greedy Policy
The greedy policy offers all providers to each patient and grants patients full autonomy. Formally, the greedy policy is $\pi^{R}(\theta)_{i,j}=1$, for every patient $i$ and provider $j$. The greedy policy maximizes the match rate under the uniform choice model because $\mathbf{X}_{\sigma_{t}}=\mathbf{1}$, and so $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}=\lVert f_{\sigma_{t}}(\mathbf{y}^{(t)})\rVert_{1}$.
For match quality, the greedy policy considers assortments myopically and ignores the structure of the match quality matrix, which can lead to poor performance. For example, if $M=1$ and $\theta_{1,1}>>\theta_{1,2}$, the greedy policy gives both patients equal chances of matching, whereas the optimal policy offers $j=1$ only to patient $i=1$. Formally, let $\pi^{*}$ be the optimal policy, $\pi^{*}(\theta)=\operatorname*{arg\,max}_{\mathbf{X}}\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$. We first prove that $\pi^{R}$ is an $\epsilon$-approximation:
{theorem}
[] Let $f_{i}$ be the uniform choice model with match probability $p$. For any $p$ and $\epsilon$, there exists a $\theta$ such that
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{R}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq\epsilon\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (8)
We prove this by generalizing the example where $\theta_{1,1}>>\theta_{2,1}$ to arbitrary $N$ and $M$. Here, the optimal policy is to offer each provider to only one patient. However, the greedy policy offers all providers to all patients, which leads to sub-optimal matches for each provider $j$. Moreover, we show that we can construct scenarios with $N=M=\frac{3}{\epsilon p}$ where the greedy policy is an $\epsilon$-approximation. For large $p$, even small $M$ and $N$ lead to situations where the greedy policy performs poorly.
3.2 Pairwise Policy
The pairwise policy inverts the greedy policy by offering each patient at most one provider. This strategy reduces patient choices while better incorporating match quality structure. The pairwise policy, $\pi^{P}(\theta)$, pairs patients and providers by solving the weighted bipartite matching problem:
$\max\limits_{\pi^{P}(\theta),\sum_{i=1}^{N}\pi^{P}(\theta)_{i,j}\leq 1,\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\leq 1}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (9)
Under a uniform choice model, the choice of $p$ impacts the pairwise policy performance: {theorem}[] Let $f_{i}$ be the uniform choice model with match probability $p$. Then
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\geq p\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (10)
To prove this, we first upper bound the optimal policy, $\pi^{*}(\theta)$, with the value of the underlying bipartite matching problem. We then show that the pairwise policy corresponds to the bipartite matching problem, where edges correspond to pairs of patients and providers, and each edge exists with probability $p$. Summing over edges gives that the pairwise policy is a $p$-approximation to the bipartite matching problem and is, therefore, a $p$-approximation to $\pi^{*}$.
Beyond the uniform choice model, we can show a similar guarantee in the MNL scenario. For the MNL choice model, the exit option $\gamma$ is analogous to $p$ in controlling policy performance. First, let $v_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$ and let $v_{i}=-1$ if $\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}=0$. Then the MNL choice model is equivalent to the uniform choice model with $p_{i}=\frac{\mathbbm{1}[v_{i}\geq 0]\exp(\theta_{i,v_{i}})}{\exp(\gamma)+\mathbbm{1}[v_{i}\geq 0]\exp(\theta_{i,v_{i}})}$. We can then upper bound optimal match quality as $\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$, while the pairwise policy achieves $\sum_{i=1}^{N}\frac{p_{i}}{N}\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$. Let $a_{i}=\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$, then the pairwise policy achieves an approximation ratio of:
$\frac{\sum_{i=1}^{N}a_{i}\frac{\exp(a_{i})}{\exp(\gamma)+\exp(a_{i})}}{\sum_{i=1}^{N}a_{i}}\geq\frac{\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})(\frac{1}{N}\sum_{i=1}^{N}a_{i})}{(\exp(\gamma)+\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i}))(\frac{1}{N}\sum_{i=1}^{N}a_{i})}\geq\frac{\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})}{\exp(\gamma)+\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})}$ (11)
Small $\gamma$ improves pairwise performance, corresponding to patients more inclined to match.
3.3 Group-Based Policy
The pairwise policy performs poorly for small $p$, so we develop the group-based policy. When $p$ is small, administrators should offer patients larger assortments because a patient’s top preference is more likely to be available. For example, consider a scenario with $N$ patients and $M$ providers. We demonstrate a performance gap between greedy and pairwise policy dependent on $p$:
Proposition 3.1.
Let $f_{i}$ be the uniform choice model with match probability $p$. If $\theta_{i,j}\sim U(0,1)$ and $M\leq N$ then
$\frac{\mathbb{E}_{\sigma,\theta}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{R}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]}{\mathbb{E}_{\sigma,\theta}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]}\geq\frac{1-(1-p)^{N/M}}{2p}$ (12)
We demonstrate this by showing that the pairwise policy achieves a match quality of $Mp$. We then compute the match probability for the greedy policy as $1-(1-p)^{N/M}$ for each of the $M$ providers, and noting that the match quality for each matched provider is at least $\frac{1}{2}$ in expectation. The pairwise policy performs poorly for small $p$, so we develop a new policy that modulates between the pairwise policy (for large $p$) and the greedy policy (for small $p$). To do this, we add a provider $j^{\prime}$ to the assortment of patient $i$ if $\theta_{i,j^{\prime}}\geq\theta_{i,v_{i}}$; that is, patients should be assigned preferred providers. However, such a policy could decrease match rates because some patients have empty assortments.
Example 2.
Let $\theta$ be the following:
$\theta=\begin{bmatrix}0.6&0.7\\ 0.3&0.6\\ \end{bmatrix}$ (13)
Here, $\pi^{P}(\theta)_{1,1}=\pi^{P}(\theta)_{2,2}=1$. We add provider $j=2$ to patient $i=1$, so $X_{i,2}=1$. Let $\sigma=[1,2]$. Then, with probability $p$, provider $j=2$ is selected by patient $i=1$ and $\mathbf{y}^{(2)}=[1,0]$ so patient $i=2$ has no options. This reduces the match rate from $p$ to $p-\frac{p^{2}}{2}$.
We overcome this issue by developing the group-based policy, $\pi^{G}(\theta)$. The group-based policy constructs groups so that each group receives the same assortment. In Example 2, this would mean that $\mathbf{X}_{1}=\mathbf{X}_{2}=[1,1]$. We demonstrate that such a grouping strategy is the only way to maintain a match rate while building upon the pairwise policy: {theorem}[] Let $\pi$ be a policy that augments the pairwise policy; $\pi(\theta)_{i,j}\geq\pi^{P}(\theta)_{i,j}$ for any $\theta$ for all $i,j$. Let $G=(V,E)$ be a directed graph with $N$ nodes such that nodes $i$ and $i^{\prime}$ are connected if $\pi(\theta)_{i,v(\theta)_{i^{\prime}}}=1$ Here, $v(\theta)_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$. If $N=M$, then
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}\left(\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\rVert_{1}]=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\rVert_{1}]=p$ (14)
if and only if each component in $G$ is a complete digraph. We prove this by showing that when $N=M$, all patients can match with a provider with probability $p$, leading to an overall match rate of $p$. To construct groups, we first compute edge weights for each pair of patients in $G$, denoted as $\alpha_{i,i^{\prime}}$. Here, $\alpha_{i,i^{\prime}}$ captures the benefit when placing $(i,v_{i})$ and $(i^{\prime},v_{i^{\prime}})$ in the same assortment. Next, we construct groups to maximize the sum of pairwise edges within each group. We do so through a linear program, with $z_{i,j}$ representing whether we include edge $(i,j)$ and $q_{i}$ representing whether we include node $i$ in the group. We place details on computing $\alpha_{i,i^{\prime}}$ and constructing the assortments in Algorithm 1.
Algorithm 1 Group-based policy ($\pi^{G}$)
1:  Input: Pairwise policy, $\pi^{P}$ and match quality $\theta$
2:  Output: Assortment, $\mathbf{X}$
3:  Initialize $\mathbf{X}=\mathbf{X}^{P}=\pi^{P}(\theta)$, $v_{i}=j$ if $X_{i,j}^{P}=1$, and $\mathbf{c}=\mathbf{1}$
4:  for all $(i,i^{\prime})\subseteq[N]$ do
5:     Let $\mathbf{X}^{\prime}=\mathbf{X}^{P}$
6:     Let ${X}^{\prime}_{i,v_{i^{\prime}}}=1$ and $X^{\prime}_{i^{\prime},v_{i}}=1$
7:     Let $\alpha_{i,i^{\prime}}=\mathbb{E}_{\sigma\sim[\{i,i^{\prime}\},\{i^{\prime},i\}]}[\sum_{t=1}^{2}(f_{\sigma_{t}}\left(\mathbf{X}^{\prime}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}-f_{\sigma_{t}}\left(\mathbf{X}^{P}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$
8:  end for
9:  Let $s=\infty$ and $r=\{1,\ldots,N\}$
10:  while s ¿ 0 do
11:     Let $s=\max\limits_{\mathbf{z}}\sum_{i\in r}\sum_{i^{\prime}\in r}z_{i,j}\alpha_{i,i^{\prime}}$, subject to $z_{i,j}=q_{i}q_{j}$, and let $\mathbf{z}$ be the corresponding solution
12:     for $i,i^{\prime}\in r$ with $z_{i,i^{\prime}}=1$ do
13:        Let $X_{i,v_{i^{\prime}}}=X_{i^{\prime},v_{i}}=1$
14:        Remove $i,i^{\prime}$ from $r$
15:     end for
16:  end while
3.4 Gradient Descent Policy
The group-based policy makes local improvements to the pairwise policy but can fail because it only considers pairwise interactions. To fix this, we develop the gradient descent policy by considering the global structure in $\theta$. Our gradient descent policy uses gradient descent to optimize an objective function $g(f(\mathbf{X}))$, which represents a lower bound on the match quality of an assortment $\mathbf{X}$ under the uniform choice model. Here, $f(\mathbf{X})_{i,j}$ lower bounds the probability that provider $j$ is available, while $g(f(\mathbf{X}))_{i,j}\in[0,1]^{N\times M}$ lower bounds the probability that provider $j$ is the top choice for patient $i$: $g(f(\mathbf{X}))\leq\mathrm{Pr}[j=\operatorname*{arg\,max}_{j^{\prime}}\theta_{i,j^{\prime}}\mathbf{y}^{(t)}_{j^{\prime}}]$. We first detail how to compute $f$ and $g$, then prove that this lower bounds the match quality (see Algorithm 2):
  1. 1.
    $f(\mathbf{X})$ lower bounds the availability probability for provider $j$. We first note that $f(\mathbf{X})_{i,j}=0$ if $X_{i,j}=0$. Next, availability depends on the response order $\sigma$. For example, if $\sigma_{1}=i$, then provider $j$ is available if $X_{i,j}=1$. Next, if $\sigma_{t}=i$, then under a uniform choice model, there is at least a $(1-p)^{t-1}$ availability probability. This holds because, in the worst case, the $t-1$ preceding patients also prefer provider $j$, so the probability none of the $t-1$ patients select $j$ is $(1-p)^{t-1}$. Next, if $n$ out of $N$ patients have provider $j$ on their assortment, then the probability that a given patient $i^{\prime}\neq i$ matches to $j$ is $p\frac{n-1}{N-1}$. That is, if patient $i$ is in the t-th position, then the availability probability is at least $(1-p\frac{n-1}{N-1})^{t-1}$. Averaging over all $N$ potential values for $t$ gives:
    $f(\mathbf{X})_{i,j}=X_{i,j}h(\lVert\mathbf{X}_{*,j}\rVert_{1}),h(n)=\frac{1}{N}\sum_{t=1}^{N}(1-p\frac{n-1}{N-1})^{t-1}$ (15)
  2. 2.
    $g(f(\mathbf{X}))$ computes the probability that provider $j$ is the top available provider. Let $u_{i,1},u_{i,2},\ldots,u_{i,N}$ be coefficients so $\theta_{i,u_{1}}\geq\theta_{i,u_{2}}\cdots\theta_{i,u_{M}}$. Then provider $u_{i,1}$ is the top option whenever it is available; $g(f(\mathbf{X}))_{i,u_{i,1}})=f(\mathbf{X})_{i,u_{i,1}}$. Similarly, $u_{i,2}$ is only selected if $u_{i,1}$ is not selected; we can estimate $g(f(\mathbf{X}))_{i,u_{i,2}})=f(\mathbf{X})_{i,u_{i,2}}(1-f(\mathbf{X})_{i,u_{i,1}})$. We can generalize this as:
    $g(f(\mathbf{X}))_{i,u_{i,k}}=f(\mathbf{X})_{i,u_{i,k}}\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{i,k^{\prime}}})$ (16)
    The expected match quality is then $p(g(f(\mathbf{X}))\cdot\theta)$.
Algorithm 2 Gradient Descent Objective
  Input: Match quality $\theta$ and Assortment $\mathbf{X}$
  Output: Estimated total match quality
  Let $h(n)=\frac{1}{N}\sum_{t=1}^{N}(1-p\frac{n-1}{N-1})^{t-1}$
  Let $f(\mathbf{X})_{i,j}=h(\lVert\mathbf{X}_{*,j}\rVert_{1})\mathbf{X}_{i,j}$
  Let $u_{i,1},u_{i,2},\ldots,u_{i,M}$ be a permutation of $[M]$, so $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,1}}\cdots\theta_{i,u_{i,M}}$
  Let $g(f(\mathbf{X}))_{i,u_{k}}=f(\mathbf{X})_{i,u_{k}}\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{k^{\prime}}})$
  return  $p*g(f(\mathbf{X}))\cdot\theta$
{theorem}
The following holds for any $\mathbf{X}$:
$p*(g(f(\mathbf{X}))\cdot\theta)\leq\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (17)
Moreover, we demonstrate tightness when using the pairwise policy: {lemma}[] When $\mathbf{X}=\pi^{P}(\theta)$, then
$p*(g(f(\mathbf{X}))\cdot\theta)=\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (18)
We demonstrate this by showing that for any $i$, $\sum_{j=1}^{m}pg(f(\mathbf{X}))_{i,j}$ is an underestimation for the probability that any of the top-$m$ providers are available. We then show that because $\theta_{i,u_{i,j}}$ is sorted from high to low, we can similarly prove that $\sum_{j=1}^{m}p(g(f(\mathbf{X}))_{i,j}\theta_{i,j}$ is also an underestimation for each patient $i$. In the scenario where $\mathbf{X}=\pi^{P}(\theta)$, we can explicitly compute $g(f(\mathbf{X}))$ because each patient matches with at most one provider.
Key to our policy is the tightness of the lower bound. When each patient receives more providers the lower bound is looser so the gradient descent policy performs worse. Conversely, when $N>>M$, each patient can only receive a few providers, so our policy performs better. Essentially, when there is low “inter-provider interference”, or patients receiving large assortments (relative to the number of providers), our policy performs better. The gradient descent policy can also naturally extend to optimize match rate if $\theta_{i,j}$ is constant; in that scenario, we optimize for $\sum_{i=1}^{N}\sum_{j=1}^{M}g(f(\mathbf{X}))_{i,j}$.
3.5 Worked Example with Policies
We motivate the need for different policies beyond greedy through a simple example with $N=3$ patients and $M=1$ providers.
Example 3.
Consider a scenario with $p=0.75$ and $\theta=[0.7,0.7,0.1]$. According to Theorem 2.3, the optimal assortment in such a scenario is $\mathbf{X}=[1,1,0]$, which achieves a match quality of $0.22$. We next compare our four policies:
  1. 1.
    Greedy offers $\mathbf{X}=[1,1,1]$, which leads to an average match quality of $\frac{(1-(1-p)^{N})\frac{1}{N}\sum_{i=1}^{N}\theta_{i,1}}{N}=0.16$. Greedy achieves a low match quality because it offers a provider to patient $i=3$, which drags down the average match quality.
  2. 2.
    Pairwise offers $\mathbf{X}=[1,0,0]$, which achieves a match quality of $0.18$. While the pairwise policy improves upon the greedy policy, the symmetry between patients one and two implies that it should offer a provider to patient $i=2$, as it can improve the match probability.
  3. 3.
    Group-Based first computes $\alpha_{1,2}=0.13$ and $\alpha_{1,3}=-0.15$ and constructs a group with patients 1 and 2. It then offers $\mathbf{X}=[1,1,0]$, the optimal assortment for match quality.
  4. 4.
    Gradient Descent optimizes for $g(f(\mathbf{X}))\cdot\theta=f(\mathbf{X})\cdot\theta=\sum_{i=1}^{N}X_{i,1}\theta_{i,1}h(\lVert\mathbf{X}_{*,j}\rVert_{1})$, which is optimizes at $\mathbf{X}=[1,1,0]$ and is the optimal assortment for match quality.
The greedy policy fails because it offers the provider to patients with a low match quality, while the pairwise policy fails because it offers providers to too few patients. The group-based and gradient descent policies rectify this situation by adapting assortment size based on the match quality and correctly offering the provider to patients 1 and 2.
4 Extending Policies to Real-World Considerations
We extend our model to incorporate three real-world matching phenomena: 1) batch-offering assortments, 2) patient cognitive load, and 3) fairness considerations.
4.1 Batch Offering Patients
Thus far, we have assumed that all menus are constructed up front, in a single “batch.” Healthcare administrators can sacrifice logistical ease to improve match quality by making offers to groups of patients in batches rather than all at once. By offering assortments in batches, administrators increase control over the response order but increase the total response time for all patients to respond. By assumption, the response time for each patient is exponentially distributed (see Section 2.1), and the expected maximum of $N$ exponential variables is proportional to $\log(N)$ (Eisenberg, 2008). By offering $L$ batches, our total response time increases from $\log(N)$ to $L\log(N/L)$.
We aim to construct batches, $b_{1},b_{2},\ldots,b_{L}\subseteq[N]$, so that the resulting response order maximizes match quality. Formally, let $S(b_{1},b_{2},\ldots,b_{L})$ be a distribution over response orders so that any patient $i\in b_{k}$ is before any patient $i^{\prime}\in b_{k^{\prime}}$ with $k<k^{\prime}$. Next, let $\sigma^{*}$ be the optimal ordering for some fixed policy $\pi$: $\sigma^{*}=\operatorname*{arg\,max}_{\sigma}\frac{1}{N}\sum_{t=1}^{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi(\theta,\sigma)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})$. We aim to set $b_{1},b_{2},\ldots,b_{L}$ so that $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ and $\sigma^{*}$ are similar.
To do so, we first characterize $\sigma^{*}$ for a policy $\pi^{A}(\theta,\sigma)$, which augments the pairwise policy using knowledge of $\sigma$, $\pi^{A}(\theta,\sigma)_{i,j}\geq\pi^{P}(\theta)_{i,j}\forall i,j$. Recall that $v_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$, and let $v^{-1}_{j}=i$ if $v_{i}=j$. Then, $\pi^{A}(\theta)_{i,j}=1$ if $\theta_{i,j}\geq\theta_{i,v_{i}}$ and $\sigma^{-1}_{i}\geq\sigma^{-1}_{v^{-1}_{j}}$, where $\sigma^{-1}_{i}=t$ if $\sigma_{t}=i$. Essentially, $\pi^{A}(\theta)$ augments the pairwise policy $\pi^{P}(\theta)$ to add pairs $i,v_{i^{\prime}}$ for patients $i^{\prime}$ preceding $i$ in $\sigma$. We focus on such a policy because it represents a natural extension of the pairwise policy to incorporate the order $\sigma$. We characterize $\sigma^{*}$ for the policy $\pi^{A}(\theta)$ then use this to construct the batches $b_{k}$:
Lemma 4.1.
Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v_{i}=-1$ and $v_{i^{\prime}}\neq-1$. Let $\sigma^{*}$ be the optimal ordering for the policy $\pi^{A}$. Then, traversing the nodes defined by $\sigma^{*}$ is a reverse topological ordering on $G$.
We demonstrate this by considering a pair of patients $i$ and $i^{\prime}$, so that $i$ precedes $i^{\prime}$ in any reverse topological order, but $i$ precedes $i^{\prime}$ in $\sigma^{*}$. We then use case work to show that we can either a) show that there is a cycle in the graph $G$, leading to a contradiction, b) recurse on a smaller pair of patients, $w,i$ or $i^{\prime},w$, or c) show that swapping $i$ and $i^{\prime}$ in $\sigma$ can only increase the match quality.
We next construct $b_{1},b_{2},\ldots,b_{L}$, so $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ is close to the optimal ordering $\sigma^{*}$. The idea is to partition $\sigma^{*}$ into $L$ batches; then we know that all patients in $b_{1}$ precede $b_{2}$. To do this, we first demonstrate that if no pairs of patients within a batch share a common descendant, then such a batching preserves the match quality of $\sigma^{*}$: {theorem}[] Let $\sigma^{*}$ be an optimal ordering. Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v_{i}=-1$ and $v_{i^{\prime}}\neq-1$. Suppose that there exists a partition of $\sigma^{*}$ into $K$ batches, $b_{1},b_{2},\ldots,b_{L}$, such that for any partition $b_{k}$, no $i,i^{\prime}\in b_{k}$ have a common descendant in $G=(V,E)$. Then $\pi^{A}$ achieves the same match quality under partition $g$ as under the optimal ordering:
$\mathbb{E}_{\sigma\sim S(b_{1},b_{2},\ldots,b_{L})}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{A}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]=\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma^{*}_{t}}\left(\pi^{A}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})$ (19)
This holds because the selection made by patient $i$ depends only on the descendants of patient $i$ on the graph G. If $b_{1},b_{2},\ldots,b_{K}$ ensures that any pair of descendants are in different batches, then the order of descendants is maintained, so patients select the same providers in $\sigma$ and $\sigma^{*}$.
In practice, we compute the batches $b_{k}$ through the following:
  1. 1.
    Computing the graph $G=(V,E)$ with $i$ connected to $i^{\prime}$ if $\theta_{i,v_{i^{\prime}}}>\theta_{i,v_{i}}$.
  2. 2.
    Computing a reverse topological ordering $\sigma^{*}$.
  3. 3.
    Selecting batches with the minimum number of edges within a batch (via dynamic programming).
We note that this is an approximation of optimal batches for two reasons: 1) while $\sigma^{*}$ is a reverse topological ordering, there could be multiple reverse topological orderings, and 2) minimizing the number of edges within a batch does not guarantee that the order of descendants is maintained.
4.2 Maximum Assortment Sizes
To incorporate patient cognitive load (Chernev et al., 2015), we incorporate restrictions on the maximum assortment size into our model. Large assortments might be undesirable or practically infeasible because patients only consider a subset of the options they are presented due to logistical or time constraints. Formally, we suppose that patients only consider a random $d$-subset $\mathbf{z}^{(d)}$ of the options they are presented at random, $z^{(d)}_{j}\leq X_{i,j}y^{(t)}_{j}$ with $\lVert\mathbf{z}^{(d)}\rVert_{1}\leq d$. Patients make decisions according to $f_{i}(\mathbf{z}^{(d)})$ rather than $f_{i}(\pi(\theta)\odot\mathbf{y}^{(t)})$. When $d$ is small, policies should offer small tailored assortments, as large assortments lead to patients ignoring many options.
4.3 Fair Matches
Policies should aim to achieve fair matches to encourage participation and maintain equity. We construct examples of fairness metrics through the prior literature on fair matches and list these below (Xinying Chen and Hooker, 2023; García-Soriano and Bonchi, 2020).
  • Minimum Match Quality - One natural objective is to maximize the minimum match quality across all matched patients. Such an objective is natural when administrators aim to maintain a certain level of care across patients.
  • Variance in Match Quality - Another approach minimizes variance in match quality to ensure that the discrepancy in match quality between patients is small.
  • Match Quality Range - One final approach to match quality fairness is to minimize the absolute difference between the maximum and minimum match quality for matched patients.
Each notion of fairness has different implications for the optimal policy. For example, optimizing for the minimum match quality might lead policies to only offer assortments to high match quality patient-provider pairs, whereas minimizing the variance in match quality might yield many low quality matches. We explore tradeoffs in match quality and fairness in Section 5.4.
5 Empirical Analysis
We analyze policies in synthetic and real-world settings to characterize how assortment policy impacts match quality and rate. First, we demonstrate that gradient descent is the best policy when patients outnumber providers. We then show that group-based and greedy policies perform well for large and small values of match probability $p$, respectively. Our results demonstrate that assortment policies should be selected based on problem specifics.
5.1 Experimental Details
We compare the four policies from Section 3 (greedy, pairwise, group-based, and gradient descent) and a random baseline with $\pi(\theta)_{i,j}\sim\text{Ber}(\frac{1}{2})$. For a fixed problem setup, we randomly sample $T$ permutations of patients and compute the match quality and match rate for the corresponding patient order ($\sigma$). We average runs across 15 seeds and $T=100$ trials and plot the standard error across these seeds. Because the scale of match quality and rate might vary between experiments, we report the normalized match quality and match rate (norm. MQ and norm. MR for short), which divides each quantity by that of the random policy.
5.2 Synthetic Examples
We construct synthetic examples to gain insight into our policies across different scenarios. For each example, we specify values for $M$, $N$, $f_{i}$, and $\theta$.
Example 4.
Figure 2: We find the optimal policy and compute the approximation ratio for match quality compared to optimal. While all non-random policies achieve approximation ratios $\geq 0.9$ when $M\geq N$ (left), the pairwise policy performs poorly and achieves a low approximation ratio when $N\geq M$ (right).
Few patients and providers - We first analyze policies with a small number of patients $N$ and number of providers $M$. Small $N$ and $M$ allow us to compare to the optimal policy $\pi^{*}$ because we can compute $\pi^{*}$ by enumerating all $2^{NM}$ assortments and selecting the one that maximizes match quality. We run two experiments: the first fixes $N=2$ while varying $M$ from $2$ to $5$, and the second fixes $M=2$ while varying $N$ from $2$ to $5$. For both, we uniformly distribute $\theta$, $\theta_{i,j}\sim U(0,1)$, and let the choice model be uniform choice with $p=1/2$. We compute the match quality approximation ratio to understand how close each policy is to the optimal.
In Figure 2, we demonstrate that when $M\geq N$ (left), all non-random policies perform well and achieve within 3.5% of optimal, while for $N\geq M$ (right), the pairwise policy performs poorly. When $M\geq N$, providers are abundant so each patient can match with their most preferred provider. However, when $N\geq M$, the pairwise policy performs up to 28% worse than optimal. Providers are scarce in this scenario, so policies should offer the same provider to multiple patients. The pairwise policy only offers each provider one patient, which is suboptimal and leads to poor match quality.
Example 5.
Figure 3: When there are more patients than providers, the gradient descent policy performs best, both for uniformly and normally distributed $\theta$.
Varied patient/provider ratio - In real-world situations, patients outnumbers providers (HHS, 2013), so we analyze how the patient/provider ratio impacts policy performance. We fix the number of providers $M=25$, while varying the number of patients $N$ from $25$ up to $200$. We compare policies under two match quality distributions: the first uniformly distributes $\theta\sim U(0,1)$, while the second normally distributes $\theta_{i,j}\sim\mathcal{N}(\mu_{j},0.1)$. The former corresponds to heterogeneous preferences where patient preferences are non-correlated, while the latter corresponds to homogenous preferences, where patient preferences are influenced by some signal of provider quality $\mu_{j}$. For both experiments, we fix $p=0.5$ for the uniform choice model.
In Figure 3, we demonstrate that the gradient descent policy performs best when patients outnumber providers. When $N/M=8$, the gradient descent policy performs 5% better than alternatives for uniform $\theta$ and 9% for normal $\theta$. The gradient descent policy performs best for large $N/M$ because $g(f(\mathbf{X}))$ better approximates match quality when $N>M$; see Section 3.4 for details. Both greedy and group-based policies perform similarly, while the pairwise policy performs worst, mirroring trends from Example 4. In Appendix A, we similarly compare policies according to the match rate and find that the gradient descent and greedy policies both maximize the match rate.
Example 6.
Figure 4: The greedy policy performs poorly when $p=0.9$ compared to smaller $p$, for both uniform and normal $\theta$, while the pairwise policy performs worse for smaller $p$, especially when $\theta$ is normally distributed. The group-based policy gets around these issues and performs well when $p$ is large and small.
Impact of $p$ and $\theta$ We assess the impact of match quality $\theta$ and match probability $p$ on policy performance with balanced patient and provider counts. We fix $N=M=25$, vary $p\in\{0.1,0.25,0.5,0.75,0.9\}$, and distribute $\theta$ either uniformly or normally.
In Figure 4, we demonstrate that the group-based policy achieves high match quality. As noted in Proposition 3.1, the group-based policy can improve upon the pairwise policy when $p$ is small. The pairwise policy performs especially poorly when $\theta$ is normally distributed because patient preferences are homogeneous, as all patients prefer some provider $j^{*}=\operatorname*{arg\,max}_{j}\mu_{j}$. The pairwise policy only places $j^{*}$ in one patient’s assortment, whereas $j^{*}$ should optimally be in multiple assortments due to its preferability. We also note the greedy and group-based policies perform well across $p$ and $\theta$; both policies are always within 6% of the best-performing policy. In Appendix A, we demonstrate that policies achieve similar match rates across $\theta$ and $p$.
(a) Uniform
(b) Normal
Figure 5: We characterize which policy maximizes match quality when varying patient/provider ratio and match probability $p$ when $\theta$ is uniformly (left) and normally (right) distributed. The group-based policy performs best for large $p$, the greedy policy performs best for small $p$, and the gradient descent policy performs best when the patient/provider ratio is large.
To understand which policy performs best across $p$, $\theta$, and $N/M$, we plot the policy which maximizes match quality when varying $p\in\{0.1,0.3,0.5,0.7,0.9\}$ and $N/M\in\{1,2,3,4,5,6,7,8\}$ with $M=10$, and distributing $\theta$ uniformly (Figure 5a) and normally (Figure 5b). The gradient descent policy performs best for large patient/provider ratios, which matches the results from Example 5. The group-based policy performs well for small $p$ because the problem more closely resembles bipartite matching (see Theorem 9), while the greedy policy performs well for small $p$ because it becomes advantageous to offer larger assortments (see Proposition 3.1). We finally note the random policy performs best when $p=0.1$ and $N/M=1$ because policies have high variance in that situation, making it difficult to ascertain which policy performs best.
Example 7.
Varied Choice Model - We evaluate the impact of choice models on policy performance by investigating the MNL and threshold choice models. We fix $N=M=25$, let $\theta$ be distributed uniformly, and vary $\gamma\in\{0,0.1,0.25,0.5\}$ for the MNL choice model and vary $\alpha\in\{0,0.1,0.25,0.5,0.75\}$ with $p=0.5$ for the threshold choice model. .
Figure 6: We compare policy performances when varying the exit option $\gamma$ for the MNL choice model, and the threshold $\alpha$ for the threshold choice model. For the MNL choice model, when $\gamma$ is small, the greedy policy performs poorly, while for large $\gamma$, all non-random policies perform similarly. For the threshold choice model, all policies perform similarly across choices of $\alpha$.
In Figure 6, we show that $\gamma$ determines the performance of the greedy policy for the MNL choice model, while policies perform similarly across $\alpha$ for the threshold choice model. As noted in Section 3.2, the choice of $\gamma$ is inversely related to the setting of $p$, with small $\gamma$ analogous to large $p$. Similar to the trends in Figure 4, we find that the greedy policy performs 22% worse than the best policy when $\gamma=0.1$. For the threshold choice model, $\alpha$ has little impact on the best policy, as greedy, pairwise, and gradient descent perform similarly across $\alpha$. In Appendix A, we explore variants of the uniform choice model where the match probability $p$ is misspecified.
5.3 Real-World Dataset
We work with a partner healthcare organization to assess the performance of our policies in a simulation that mimics situations faced in practice.
5.3.1 Constructing Data from Healthcare Providers
We work with a large healthcare provider in Connecticut to develop a simulation that reflects their system dynamics. We select values of $N$, $M$, $\theta$, and $f_{i}$:
  1. 1.
    $N$ and $M$ We estimate $N=1225$ and $M=700$ from the average panel size and provider count of the organization.
  2. 2.
    Choice Model $f_{i}$ We select the choice model based on previous work which shows that patients are low-effort decision makers who primarily make decisions based on geographic proximity (Salisbury, 1989; Buzza et al., 2011). We incorporate these two factors and let $f_{i}$ be a threshold model, with $\alpha$ representing the maximum distance patients are willing to travel.
  3. 3.
    Match Quality $\theta$ We construct $\theta$ to reflect geographic proximity and the presence of comorbidities. Geographic proximity is the top factor impacting patient match quality (Yen and Mounts, 2013; Buzza et al., 2011), while patients with comorbidities are best served by providers with the corresponding specialized training. Formally, for a patient $i$ and provider $j$, let $d_{i,j}$ be the distance between the patient and provider, and let $\beta_{i,j}$ denote whether patient $i$’s comorbidity and $j$’s specialty match. For example, if patient $i$ has a heart-related comorbidity and provider $j$ has a cardiology speciality, then $\beta_{i,j}=1$. We manually match patient comorbidities to provider specialties. We therefore let match quality $\theta_{i,j}=\alpha+(1-\alpha)(\Delta\beta_{i,j}+(1-\Delta)(\frac{\bar{d}}{d_{i,j}}-1)$. Here, $\Delta$ weights between comorbidities and distance and $\alpha$ sets the threshold match quality for a patient at distance $d_{i,j}=\bar{d}$.
We compute $d_{i,j}$ by sampling patient locations from Connecticut zip codes and obtaining provider locations from a Medicare dataset (Medicare & Medicaid, 2025). We obtain $\beta_{i,j}$ from prior work on comorbidity rates  (Piccirillo et al., 2008) and using information on provider specialties from a Medicare dataset (Medicare & Medicaid, 2025). We let $p=0.75$ because most patients are low-effort, and $\Delta=\alpha=\frac{1}{2}$ to balance proximity and comorbidities. Finally, we set $\bar{d}=20.2$ since prior work demonstrates the average distance threshold for patients is 20.2 miles (Yen and Mounts, 2013).
5.3.2 Real-World Simulation Results
(a) Comparison
(b) Comorbidities
(c) Match Rate vs. Population
(d) Match Rate by Zipcode
Figure 7: (a) We construct a real-world scenario with $N=1225$ and $M=700$. We find that the gradient descent policy performs best because of the large patient/provider ratio. (b) The gradient descent policy achieves a higher match quality for patients with comorbidities because these patients match with specialist providers. (c) Geographically, the gradient descent policy tends to match patients from more populated zip codes, (d) with the highest match rates occurring in cities such as Hartford and New Haven.
In Figure 7a, we show that the gradient descent policy achieves the highest match quality and outperforms alternatives by 13%. This mirrors Figure 5, where gradient descent is the best policy for large patient/provider ratios. Next, we analyze which patients achieve high-quality matches under the gradient descent policy. We find that patients with comorbidities achieve higher match quality ($p<10^{-17}$) because they tend to match with specialist providers (Figure 7b). Additionally, patients from more populated zip codes achieve higher match rates (Figure 7c), potentially due to the lack of providers in rural areas (Hart et al., 2002). Pictorially, this corresponds to high match rates in urban areas such as Hartford and New Haven (Figure 7d).
5.4 Real-World Considerations
We extend our scenario to explore the real-world considerations introduced in Section 4: patient batching, maximum assortment sizes, and fairness considerations.
5.4.1 Batched Patients
(a) Batch Size
(b) Max Assortment Size
(c) Fairness
Figure 8: (a) Offering patients in two batches improves match quality by 7% while increasing response time by 90%. Beyond two batches, the increases in match quality are diminishing. (b) We simulate cognitive overload by assuming patients consider up to $d$ options. Smaller $d$ favors the gradient descent policy because it offers carefully tailored assortments without being too large. (c) While the gradient descent policy maximizes match quality, it achieves worse fairness because it neglects the worst-off patient. In contrast, pairwise and group-based policies achieve higher fairness by only suggesting high-quality matches.
Offering assortments to batches of patients rather than one-shot allows administrators to exercise some control over patient response orders. We apply the batching algorithm from Section 4.1 and analyze the impact of the number of batches, $L$, on the match quality in our real-world scenario. In Figure 8a, we demonstrate that batch offerings allow administrators to trade off logistical burdens for match quality. While offering assortments in two batches can improve match quality by 8% compared to the single-batch setting, the total response time also increases by 74%. Moreover, larger batch sizes lead to diminishing increases in match quality with a near linear increase in the time taken.
5.4.2 Varied Assortment Sizes
Patients might only consider a subset of providers offered due to choice overload (Chernev et al., 2015). To simulate this, we incorporate the maximum assortment size, $d$, to simulate choice overload (see Section 4.2 for details). In Figure 8b, we vary $d\in\{5,10,15,20,25\}$ and find that the gradient descent policy is the best policy across values of $d$, especially for small $d$. This occurs because the gradient descent policy offers smaller assortments compared to the greedy policy, and so restricting assortment size has less of an impact.
5.4.3 Fairness Considerations
To understand how policies trade-off between fairness and match quality, we plot the minimum match quality against the average match quality in Figure 8c (see Section 4.3 for a more thorough discussion of different notions of fairness). In our fairness computation we exclude patients who fail to match, because including them would make the fairness metric always 0 and would not reflect the difference in fairness across policies. We find that the fairest policy, pairwise, also achieves the lowest average match quality. Fairer policies maximize tend to match fewer patients, which keeps the minimum match quality high but lowers the average match quality. If healthcare administrators are focused on fairness as opposed to match quality, then they can select a fairer policy such as the pairwise policy. We compare policies across one definition of fairness for simplicity, but we note that in reality, there are multiple dimensions of fairness that need to be considered (see Section 4.3).
6 Conclusion and Discussion
The patient-provider relationship is key to quality healthcare, yet provider turnover frequently leads to patients being without providers. Our work studies how to re-match patients in such a scenario through assortment optimization. Our study proposes algorithmic approaches for patient-provider matching systems and offers analytical and empirical insights to inform system design. We further offer contributions to one-shot matching under random response order, an understudied phenomenon within the assortment literature. We conclude by providing recommendations for deploying assortment-based systems into practice and discussing extensions of our work.
Recommendations for Deployment
We translate our takeaways into real-world implications:
  1. 1.
    Larger assortments become increasingly useful as patients become more selective because patients are more likely to reject matches, making it important to give more options. Moreover, assortment sizes should be tailored to patient behavior; the choice model and match probability impact which policy performs best.
  2. 2.
    The level of heterogeneity impacts the amount of work needed to design a good algorithm. For example, patients with diverse preferences require less work to find the ideal match, and even relatively simple baselines would work well. However, homogeneous scenarios (e.g., a small city with few providers) require care when matching patients and providers. For example, when patients all prefer the same provider, simple solutions such as the pairwise policy can perform poorly.
  3. 3.
    Healthcare administrators should be aware of the underlying factors that dictate which patients get matched. For example, providers are scarce in rural areas, so assortments might naturally lead to low match rates in rural areas. Algorithm designers should work with administrators during development to minimize these biases.
Model Extensions
We detail several extensions of this model that better capture real-world patient-provider matching. As noted in Section 2.1, we assume that the response time for each patient is i.i.d. A natural extension would be to develop policies that work well independent of the response order distribution. Our experiments also assume that the choice model, $f_{i}$, is constant across patients. However, in reality, patients might have heterogeneous choice models. For example, in the uniform choice model, the value of $p$ could vary across patients. It would be interesting to see how extensions of our policies would perform under heterogeneous choice models.
Extensions to Other Domains
Future work can extend our framework to other scenarios with random response orders. Examples include matching between students and courses, where universities could offer each student an assortment of courses, and between schools and children, where school choice might lead counties to offer each child multiple options (Chen and Sönmez, 2006).
Acknowledgements
We thank Aakash Lahoti, Keegan Harris, and George Chen for their comments. We thank the reviewers at ML4H for their feedback on a preliminary version of this work. We thank Susan Barrett and Karen Hall for their input from a health system perspective. Co-author Raman is supported by an NSF GRFP Fellowship.
Appendix A Extended Experiments
More patients than providers
Figure 9: We compare the match rate across policies when varying the patient/provider ratio, and find that the greedy and gradient descent policies perform best, while the group-based and pairwise policies lag.
We compare the match rate across policies from Example 5 and find that the gradient descent and greedy policies achieve the highest match rate for large patient/provider ratios (see Figure 9). As shown in Section 3.1, the greedy policy achieves the highest match rate for the uniform choice model, so it is unsurprising that it achieves the highest match rate. Beyond that, the high performance of gradient descent matches the trends with match quality in Example 5.
Impact of $p$ and $\theta$
Figure 10: For $p\leq 0.5$, we find that all four policies achieve the same match rate, while for $p\geq 0.75$, we find that gradient descent achieves a slightly lower match rate compared to the greedy, pairwise, and group-based policies.
We extend Example 6 to compare the match rates across policies when varying $p$ and $\theta$ in Figure 10. All non-random policies have similar match rates for $p\leq 0.5$, implying that match quality is better at differentiating policies than match rate. For larger $p$, the gradient descent policy performs slightly worse in match rate, while the greedy, pairwise, and group-based policies achieve the same match rate. As in Example 6, we find that the best-performing policies for match quality also tend to exhibit the highest match rate (in this case, the group-based policy is optimal for both).
Misspecified Choice Models
Figure 11: We evaluate the performance of policies when the observed match probability, $\hat{p}$, differs from the true match probability, $p$. We find that the observed match probability has little impact on the relative ordering of policies, although the group-based policy improves as $p$ and $\hat{p}$ get closer.
Our experiments in Section 5 consider situations with known values of $p$, but real-world scenarios frequently involve misspecified choice models. To model this, we run policies with some observed match probability $\hat{p}$ while letting $p$ be the match probability. This setup allows us to simulate real-world scenarios where match probability is unknown. We compare policies across $\hat{p}\in\{0.1,0.25,0.5,0.75,0.9\}$ and $p\in\{0.1,0.25,0.5,0.75,0.9\}$.
In Figure 11, we demonstrate that most policies perform similarly across $\hat{p}$ for fixed $p$. This trend holds because the random, greedy, and pairwise policies are unimpacted by $\hat{p}$. We find that the group-based policy gets worse as $\hat{p}$ and $p$ differ; for example, when $p=0.9$ and $\hat{p}=0.1$, the group-based policy performs worse than the pairwise policy, while when $\hat{p}=0.9$, the policies perform similarly. We note that the differences are slight even for the group-based policy, so our policies are generally robust to misspecification in match probabilities.
Appendix B Proofs
See 2.3
Proof.
Suppose that $s$ patients are offered the provider $j=1$: $\lVert\mathbf{X}\rVert_{1}=s$. In the uniform choice model, the probability that any of the $s$ patients select the single provider is $p$, so the probability of provider $j=1$ matching is $1-(1-p)^{s}$. By symmetry in the response order, the chance that each patient is selected is equal, and so the match quality is
$(1-(1-p)^{s})\frac{\sum_{i=1}^{N}\theta_{i,1}\mathbf{X}_{i,1}}{s}$ (20)
For fixed $s$, the optimal assortment selects the $s$ largest values of $\theta_{i,1}$. Next, we note that larger values of $s$ increase the match probability $(1-(1-p)^{s})$, but could decrease the average match quality for the selected patient, $\frac{\sum_{i=1}^{N}\theta_{i,1}\mathbf{X}_{i,1}}{s}$. To balance between the two factors, we can iterate through values of $s$ and compute $(1-(1-p)^{s})$ and $\frac{\sum_{i=1}^{N}\theta_{i,1}\mathbf{X}_{i,1}}{s}$. ∎
See 3.1
Proof.
We construct a problem instance where the greedy policy achieves an $\epsilon$ fraction of the optimal match quality. Let $N=M$ and construct $\theta$ as follows: let $\theta_{1,1}=1$, while $\theta_{i,1}=2\Delta$ for $i\neq 1$, where $\Delta\leq\frac{1}{2}$. Let $\theta_{i,j}=\Delta$ for all $i$ and $j\neq 1$.
Let $\pi$ be the policy so that $\pi(\theta)_{i}=\mathbf{e}_{i}$. The expected match quality for this policy is $p(\Delta(N-1)+1)$; for patients 2 to $N$, it achieves an expected match quality of $p(\Delta(N-1))$ in total, while for patient $1$ it achieves an expected match quality of $p$. Because $\pi^{*}(\theta)$ is optimal, we have
$p(\Delta(N-1)+1)=\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (21)
Next, consider the greedy policy where $\pi^{R}(\theta)=\mathbf{1}$. All patients prefer provider $1$, and because of this, all patients have an equal likelihood of matching with provider $1$. Each patient has a $\frac{1}{N}$ chance of matching with provider $1$, and therefore the expected match quality for provider $1$ is $\frac{1}{N}+\frac{N-1}{N}(2\Delta)$. For all the $N-1$ other providers, we receive a reward of $\Delta$ upon matching, and each patient matches with probability $p$; therefore, we receive a total match quality of at most $\frac{1}{N}+2\frac{N-1}{N}\Delta+p(N-1)\Delta$. Taking the ratio of the greedy and optimal policies yields the following:
$\frac{\frac{1}{N}+\frac{2\Delta(N-1)}{N}+p(N-1)\Delta}{p(\Delta(N-1)+1)}$
$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p(\Delta(N-1)+1)}$
$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p}$
$\leq\frac{1}{Np}+\frac{2\Delta}{p}+N\Delta$
Finally, letting $N=\frac{3}{\epsilon p}$ and $\Delta=\min(\frac{\epsilon}{3N},\frac{p\epsilon}{6})$ yields that
$\frac{1}{Np}+2\frac{\Delta}{p}+N\Delta\leq\frac{\epsilon}{3}+\frac{\epsilon}{3}+\frac{\epsilon}{3}=\epsilon$ (22)
Therefore, for any choice of $\epsilon$ and $p$, there exists a choice of $\theta$ so the greedy policy is at most an $\epsilon$ approximation. ∎
See 9
Proof.
The pairwise policy constructs an assortment so each patient is only offered one provider, and no provider is offered to more than one patient. Under this scenario, each patient offered a non-zero assortment matches with probability $p$; that is, if $\pi^{P}(\theta)_{i,j}=1$, then the expected match quality is $p\theta_{i,j}$. The match quality is then $\frac{p}{N}\sum_{i=1}^{N}\pi^{p}(\theta)_{i,j}\theta_{i,j}$
Next, we upper bound the performance of the optimal policy. The optimal policy achieves a match quality of
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ (23)
For any order $\sigma$, we have that no provider can be selected by two patients, and each patient can select at most one provider. That is, for any ordering, we can bound the match quality as
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]\leq\max\limits_{\mathbf{X},\sum_{i=1}^{N}X_{i,j}\leq 1,\sum_{j=1}^{M}X_{i,j}\leq 1}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}X_{i,j}\theta_{i,j}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (24)
The second inequality comes from the definition of $\pi^{P}(\theta)$. Therefore:
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]\leq\frac{1}{N}\sum_{i=1}^{N}\pi^{p}(\theta)_{i,j}\theta_{i,j}=p\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{p}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ (25)
See 2
Proof.
We will first prove the reverse direction; that if $G$ is a complete digraph, then $\pi$ and $\pi^{P}$ have the same match rate. Consider patient $i$ in a component that is a connected component of size $k$. This means that $\sum_{j}\pi(\theta)_{i,j}=1$. We will prove the set of providers unmatched and offered to patient $i$ is non-empty; that is $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}>0$ for any ordering $\sigma$ and $\mathbf{y}^{(t)}$ with $\sigma_{t}=i$. $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}=k$ when $\mathbf{y}^{(t)}=\mathbf{1}$, and decreases by one each time a provider from $\pi(\theta)_{\sigma_{t}}$ is selected. For any $j$ with $\pi(\theta)_{i,j}=1$, there exists $k-1$ other $i^{\prime}$ with $\pi(\theta)_{i^{\prime},j}=1$ because all patients within the complete graph have provider $j$ on their assortment. Therefore, at most $k-1$ other patients can select providers from $\pi(\theta)_{i}$, and therefore $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}$ can decrease by at most $k-1$ before $\sigma_{t}=i$. Therefore, $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}\geq k-(k-1)=1>0$. Moreover, because each patient has a non-empty assortment, in expectation, $Np$ patients match, and so both $\pi$ and $\pi^{P}$ achieve $Np$ matches.
Next, we will prove the forward direction: if the match rates are equal between $\pi$ and $\pi^{P}$, then $G$ consists of components that are complete digraphs. We note that $\pi^{P}$ achieves a match rate of $Np$ because each patient is offered a non-empty assortment of size one. Therefore, the only way for $\pi$ to achieve the same match rate is to offer each patient a non-empty assortment. To show this, we first consider some node $u$ in the graph, corresponding to a patient. Suppose that patient $i$ has an assortment of size $k$, indicating that there exist $k$ edges from $i$ to some node. Suppose that there exists a node, $i^{\prime}$ such that $(i^{\prime},i)$ is an edge, but $(i,i^{\prime})$ is not. Then consider the ordering that places $i$ after its neighbors and $i^{\prime}$. In such an ordering, suppose $i^{\prime}$ selects $v_{i}$ and each neighbor of $i$, $u$, select $v_{u}$. This results in neither $i$ nor its neighbors being available when $i$ must select a patient, leaving an empty assortment. Therefore, if patient $i$ has $v_{i^{\prime}}$ on its assortment, then $i^{\prime}$ must have $v_{i}$ on its assortment for the policy $\pi$.
Next, we consider the scenario where there exists a node $w$ such that $w$ is a two-hop neighbor of $i$ but not an immediate neighbor of $i$. Suppose $w$ is adjacent to $i^{\prime}$, so that $i^{\prime}$ is on $w$’s assortment. Order the patients such that $w$ comes first, then $i$’s neighbors, then $i$. Let $w$ select $i^{\prime}$, let $i^{\prime}$ select $i$, and let all of $i$’s other neighbors select themselves. This results in an empty assortment for $i$; therefore, when all assortments are nonempty, there must not exist any two-hop neighbors that are not also one-hop neighbors. In other words, every node in $G$ only has one-hop neighbors. Therefore, in any component, all neighbors are connected, and so $G$ consists of components that are complete digraphs. ∎
See 3.1
Proof.
We first compute the expected match quality for the pairwise policy. Each of the $M$ matched patients has a $p$ chance of selecting their assigned provider, and $\theta_{i,j}\leq 1$; therefore, $\mathbb{E}_{\sigma,\theta}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq pM$.
[…]
$\mathbb{E}[\theta_{i,j}|\text{j is top avail. for i}]=\frac{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}\mathbb{E}[(\theta_{i})_{(j)}]}{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}}$ (26)
By Chebyshev’s sum inequality, we have that
$\frac{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}\mathbb{E}[(\theta_{i})_{(j)}]}{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}}\geq\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}[(\theta_{i})_{(j)}]=\frac{1}{2}$ (27)
Therefore, summing across all providers gives the ratio as
$\frac{1-(1-p)^{N/M}}{2p}$ (28)
See 2
Proof.
We compute a matrix, $\hat{\mathbf{Y}}$, such that $Y_{i,j}$ corresponds to the probability that patient $i$ is matched with $j$. Then $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]=\hat{\mathbf{Y}}\cdot\theta$.
We then need to prove that
$(p*g(f(\mathbf{X}))-\hat{\mathbf{Y}})\cdot\theta\leq 0$ (29)
To prove this, first let $\ell_{u_{i,j}}=(p*g(f(\mathbf{X}))_{i,j}-\hat{Y}_{i,j})$.
We will first bound $\sum_{j=1}^{M}\ell_{u_{i,j}}$ for a fixed $i$. Let $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,2}}\ldots\theta_{i,u_{i,M}}$ for some coefficients $u$. For any $m$, $\sum_{j=1}^{m}\ell_{u_{i,j}}=\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}-\hat{\mathbf{Y}}_{i,u_{i,j}}$. We can next compute $\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}$. Recall from Section 3.4 that we can explicitly compute $g(f(\mathbf{X}))_{i,j}$ through the function $h(n)=\frac{1}{N}\sum_{k=1}^{N}(1-p\frac{n-1}{N-1})^{k-1}$. Then by the of definition $g(f(\mathbf{X})_{i,j})$, we have $p\sum_{j=1}^{m}g(f(\mathbf{X}))_{i,j}=p(\sum_{j=1}^{m}X_{i,u_{i,j}}h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,l}}\rVert_{1})))$.
Next, we note that $\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ represents the probability that any of the top-$m$ options are available and selected. For any provider $j$, the probability that $j$ is available is at least $h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (see Section 3.4). Therefore,
$\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ (30)
[…]
$\geq p\sum_{j=1}^{m}X_{i,u_{i,j}}h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (32)
Combining our statements for $\hat{\mathbf{Y}}$ and $p*g(f(\mathbf{X}))$ gives that $\sum_{j=1}^{m}X_{i,u_{i,j}}\hat{\mathbf{Y}}_{i,u_{i,j}}\geq\sum_{j=1}^{m}h(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})\prod_{l=1}^{m-1}(1-(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})=\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}$. Therefore, we have that $\sum_{j=1}^{m}\ell_{u_{i,j}}\leq 0$.
Next, we will show that $(p*g(f(\mathbf{X}))_{i}-\hat{\mathbf{Y}}_{i})\cdot\theta_{i}=\sum_{j=1}^{M}\ell_{u_{i,j}}\theta_{i,j}\leq 0$ for all $i$. We will prove this inductively; first note that $l_{i,u_{i,1}}\theta_{i,u_{i,1}}\leq 0$ Next, assume that $\sum_{j=1}^{M-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$, then:
$\sum_{j=1}^{m}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$ (33)
$=\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m}}$ (34)
$\leq\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m-1}}$ (35)
$\leq(\sum_{j=1}^{m-1}\ell_{u_{i,j}})(\sum_{j=1}^{m-1}\theta_{i,u_{i,j}}-\theta_{i,u_{i,m-1}})\leq 0$ (36)
Therefore, $p*(g(f(\mathbf{X}))_{i}\cdot\theta_{i})\leq 0$ for all $i$, and so summing across all $i$ gives $p*(g(f(\mathbf{X}))\cdot\theta)\leq 0$
See 2
Proof.
Let $\mathbf{X}=\pi^{P}(\theta)$ for some given $\theta$. Then because $\pi^{P}(\theta)$ is computed by solving a 1-1 bipartite matching problem, we know that $\sum_{i=1}^{N}X_{i,j}\leq 1$ for all $j$ and $\sum_{j=1}^{M}X_{i,j}\leq 1$ for all $i$. Next, we note that $0\leq\lVert\mathbf{X}_{*,j}\rVert\leq 1$, and moreover, if $\mathbf{X}_{i,j}=1$, then $\lVert\mathbf{X}_{*,j}\rVert=1$. Because $h(1)=1$, we have that $f(\mathbf{X})_{i,j}=\mathbf{X}_{i,j}$. Finally, because each row $i$ of $f(\mathbf{X})_{i,j}$ has at most one $j$ with $\mathbf{X}_{i,j}=1$, and because $\bar{\mathbf{X}}_{i,u_{k}}=f(\mathbf{X})_{i,u_{k}}p\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{k}})$, we have that $g(f(\mathbf{X}))=f(\mathbf{X})=p\mathbf{X}$. Therefore, $g(f(\mathbf{X}))\cdot\theta$ is exactly p times the value of the bipartite match. Finally,
$\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]=p\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}=p(f(g(\mathbf{X}))\cdot\theta).$ (37)
Therefore, the expected match quality for the pairwise policy $\pi^{P}(\theta)$ is the same as $p(\mathbf{X}\cdot\theta)$. ∎
See 4.1
Proof.
We first note that $G$ is acyclic; if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ and $\theta_{i^{\prime},v_{i^{\prime}}}\leq\theta_{i^{\prime},v_{i}}$, then swapping $v_{i}$ and $v_{i^{\prime}}$ improves the pairwise policy $\pi^{P}(\theta)$. We note that $\pi^{P}(\theta)$ is the optimal solution for $\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)\theta_{i,j}$ by definition, and so such a swap would violate this property, implying that $G$ is acyclic.
Next, suppose that $\sigma^{*}$ is not a reverse topological ordering. Then there exists nodes $i$ and $i^{\prime}$, so that $i$ comes before $i^{\prime}$ in $\sigma^{*}$ and there exists an edge from $i$ to $i^{\prime}$. Consider the following cases:
  1. 1.
    Case 1: There exists no node $w$, between $i^{\prime}$ and $i$ in $\sigma$ so that $w$ and $i^{\prime}$ share an edge (in either direction). Then $i$ and $i^{\prime}$ can be placed in sequence ($i$ comes immideatly before $i^{\prime}$) without impacting the expected reward. This is because all nodes $w$ between $i$ and $i^{\prime}$ have no impact on the preferences of $i^{\prime}$ because $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{w}}$. Moreover, swapping $i$ and $i^{\prime}$ can only increase the expected match quality; placing $i$ before or after $i^{\prime}$ has no impact on the match quality for patient $i^{\prime}$ (as $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{i}}$), but placing $i$ after $i^{\prime}$ can increase the expected match quality if $v_{i}$ goes unmatched.
  2. 2.
    Case 2: Suppose there exists a node $w$, such that there is an edge from $w$ to $i^{\prime}$ Then the pair $(w,i^{\prime})$ violates the reverse topological order, and so we can recurse on this. Note that this is a smaller length pair of nodes within the ordering $\sigma^{*}$. Similarly if there is an edge from $i$ to $w$, then $(i,w)$ violates the reverse topological order ,and so we can recurse on this
  3. 3.
    Case 3: Suppose that there exists a node $w$ such that there are edges from $i^{\prime}$ to $w$ and $w$ to $i$. Then this creates a cycle, breaking the acyclic property mentioned earlier.
  4. 4.
    Case 4: Suppose that there exists a node $w$ such that there is an edge from $i^{\prime}$ to $w$. There exists no path from $i^{\prime}$ to $i$ due to the acyclic nature of the graph. Additionally, $w$ does not have an edge to $i$ or any of the ancestors of $i$. Move $i^{\prime}$ and all of its descendants before $i$ in the graph; there are no edges from the descendants of $i^{\prime}$ to $i$. Therefore, this flips the order of $i$ and $i^{\prime}$ without impacting any of the descendants; because only $i$ and $i^{\prime}$ are impacted by this change (as the descendants of $i^{\prime}$ are not impacted by $i$), such a change can only increase the expected match quality, as $i$ can potentially match with $v_{i^{\prime}}$.
  5. 5.
    Case 5: Suppose there exists a node $w$ such that there exists an edge from $w$ to $i$. Due to the acyclic nature of the graph, there are no edges from $i^{\prime}$ (or any of its ancestors) to $w$ or any of the ancestors of $i$. Now consider moving $i$ and its ancestors after $i^{\prime}$. Because there are no edges from $i^{\prime}$ to the ancestors of $i$, the result of the assortment from the ancestors of $i$ has no impact on $i^{\prime}$. Therefore, while patient $i$ can now potentially match with provider $v_{i^{\prime}}$, none of the ancestors of $i$ are negatively impacted, improving aggregate match quality.
Our cases cover every scenario for nodes between $i$ and $i^{\prime}$, and in all cases, we can make a minor change to the order $\sigma^{*}$ to bring it closer to a reverse topological ordering while improving match quality or recurse upon a subset of the ordering. ∎
See 4.1
Proof.
Let $\sigma^{*}$ be the optimal ordering and let $\sigma$ be any ordering from batches $b_{1},b_{2},\ldots,b_{L}$. Then we will prove that for any node $i$, the order of descendants of $i$ will be the same in $\sigma^{*}$ and $\sigma$.
To prove this, consider a node $i$. Let the descendants of node $i$ in the optimal ordering be $d_{1},d_{2},\ldots,d_{i}$. We note that in $\sigma$, each $d_{i^{\prime}}$ is in a separate partition by properties of $b_{k}$. Additionally, because $b_{1},b_{2},\ldots,b_{L}$ is a partition of $\sigma^{*}$, the order of $d_{1},d_{2},\ldots,d_{i}$ is maintained. Therefore, for any node, the same set of ancestors is maintained through the partition $b_{k}$.
Next, we will show that the descendants of patient $i$ dictates what patient $i$ selects in $\pi^{A}$. Let $Z_{1},Z_{2},\ldots,Z_{N}$ be a set of Bernoulli random variables, each of which is 1 with probability $p$. If $Z_{t}=1$ then patient $\sigma_{t}$ will select the highest available provider in their assortment, and if $Z_{t}=0$, then patient $\sigma_{t}$ will select no provider. Additionally, let $V_{1},V_{2},\ldots,V_{N}$ be a set of random variables, where $Z_{t}$ denotes the provider selected by patient $\sigma_{t}$ ($V_{t}=0$ if $Z_{t}=0$). Note that $V_{1}$ is only a function of $Z_{1}$; if $Z_{1}=0$, then $V_{1}=0$, and otherwise, $V_{1}=\operatorname*{arg\,max}_{j}\theta_{\sigma_{1},j}$. Next, let $D_{\sigma_{t}}\subseteq[N]$ be the set of descendants for patient $\sigma_{t}$; that is, $d\in D_{\sigma_{t}}$ means there exists a path from $\sigma_{t}$ to $d$ in $G$. Suppose that $V_{\sigma_{t^{\prime}}}=f(\{Z_{d}\}_{d\in D_{\sigma_{t^{\prime}}}})$ for all $t^{\prime}\leq t$. Then patient $\sigma_{t+1}$ will only select some $j$ so $\theta_{\sigma_{t+1},j}\geq\theta_{\sigma_{t+1},v_{\sigma_{t+1}}}$. By the construction of the graph $G$, this corresponds to edges connected to patient $\sigma_{t+1}$, and so $V_{t+1}=g(\{V_{d}\}_{d\in D_{\sigma_{t+1}})}$. Next, note that $V_{d}=f(\{Z_{d^{\prime}}\}_{d^{\prime}\in D_{d}})$. Therefore, we can rewrite $V_{t+1}$ as
$V_{t+1}=g(\{f(\{Z_{d^{\prime}}\}_{d^{\prime}\in D_{d}})\}_{d\in D_{t+1})}.$ (38)
We finally collect like terms and note that the set of descendants of the children of $\sigma_{t+1}$ is the set of descendants of $\sigma_{t+1}$. Therefore, we rewrite $V_{t+1}$ as
$V_{t+1}=g(\{Z_{d}\}_{d\in D_{t+1}})$ (39)
Therefore, $V_{t+1}$ only depends on the descendants of $t+1$. Therefore, if the ordering of descendants is fixed, and $Z_{1},Z_{2},\ldots,Z_{N}$ are decided a priori, then $V_{\sigma_{t}}$ is also fixed. The ordering of descendants is the same between $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ and $\sigma^{*}$, so both achieve the same match quality. ∎

### v2

@@
 Abstract
-Rising provider turnover forces healthcare administrators to frequently rematch patients to available providers, which can be cumbersome and labor-intensive.
-To reduce the burden of rematching, we study algorithms for matching patients and providers through assortment optimization.
-We develop a patient-provider matching model in which we simultaneously offer each patient a menu of providers, and patients subsequently respond and select providers.
-By offering assortments upfront, administrators can balance logistical ease and patient autonomy.
-We study policies for assortment optimization and characterize their performance under different problem settings.
-We demonstrate that the selection of assortment policy is highly dependent on problem specifics and, in particular, on a patient’s willingness to match and the ratio between patients and providers .
-On real-world data, we show that our best policy can improve match quality by 13% over a greedy solution by tailoring assortment sizes based on patient characteristics.
-We conclude with recommendations for running a real-world patient-provider matching system inspired by our results.
+Rising provider turnover results in frequently needing to rematch patients with available providers.
+However, the rematching process is cumbersome for both patients and health systems, resulting in labor-intensive and ad hoc reassignments.
+We propose a novel patient-provider matching approach to address this issue by offering patients limited provider menus.
+The goal is to maximize match quality across the system while preserving patient choice.
+We frame this as a novel variant of assortment optimization, where patient-specific provider menus are offered upfront, and patients respond in a random sequence to make their selections.
+This hybrid offline-online setting is understudied in previous literature and captures system dynamics across various domains.
+We first demonstrate that a greedy baseline policy–which offers all providers to all patients–can maximize the match rate but lead to low-quality matches.
+Based on this, we construct a set of policies and demonstrate that the best policy depends on problem specifics, such as a patient’s willingness to match and the ratio of patients to providers.
+On real-world data, our proposed policy improves average match quality by 13% over a greedy solution by tailoring assortments based on patient characteristics.
+Our analysis reveals a tradeoff between menu size and system-wide match quality, highlighting the value of balancing patient choice with centralized planning.
 1 Introduction
-Primary care providers are essential to the healthcare ecosystem because they are the first point of contact for many patients (Pearson and Raeke, 2000; Wu et al., 2022).
-Patients rely on primary care providers for routine checkups and referrals to specialists.
-Moreover, care continuity can instill trust and improve medication uptake rates and patient health (Wu et al., 2022).
-Unfortunately, high provider turnover rates frequently lead to patients without an assigned provider (Reddy et al., 2015).
-Provider turnover disrupts patient care and leads to worse care (Reddy et al., 2015).
-In principle, healthcare administrators reassign unmatched patients to other providers; however, in practice, the process takes months due to provider scarcity and the logistical burden of rematching and coordinating patient matches (Hedden et al., 2021).
-While many patients find their new provider quickly, others have to wait years to find a new provider due to large numbers of patients, high turnover rates, and provider scarcity (Hedden et al., 2021; Shanafelt et al., 2012).
-Algorithms that automatically match patients and providers can reduce logistical hassle but require balancing patient autonomy and system-wide utility.
-For example, while automatically assigning each patient to a provider would decrease wait times, it also reduces patient autonomy because patients cannot select their provider (Entwistle et al., 2010; Gaynor et al., 2016).
-In contrast, offering patients full autonomy could lead to suboptimal matches; for example, patients who rematch quickly might prevent better matches for late-responding patients.
-In addition, granting patients full independence can be overwhelming and delay decisions (Bate and Robert, 2005).
-We propose to match patients and providers through an assortment framework (Shi, 2016; Rios and Torrico, 2023; Davis et al., 2013).
-Assortments are menus of options given to customers who select an option from such a menu (Bertsimas and Mišic, 2015).
-Assortments are used in e-commerce settings where platforms give customers a suggested list of products and have them select from this list (Tian et al., 2024).
-In an assortment setting, each patient receives a menu of providers and selects a provider from such a menu.
-These curated offer sets allow administrators to balance patient autonomy and system-wide match utility.
-We devise policies for offering assortments in a patient-provider matching scenario.
-To do so, we first develop a model of patient-provider matching then construct a set of policies for offering assortments.
-We characterize how policy performance depends on problem characteristics such as patient-to-provider ratio, and we empirically detail the regimes under which each of our policies performs best.
-We then evaluate our policies in a real-world scenario constructed on top of a Medicare dataset to assess the performance of our policies in a real-world example.
-[^1] We include all code at https://github.com/naveenr414/patient-provider
-Our key question is the following: How should healthcare administrators design provider menus for patients to optimize system-wide match rates and quality?
+Primary care providers (PCPs) are essential to the healthcare ecosystem because they are the first point of contact for many patients (Pearson and Raeke, 2000; Wu et al., 2022).
+Patients rely on PCPs for routine checkups and referrals to specialists.
+PCPs serve as a regular touchpoint for patients and provide care continuity, which has been shown to instill trust and improve medication uptake rates and patient health (Wu et al., 2022).
+Furthermore, high-quality patient-provider matches compound the benefit of primary care, where the notion of a high-quality match is heterogeneous across patients and can include factors such as geographical proximity, provider specialties, and concordance along language, race, and gender (Greenwood et al., 2020, 2018; Manson, 1988).
+As one such example, Alsan et al.
+(2019) found that access to race-concordant doctors could reduce cardiovascular disease by 19% (Alsan et al., 2019).
+Unfortunately, high provider turnover rates frequently lead patients to lose their PCP, which disrupts patient care and can lead to worse health outcomes (Reddy et al., 2015).
+In principle, healthcare administrators reassign unmatched patients to other providers.
+However, in practice, the process takes months due to provider scarcity and the administrative burden of patient coordination (Hedden et al., 2021).
+As an alternative to manual reassignments, algorithmically matching patients and providers can reduce logistical hassle.
+However, this approach requires carefully balancing patient autonomy and system-wide utility.
+While automatically assigning each patient to a provider would decrease wait times, it also eliminates patient choice (Entwistle et al., 2010; Gaynor et al., 2016).
+In contrast, offering patients full autonomy to choose between all available providers could lead to suboptimal matches when considering system-wide match quality; for example, patients who choose quickly might prevent better matches for late-responding patients.
+In addition, granting patients full independence can be overwhelming and lead to decision delay (Bate and Robert, 2005).
+In this work, we formulate the patient-provider matching problem as a variant of assortment optimization (Shi, 2016; Rios and Torrico, 2023; Davis et al., 2013); we illustrate this problem in Figure 1.
+Assortments are menus of options given to customers, who then make selections from this menu (Bertsimas and Mišic, 2015).
+This framework is commonly seen on e-commerce platforms (Tian et al., 2024).
+By offering limited provider menus, patients have the autonomy to select their preferred provider, while healthcare administrators can curate these assortments to maximize system-wide metrics.
+Our work studies a unique variant of assortment optimization with offline assortment and online response: assortments are created and offered simultaneously, and patients then respond sequentially in a random order to make provider selections.
+This variant is especially interesting because it captures the logistical complexities of the patient-provider matching problem.
+Dynamically varying assortments online are logistically taxing for healthcare administrators, yet it is difficult to control the patient response order, necessitating a combination of offline and online modeling.
+Moreover, this setting generalizes beyond patient-provider matching; similar situations can be found in other two-sided matching markets such as food delivery, where companies may propose a set of order options (menus) to drivers who respond in an online fashion, and where there is a tradeoff between driver autonomy and system utility (i.e., cost).
+Our contributions answer this key question: How should healthcare administrators design provider menus for patients to optimize system-wide match rates and match quality?
+Our results show that limiting assortment sizes and offering carefully tailored menus enables a tradeoff between match rate and match quality, underscoring the balance between patient autonomy and centralized planning.
 1.1 Contributions
-We develop a model of patient-provider matching using assortment offerings and analyze its impact on match rate and match quality.
-We characterize well-performing policies when varying properties such as the patient/provider ratio.
-We then empirically analyze these policies in a synthetic and real-world dataset built using a Medicare dataset.
-We use the insights from these studies to propose recommendations for designing patient-provider matching systems in practice.
-Developing a Model of Patient-Provider Matching
-We develop a model of patient-provider matching where administrators offer assortments upfront, then patients sequentially respond in a random order and select providers one by one.
-We do so because it captures two critical elements: 1) patients have autonomy in selecting providers and 2) administrators have some control over the matching process by varying assortments offered to patients.
-We model patient decisions through a choice model that captures how the set of providers offered impacts the final provider chosen.
-The selection of choice model is key to the fidelity of our model, so we analyze various choice models, including the uniform choice model and the multinomial logit choice model.
+Overall, we make three contributions to the patient-provider matching and assortment optimization literature: we i) model patient-provider matching using a hybrid offline-online variant of assortment optimization, ii) characterize the performance of assortment policies theoretically and empirically, and detail how the best-performing policy depends on problem characteristics such as patient match probability and patient/provider ratio, and iii) demonstrate that carefully tailoring assortment menus allows us to tradeoff between match rate and match quality.
+Patient-Provider Matching as Assortment Optimization
+We develop a model of patient-provider matching where administrators offer assortments upfront, and then patients respond sequentially in a random order.
+We do so to capture two critical elements: 1) patients have autonomy in selecting providers, and 2) the system has some control over the matching process in selecting assortments for each patient.
+Additionally, such a model can be broadly applicable to two-sided marketplaces where one side (users) selects online from an assortment, yet dynamic variation of the assortment is infeasible.
+We model patient decisions through a choice model that captures how the set of offered providers impacts the resultant matches.
+The selection of the choice model is key to the fidelity of our model, so we analyze various choice models, including the uniform choice model and the multinomial logit choice model.
 Finally, we capture patient heterogeneity through a match quality matrix, which captures differences in match quality between pairs of patients and providers.
-Match quality can encompass geographical proximity, provider specialties, and concordance along language, race, and gender (Greenwood et al., 2020, 2018; Manson, 1988).
-Patients tend to match with higher quality providers, though the exact match probabilities are dictated by their choice model.
-Constructing Assortment Policies
-We characterize the optimal policy in the single-provider scenario and use these insights to construct a set of policies for assortment optimization.
-We first construct a baseline policy, greedy, which offers every provider to every patient.
-We demonstrate that the greedy policy can do arbitrarily poorly, so we build upon this by developing three new policies that adaptively vary assortment sizes: pairwise, group-based, and gradient descent.
-The pairwise policy offers each patient at most one provider, and we demonstrate an instance-dependent approximation guarantee for this policy.
-We then develop the group-based policy by augmenting the pairwise policy so we offer assortments to groups of patients.
-We demonstrate that such group-based maintains the match rate of the pairwise policy while potentially improving the match quality.
-For the gradient descent policy, we construct an objective function that provably lower bounds the match quality, and we use gradient descent to optimize this quantity.
-Characterizing Policy Performance
-We analyze the performance of our policies in a synthetic setting and empirically characterize the conditions under which each policy performs best.
-We demonstrate that the best-performing policy depends on two factors: the match probability and the ratio of patients to providers.
-When patients outnumber providers, as is common in many real-world health systems, the gradient descent policy performs best because the objective function is closer to the lower bound.
-When patient and provider counts are balanced, we show that the greedy policy achieves a high match quality when patients have low match probabilities, while the group-based policy achieves a high match quality when patients have high match probabilities.
-We evaluate policies in a real-world scenario constructed via Medicare data (Medicare & Medicaid, 2025) and we show here that the gradient descent policy performs best.
-We extend this scenario to real-world considerations such as fairness and cognitive overload.
-Contribution Summary
-Overall, we make three contributions to the patient-provider matching and assortment optimization literature: we i) develop a model of patient-provider matching under assortments, ii) characterize the performance of assortment policies, and detail how the best-performing policy depends on problem characteristics such as patient match probability and patient/provider ratio, and iii) make recommendations for the implementation of a real-world patient-provider matching system based on the results of our theoretical and empirical analysis.
-Figure 1: A) In this work, we match patients with providers B) by offering a set of assortments to patients, where each assortment consists of a set of providers.
-C) Patients then respond sequentially and either select a provider from their assortments or abstain.
+Characterizing Assortment Policies
+We study various policies for assortment optimization and characterize their performance theoretically and empirically.
+We quantify policy performance through two metrics: match rate and match quality.
+We first propose a baseline greedy policy that offers all providers to all patients, then demonstrate that this policy can perform arbitrarily poorly for match quality.
+We then build upon this by developing new policies that adaptively vary assortment sizes.
+We demonstrate that tailored assortment offerings can improve match quality compared to baseline greedy solutions.
+We empirically characterize the best policy as a function of model parameters such as the ratio of patients to providers and the underlying model assumptions.
+[^1] We include all code at https://github.com/naveenr414/patient-provider/
+Real-World Matching System Insights
+We apply our policies in a semi-synthetic simulation based on a real-world healthcare system in Connecticut.
+We demonstrate that the best-performing policy here outperforms the greedy baseline by 13% for match quality.
+We extend this to consider metrics beyond match quality and match rate, such as fairness and regret.
+These findings yield actionable recommendations and highlight the tensions between various metrics when designing real-world patient-provider matching systems.
+Our key takeaway is that the best set of assortments offered depends on problem specifics, including the ratio of patients to providers and the willingness of patients to match.
+As patients are more willing to match, assortments should be smaller and more tailored to maximize match quality.
+Figure 1: Overview of patient-provider matching model.
+A) We match patients with providers B) by offering a set of assortments to patients, where each assortment consists of a set of providers.
+C) Patients then respond sequentially in a random order and either select a provider from their assortments or abstain.
 D) This results in matches between patients and providers.
@@
 Moreover, such issues are exacerbated recently due to high rates of provider burnout (Shanafelt et al., 2012) and the Covid-19 pandemic (Shanafelt et al., 2022).
-Once a provider drops out, it is difficult for patients to find a new one; while 54% of patients find their new provider within 12 months, 6% fail to find a new provider even after 36 months (Hedden et al., 2021).
-High levels of provider turnover can worsen primary care because patients lack access to routine checkups or reminders to take medication (Reddy et al., 2015; Levine et al., 2019; Nguyen et al., 2020; Alsan et al., 2019).
-While provider turnover impacts patient satisfaction, quantifying “match quality” is hard.
+When a provider leaves, it is difficult for patients to find a new one; only 54% of patients find their new provider within one year, and 6% fail to find a new provider even after three years (Hedden et al., 2021).
+High levels of provider turnover can worsen primary care because patients lack access to routine checkups which can reduce accountability, leading to consequences such as lower medication adherence (Reddy et al., 2015; Levine et al., 2019; Nguyen et al., 2020; Alsan et al., 2019).
+Provider turnover itself impacts patient satisfaction and health outcomes, but the quality of the new match is also consequential in both dimensions.
+Nevertheless, it is difficult to quantify what constitutes a “high-quality” match.
 Patients care mainly about geographic proximity (Dunlea and Lenert, 2015), with concordance along race (Greenwood et al., 2020), gender (Greenwood et al., 2018), and language (Manson, 1988) playing a secondary role.
 Communication style alignment can improve match quality (Street et al., 2008).
-These factors build trust and lead patients to more truthfully report symptoms (Persky et al., 2013).
-Algorithms for Patient-Provider Matching Existing work in patient-provider matching uses either a one-shot matching framework or a genetic programming-based approach.
-Within the one-shot framework, linear programs can manage provider panel sizes (Harrington et al., 2021), the deferred acceptance framework can perform two-stage matching (Chen et al., 2020), and scheduling algorithms are used to model optimal rates of on-demand scheduling (Qu et al., 2007).
-A second approach to patient-provider matching builds matches through genetic programming.
-Genetic programming algorithms learn matches over time through a fitness function.
+These factors build trust which translates to better care, such as leading patients to more truthfully report symptoms (Persky et al., 2013).
+Algorithms for Patient-Provider Matching Existing work in patient-provider matching typically adopts either a one-shot matching framework or a genetic programming-based approach.
+Within the one-shot framework, linear programs are used to manage provider panel sizes (Harrington et al., 2021); deferred acceptance frameworks enable two-stage matching (Chen et al., 2020); and scheduling algorithms model optimal rates of on-demand scheduling (Qu et al., 2007).
+An alternative approach to patient-provider matching uses genetic programming, which learns matches over time through a fitness function.
 Within this framework, one line of work uses genetic programming to find fair matches between patients and providers (Chen et al., 2020), while another approach uses genetic programming to balance workloads between providers (Zhu et al., 2023).
-We build on these lines of work by incorporating patient autonomy into the process, which motivates the need for an assortment-based approach.
-Assortment Optimization Our work analyzes patient-provider matching through assortment optimization, where retailers offer a set of menus to customers.
-In assortment optimization, retailers construct assortments and customers make decisions based on offered options (Bertsimas and Mišic, 2015).
-Retailers first construct models of customer decisions, such as the multinomial logit model (Davis et al., 2013) and nested logit models (Alfandari et al., 2021), both of which use logit probabilities to model customer decisions.
-Assortments are employed in domains including school choice (Shi, 2016), dating markets (Rios and Torrico, 2023), and e-commerce (Tian et al., 2024).
-Assortment optimization algorithms come in offline and online variants.
-In offline assortment optimization, assortments are offered one-shot and customers make decisions simultaneously and independently (Bertsimas and Mišic, 2015).
-In online assortment optimization, customers arrive sequentially, and options are offered for each (Aouad and Saban, 2023).
+In contrast to these approaches, we incorporate patient autonomy into the process, which motivates our assortment-based approach.
+Assortment Optimization Our work analyzes patient-provider matching through assortment optimization.
+In this setting, retailers construct assortments for customers, and customers make decisions based on offered options (Bertsimas and Mišic, 2015).
+Retailers use models to capture customer choice behavior, such as the multinomial logit (Davis et al., 2013) and nested logit (Alfandari et al., 2021), and generate product menus accordingly.
+Assortments are ubiquitous in e-commerce settings (Tian et al., 2024) but also employed in other domains such as school choice (Shi, 2016) and dating markets (Rios and Torrico, 2023).
+Both offline and online variants have been studied.
+In offline assortment optimization, assortments are offered one-shot and customers make decisions simultaneously and independently (Bertsimas and Mišic, 2015), while in online assortment optimization, customers arrive sequentially, and options are dynamically offered for each (Aouad and Saban, 2023).
+Other variants of assortment optimization include settings where system parameters are unknown (Rusmevichientong and Topaloglu, 2012), are dynamically learned (Caro et al., 2014), or have constraints (Feldman and Topaloglu, 2018).
 Our setting resides between offline and online, as we offer assortments offline, but patients respond online.
-Matching under Random Ordering Patient-provider matching is challenging from an optimization perspective because patients make decisions in a random order.
-Randomized response orders are found in variants of online selection problems such as secretary problems (Freeman, 1983) and online learning (Orabona, 2019).
-One key difference is the lack of flexibility in our scenario; changing assortments online for the patient-provider matching setting would involve considerable logistical hassle, so we offer assortments in a one-shot manner.
 2 Problem Formulation
 2.1 Formal Model of Patient-Provider Matching
-We focus on matching $N$ patients to $M$ providers.
+We introduce the problem of matching $N$ patients to $M$ providers.
 We present each patient $i$ with an assortment $\mathbf{X}_{i}\in\{0,1\}^{M}$ upfront.
@@
 We assume patient response times are i.i.d.
-and exponential; prior work demonstrates this can model real-world wait times (Newell, 2013).
-Under this assumption, $\sigma$ is a random permutation of $[N]$.
+so $\sigma$ is a uniformly random permutation of $[N]$; prior work demonstrates this can model real-world wait times (Newell, 2013) and we relax this assumption in Section 4.3.
+Patients select providers based on match quality $\theta_{i,j}\in[0,1]$, where $\theta_{i,j}$ is analogous to a reward for patient $i$ selecting provider $j$.
+Match quality can encompass various factors, including demographic concordance (Greenwood et al., 2020, 2018; Manson, 1988), physical proximity, and patient needs.
+We assume administrators know $\theta_{i,j}$ because $\theta_{i,j}$ can be estimated through patient surveys; we further discuss this assumption in Section 2.2.
 Each patient selects a provider from their assortment according to a choice model $f_{i}:\{0,1\}^{M}\rightarrow\{0,1\}^{M}$.
-The choice model takes as input a 0-1 vector $\mathbf{z}$ and outputs a random 0-1 vector representing the provider selected by patient $i$.
+The choice model takes as input a 0-1 vector $\mathbf{z}$ denoting available and offered providers, and it outputs a random 0-1 vector representing the provider selected by patient $i$.
 Formally, let $\mathbf{y}^{(t)}$ represent the set of unselected providers at time $t$.
@@
 If $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})=\mathbf{0}$, then no provider is selected, while if $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})_{j}=1$, then provider $j$ is selected.
-We note that a patient can match with at most one provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\leq 1$
-Patients are more likely to select providers with a higher match quality, $\theta_{i,j}\in[0,1]$, where $\theta_{i,j}$ is analogous to a reward for patient $i$ when selecting provider $j$.
-Match quality can encompass various factors, including demographic concordance (Greenwood et al., 2020, 2018; Manson, 1988), physical proximity, and patient needs.
-We assume administrators know $\theta_{i,j}$.
-Prior work demonstrates providers are able to successfully identify new high-match quality providers for patients (Lohr et al., 2013), which implies that we can predict match quality from patient data.
+We note that a patient can match with at most one provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\leq 1$.
+We assume that each provider can match with at most one patient as well.
+We assume this to simplify the presentation of our paper, though we relax this assumption in Section 4.3 and show in Appendix C that this is empirically equivalent to modifying the total number of providers.
 We present examples of choice models below:
   1. 1.
-    Uniform - Patients select their most preferred provider with probability $p$ and otherwise abstain. Let $\mathbf{e}_{k}$ denote the $k$-th standard basis vector, then:
+    Uniform - Patients select their most preferred available provider with probability $p$ and otherwise abstain. Let $\mathbf{e}_{k}$ denote the $k$-th standard basis vector, then:
     $f_{i}(\mathbf{z})=\left\{\begin{array}[]{ll}\mathbf{e}_{\operatorname*{arg\,max}_{j}(\theta_{i}\odot\mathbf{z})_{j}}&\text{with prb. p}\\ \mathbf{0}&\text{otherwise}\end{array}\right.$ (1)
@@
     Threshold - Patients follow the uniform choice model but only select providers with match quality above some threshold $\alpha$.
-If match quality corresponds to geographic proximity, $\alpha$ corresponds to maximum travel distance.
+For example, if match quality corresponds to geographic proximity, $\alpha$ corresponds to maximum travel distance.
 Formally:
@@
   3. 3.
-    Multinomial Logit (MNL) - Patients select providers according to $\theta$, and have an exit option $\gamma$ which represents selecting no provider.
-Under the MNL choice model (Train, 2009), this is:
+    Multinomial Logit (MNL) - Patients select providers according to $\theta$, and have an exit option valued at $\gamma$ which represents selecting no provider.
+Under the MNL choice model (Train, 2009), $f$ is:
     $f_{i}(\mathbf{z})=\left\{\begin{array}[]{ll}\mathbf{e}_{j}&\text{with prb.
@@
 }\frac{\exp(\gamma)}{\exp(\gamma)+\sum_{j^{\prime}}z_{j^{\prime}}\exp(\theta_{i,j^{\prime}})}\end{array}\right.$ (3)
-We aim to find a policy $\pi(\theta)$ that constructs an assortment $\mathbf{X}$, given the match quality matrix $\theta$. We evaluate policies through two metrics: match quality and match rate.
+We aim to find a policy $\pi(\theta)$ that constructs an assortment $\mathbf{X}$ from a match quality matrix $\theta$. We evaluate policies through two metrics: match quality and match rate.
   1. 1.
-    Match Quality measures $\theta$ across selected patient-provider pairs.
-Formally, it aggregates $\theta_{\sigma_{t}}$ across provider selections $f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})$ for all $N$ patients (= $N$ time periods):
-    $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})\cdot\theta_{\sigma_{t}}]$ (4)
+    Match Rate (MR) is the proportion of patients who are matched.
+Formally, it computes if patient $\sigma_{t}$ matches with a provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\geq 0$:
+    $\mathrm{MR}(\mathbf{X},\theta,f)=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}(\mathbf{X}\odot\mathbf{y}^{(t)})\rVert_{1}]$ (4)
   2. 2.
-    Match Rate is the proportion of patients who are matched.
-Formally, it computes if patient $\sigma_{t}$ matches with a provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\geq 0$:
-    $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})\rVert_{1}]$ (5)
-Example 1.
-We quantify the example in Figure 1. Suppose $N=M=3$, and let $\theta$ and $\mathbf{X}$ be
+    Match Quality (MQ) measures $\theta$ across selected patient-provider pairs.
+Formally, it aggregates $\theta_{\sigma_{t}}$ across provider selections $f_{\sigma_{t}}(\mathbf{X}\odot\mathbf{y}^{(t)})$ for all $N$ patients (= $N$ time periods):
+    $\mathrm{MQ}(\mathbf{X},\theta,f)=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}(\mathbf{X}\odot\mathbf{y}^{(t)})\cdot\theta_{\sigma_{t}}]$ (5)
+We select these two metrics because they naturally capture the scale and quality of matches, though, in Section 5, we evaluate additional metrics such as fairness and regret.
+Example 2.1.
+We quantify the example in Figure 1 to demonstrate the metric calculations. Suppose $N=M=3$, and let $\theta$ and $\mathbf{X}$ be
 $\theta=\begin{bmatrix}0.5&0.6&0.9\\ 0.7&0.5&0.4\\ 0.6&0.2&0.8\end{bmatrix},\hskip 14.22636pt\mathbf{X}=\begin{bmatrix}0&0&1\\ 1&1&0\\ 1&0&1\end{bmatrix}$ (6)
 Suppose $\sigma=[2,1,3]$.
-Then one trial could be the following: patient $\sigma_{1}=2$ abstains from selecting a provider, patient $1$ selects provider $3$ and patient $3$ selects provider $1$: $f_{2}(\mathbf{y}^{(1)}\odot\mathbf{X}_{2})=\mathbf{0}$, $f_{1}(\mathbf{y}^{(2)}\odot\mathbf{X}_{1})=\mathbf{e}_{3}$, and $f_{3}(\mathbf{y}^{(3)}\odot\mathbf{X}_{3})=\mathbf{e}_{1}$.
+Then one trial could be the following: patient $\sigma_{1}=2$ abstains from selecting a provider, patient $1$ selects provider $3$, and patient $3$ selects provider $1$: $f_{2}(\mathbf{y}^{(1)}\odot\mathbf{X}_{2})=\mathbf{0}$, $f_{1}(\mathbf{y}^{(2)}\odot\mathbf{X}_{1})=\mathbf{e}_{3}$, and $f_{3}(\mathbf{y}^{(3)}\odot\mathbf{X}_{3})=\mathbf{e}_{1}$.
 Here, $\mathbf{y}^{(1)}=[1,1,1]$, $\mathbf{y}^{(2)}=[1,1,1]$, and $\mathbf{y}^{(3)}=[1,1,0]$.
-The resulting match quality is $\frac{1}{3}(\theta_{1,3}+\theta_{3,1})=0.5$ and match rate is $\frac{2}{3}$
-2.2 Model Fidelity
-We discuss our model fidelity for real-world patient-provider matching:
-  1. 1.
-    Patient autonomy is a critical part of our model, as assortments allow patients to select between providers while allowing administrators some control.
-Prior work demonstrated that people are willing to trade autonomy for efficiency, including in online markets (Diederich et al., 2025) and domestic life (Munro et al., 2014).
-We find similar ideas during our conversations with our healthcare partners, where patients might prefer a few high-quality choices to an overwhelmingly large number of options (Chernev et al., 2015).
-  2. 2.
-    Match quality and match rate underlie assortment computation.
-We focus on match quality because it reflects the benefit patients get from matching, while match rate reflects the fraction of patients with access to providers.
+The resulting match rate is $\frac{2}{3}$ and match quality is $\frac{1}{3}(\theta_{1,3}+\theta_{3,1})=0.5$.
+2.2 Relevance to Patient-Provider Matching
+We provide additional details on the reasons behind the assumptions made in Section 2.1 and their connections to realities in healthcare.
+Offline-Online Setting
+We assume that patients are offered assortments offline and then respond in an online manner; this choice reflects logistical constraints and administrative burden.
+While pure offline assignment limits autonomy, pure online assortments or dynamically modifying the assortments online results in two problems: a) potentially misaligned incentive for patients to wait to make decisions, and b) an additional logistical burden.
+If assortments are dynamically modified, then, for example, patients can see different providers depending on the day they log in to the patient portal (e.g., A-B-C on Monday, X-Y-Z on Wednesday), which leads to confusion and incentivizes waiting.
+Dynamic offering might also be infeasible if patient notifications are sent in an analog manner (e.g., snail mail), and offering assortments in an online fashion might require administrators to continually update and reorganize assortments.
+In Section 4.3, we show that relaxing the model to a fully online setting has little impact on our overall takeaways.
+Assumptions
+We make several key assumptions in the model definition: a) providers each have unit capacity, b) patients respond in a uniform random order, and c) the match quality $\theta$ is perfectly known.
+Our core analytical results rely on these assumptions, but we perform sensitivity analyses to study our policies’ performance in more general settings.
+We detail how to relax the unit capacity assumption and provide results in Section 4.3 and demonstrate that this has little impact on our policies or findings.
+We assume that patients respond in random order to simplify the problem setup, and to better understand this assumption in practice, we experiment with non-uniform random orderings in Section 4.3.
+Finally, while healthcare administrators do not currently explicitly estimate match quality $\theta$, this notion is anecdotally incorporated into rematching decisions, such as finding proximal providers based on a patient’s location.
+To quantify match quality, administrators could collect preferences over providers for each patient and estimate rankings over providers via a learning-to-rank algorithm (Cao et al., 2007); we leave this estimation problem to future study.
+Even if $\theta$ is not perfectly known, we demonstrate in Appendix A that estimates of $\theta$ suffice to allow for good performance.
+Model Fidelity
+Our model aims to capture real-world concerns by balancing patient autonomy with system-wide utility through assortment offer decisions.
+The emphasis on patient autonomy arises from our conversations with healthcare partners, where patients might prefer a few high-quality choices to an overwhelmingly large number of options (Chernev et al., 2015), and from prior work that demonstrates people are willing to trade autonomy for efficiency, including in online markets (Diederich et al., 2025) and domestic life (Munro et al., 2014).
+Here, we intentionally leave the definition of “quality” ambiguous, as the definition of match quality should reflect patient preferences.
 For example, if patients primarily care about proximity, match quality should be based on the distance between patients and providers.
-We discuss different notions of match quality in Section 5.3.1.
-  3. 3.
-    Choice models reflect how patients select providers.
-Choice models are critical to our setting, so we analyze various choice models in Section 3 and experiment with different selections in Section 5.2.
-We focus on choice models that are clinically motivated or ubiquitous in prior work.
-For example, the uniform choice model reflects low-effort patient decisions (Salisbury, 1989), while the MNL choice model is commonly seen in prior work (Davis et al., 2013).
+We discuss this further in Section 5.1.
 2.3 Analysis in the Single Provider Scenario
-To gain insight into the optimal match quality policy structure, we analyze the $M=1$ scenario under the uniform choice model.
-We first explicitly detail the optimal policy: {theorem}[] Let $f_{i}$ be the uniform choice model with probability $p$.
-Suppose $M=1$, and let $u_{1},u_{2},\ldots,u_{N}$ be a set of coefficients such that $\theta_{u_{1}}\geq\theta_{u_{2}}\cdots\theta_{u_{N}}$.
+To gain insight into the optimal policy for both match rate and match quality, we analyze the situation when $M=1$:
+Theorem 2.1.
+Let $f$ be the uniform choice model with probability $p$.
+Suppose $M=1$, and let $u_{1},u_{2},\ldots,u_{N}$ be a permutation of $\{1,\ldots,N\}$ such that $\theta_{u_{1},1}\geq\theta_{u_{2},1}\cdots\theta_{u_{N},1}$; that is patient $u_{1}$ has the highest match quality, followed by $u_{2}$, etc..
 Let $s$ be defined as follows:
-$s=\operatorname*{arg\,max}_{s}(1-(1-p)^{s})\frac{\sum_{i=1}^{N}\theta_{u_{i},1}}{s}$ (7)
-Then the policy which maximizes match quality is $\mathbf{X}_{u_{1},1}=\mathbf{X}_{u_{2},1}\cdots\mathbf{X}_{u_{s},1}=1$, where $\mathbf{X}_{i,1}$ is 0 otherwise.
-When $M=1$, we can decompose the match quality into the probability any patient matches and the average match quality given matching.
-If we offer the single provider to $s$ patients, then the match probability is $(1-(1-p)^{s})$.
-For fixed $s$, we offer provider $j=1$ to the top-$s$ patients by match quality.
-Note that the optimal match rate is achieved with $s=N$, while the optimal match quality might arise from $s<N$ (as we’ll show in Section 3.5).
-We include all full proofs in Appendix B.
-3 Constructing Matching Policies
-We introduce policies $\pi(\theta)$ to construct assortments.
-We first introduce a baseline greedy policy and discuss why such a policy can perform poorly.
-We then propose three new policies: pairwise, which assigns each provider a single patient, group-based, which improves upon the weaknesses of the pairwise policy, and gradient descent, which optimizes for a lower bound on the match quality.
-3.1 Greedy Policy
+$s=\operatorname*{arg\,max}_{s}(1-(1-p)^{s})\frac{\sum_{i=1}^{s}\theta_{u_{i},1}}{s}$ (7)
+Then the policy which maximizes match rate is $\pi(\theta)=\mathbf{X}=\mathbf{1}$ while the policy which maximizes match quality is $\pi(\theta)=\mathbf{X},\mathbf{X}_{u_{1},1}=\mathbf{X}_{u_{2},1}\cdots\mathbf{X}_{u_{s},1}=1$, where $\mathbf{X}_{i,1}$ is 0 otherwise.
+We include all full proofs in Appendix F, and sketch the brief intuition here.
+Here $s$ represents the number of patients we offer the single provider to, and $u$ represents the ordering of patient match qualities for a given provider.
+When $M=1$, we can decompose the match quality into the probability any patient matches and the average match quality.
+When the single provider is offered to $s$ patients, the match rate is $(1-(1-p)^{s})$, while the average match quality for the top $s$-patients is $\frac{\sum_{i=1}^{N}\theta_{u_{i},i}}{s}$.
+Therefore, to maximize match rate we let $s=N$, while for match quality, we find the $s$ which maximizes the product of the average match quality and the match rate.
+Our results demonstrate a tradeoff between match rate and match quality even in the single provider scenario, and we next detail algorithms to generalize this idea.
+3 Constructing Assortment Policies
+We develop and analyze policies $\pi(\theta)$ to construct assortments that optimize for match rate and match quality.
+We first introduce a baseline greedy policy, which offers all providers to all patients.
+While the greedy policy maximizes the match rate, it can lead to poor match quality, motivating the need for more complex policies.
+As a result, we propose three new policies: pairwise, group-based, and gradient descent in Section 3.1 and analyze their match rate and match quality in Section 3.2 and Section 3.3 (summarized in Table 1).
+We first introduce an illustrative example and use this throughout this Section.
+Example 3.1.
+Consider a scenario with $p=0.75$ and $\theta=[0.7,0.7,0.1]$. According to Theorem 2.1, the optimal assortment is $\mathbf{X}=[1,1,0]$, which achieves a match quality of $0.22$.
+3.1 Assortment Policies
+3.1.1 Greedy Policy
 The greedy policy offers all providers to each patient and grants patients full autonomy.
@@
 For match quality, the greedy policy considers assortments myopically and ignores the structure of the match quality matrix, which can lead to poor performance.
-For example, if $M=1$ and $\theta_{1,1}>>\theta_{1,2}$, the greedy policy gives both patients equal chances of matching, whereas the optimal policy offers $j=1$ only to patient $i=1$.
-Formally, let $\pi^{*}$ be the optimal policy, $\pi^{*}(\theta)=\operatorname*{arg\,max}_{\mathbf{X}}\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$.
-We first prove that $\pi^{R}$ is an $\epsilon$-approximation:
-{theorem}
-[] Let $f_{i}$ be the uniform choice model with match probability $p$. For any $p$ and $\epsilon$, there exists a $\theta$ such that
-$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{R}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq\epsilon\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (8)
-We prove this by generalizing the example where $\theta_{1,1}>>\theta_{2,1}$ to arbitrary $N$ and $M$.
+For example, if $M=1$ and $\theta_{1,1}>>\theta_{2,1}$, the greedy policy gives both patients equal chances of matching, whereas the optimal policy offers $j=1$ only to patient $i=1$.
+3.1.2 Pairwise Policy
+To motivate the need for policies beyond greedy, we demonstrate that the greedy policy is sub-optimal in Example 3.1.
+Recall that $\theta=[0.7,0.7,0.1]$, so it is better for patients 1 or 2 to match rather than patient 3.
+The greedy policy has $\pi^{R}(\theta)=[1,1,1]$, which leads to an average match quality of $\frac{(1-(1-p)^{N})\frac{1}{N}\sum_{i=1}^{N}\theta_{i,1}}{N}=0.16$.
+Greedy achieves a low match quality because it offers a provider to patient $3$, which drags down the average match quality.
+More broadly, the greedy policy can offer providers to too many patients, which can lead to sub-optimal match qualities because assortments are not tailored.
+To better tailor assortments to patients, we introduce the pairwise policy $\pi^{P}(\theta)$, which contrasts with the greedy policy by offering each patient at most one provider.
+The pairwise policy, $\pi^{P}(\theta)$, pairs patients and providers by solving the weighted bipartite matching problem:
+$\max\limits_{\begin{subarray}{c}\pi^{P}(\theta),\\ \sum_{i=1}^{N}\pi^{P}(\theta)_{i,j}\leq 1,\\ \sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\leq 1\end{subarray}}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (8)
+This strategy better incorporates match quality structure at the expense of fewer patient choices. We note that we can optimize $\pi^{P}(\theta)$ by solving a linear program.
+3.1.3 Group-Based Policy
+While the pairwise policy can potentially better tailor assortments to patients, it can suffer from assortments that are too constrained.
+For example, returning to Example 3.1, the pairwise policy offers $\mathbf{X}=[1,0,0]$, which achieves a match quality of $0.18$.
+While the pairwise policy improves upon the greedy policy for match quality, it is still not the optimal assortment ($\mathbf{X}=[1,1,0]$) because it only offers each provider at most one patient.
+[…]
+$\max\limits_{\mathbf{q}}\sum_{i}\sum_{i^{\prime}>i}q_{i}q_{i^{\prime}}\alpha_{i,i^{\prime}}$ (9)
+We note that we can find such a solution quickly by letting $z_{i,j}=q_{i}q_{j}$, noting that $q_{i}\in\{0,1\}$, and solving at most $N$ linear programs (though in practice, it requires far fewer computations).
+We repeatedly solve for $\mathbf{q}$ and construct assortments from this.
+By constructing assortments in this group-based format, we guarantee that the match rate of our group-based policy matches the pairwise policy; we provide details in Section 3.2.
+We formalize this procedure in Algorithm 1.
+Algorithm 1 Group-based policy ($\pi^{G}$)
+1:  Input: Match quality matrix $\theta$
+2:  Output: Assortment, $\mathbf{X}$
+3:  Initialize the assortment $\mathbf{X}=\pi^{P}(\theta)$
+4:  for all $(i,i^{\prime})\subseteq[N]$ do
+5:     $\mathbf{X}^{\prime}=\pi^{P}(\theta),{X}^{\prime}_{i,v_{i^{\prime}}}=X^{\prime}_{i^{\prime},v(\theta)_{i}}=1$
+6:     Let $\alpha_{i,i^{\prime}}=\mathrm{MQ}(\mathbf{X}^{\prime},\theta,f)-\mathrm{MQ}(\pi^{P}(\theta),\theta,f)$
+7:  end for
+8:  Let $s=\infty$ and $r=\{1,\ldots,N\}$
+9:  while $s>0$ do
+10:     Let $s=\max\limits_{\mathbf{q}}\sum_{i\in r}\sum_{\begin{subarray}{c}i^{\prime}\in r\\ i^{\prime}>i\end{subarray}}q_{i}q_{i^{\prime}}\alpha_{i,i^{\prime}}$, $\mathbf{q}^{*}=\arg\max\limits_{\mathbf{q}}\sum_{i\in r}\sum_{\begin{subarray}{c}i^{\prime}\in r\\ i^{\prime}>i\end{subarray}}q_{i}q_{i^{\prime}}\alpha_{i,i^{\prime}},$
+11:     Let $X_{i,v_{i^{\prime}}}=1$ for all $i,i^{\prime}$ with $q^{*}_{i}=q^{*}_{i^{\prime}}=1$
+12:     Remove all $i$ with $q^{*}_{i}=1$ from $r$
+13:  end while
+3.1.4 Gradient Descent Policy
+We finally propose the gradient descent policy $\pi^{D}$, which generalizes beyond pairwise interactions by optimizing for a surrogate heuristic for match quality under the uniform choice model.
+The heuristic consists of two functions: $h(\mathbf{X})\in[0,1]^{N\times M}$, where $h(\mathbf{X})_{i,j}$ denotes the probability patient $i$ has provider $j$ available in their assortment, and $g(h(\mathbf{X}))\in[0,1]^{N\times M}$, where $g(h(\mathbf{X}))_{i,j}$ denotes the probability that provider $j$ is the most-preferred available provider for patient $i$.
+Here, $g$ and $h$ are known and defined below, so we aim to find a $\mathbf{X}$ to optimize $p\cdot\langle g(h(\mathbf{X})),\theta\rangle$.
+Under the uniform choice model, the probability that patient $i$ matches with provider $j$ is then $p\cdot g(h(\mathbf{X}))$, and so the expected match quality is $p\cdot\langle g(h(\mathbf{X})),\theta\rangle$.
+To compute $h(\mathbf{X})$ and $g(h(\mathbf{X}))$, first recall that $\sigma$ is a random ordering of patients, and $\sigma_{t}=i$ denotes that patient $i$ is the t-th patient to select their provider.
+To find $h(\mathbf{X})_{i,j}$, we multiply $X_{i,j}$, which indicates if $j$ is offered to $i$, by the probability that $j$ is still available when patient $i$ arrives.
+We approximate the latter by considering the availability probability for each fixed arrival point (i.e., if $\sigma_{t}=i$) and average over all possible $N$ arrival orders.
+Clearly, if $\sigma_{1}=i$, then $h(\mathbf{X})_{i,j}=X_{i,j}$, as no previous patient could have selected provider $j$.
+If $\sigma_{t}=i$, then the probability that none of the previous $t-1$ patients select provider $j$ is at least $(1-p)^{t-1}$; at most $t-1$ prior patients could have been offered $j$ and had $j$ as their most preferred provider.
+Because $t$ is uniformly distributed from $1$ to $N$:
+$h(\mathbf{X})_{i,j}=X_{i,j}\sum_{t=1}^{N}\frac{(1-p)^{t-1}}{N}$ (10)
+We can further tighten this by considering the number of patients who are offered provider $j$, namely $n=\lVert\mathbf{X}_{*,j}\rVert_{1}$ out of $N$ patients.
+Since only $n$ patients are offered provider $j$, we rescale the selection probabilities accordingly: $p^{\prime}=\left(\frac{n-1}{N-1}\right)p$.
+We can then update our calculation of $h(\mathbf{X})_{i,j}$ as:
+$h(\mathbf{X})_{i,j}=X_{i,j}\sum_{t=1}^{N}\frac{(1-p^{\prime}_{j})^{t-1}}{N}\ \mathrm{where}\ p^{\prime}_{j}=\left(\frac{\lVert\mathbf{X}_{*,j}\rVert_{1}-1}{N-1}\right)p$ (11)
+We then compute $g(h(\mathbf{X}))$ by first computing a preference ordering over providers for patient $i$.
+Let $u_{i,1},u_{i,2},\ldots,u_{i,M}$ be the most to least preferred providers for patient $i$; that is, $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,2}}\geq\cdots\geq\theta_{i,u_{i,M}}$.
+Then for patient $i$ to pick their kth best option, it requires that 1) $u_{i,k}$ is available, and 2) none of $u_{i,1},u_{i,2},\ldots,u_{i,k-1}$ are available.
+Formally, we can estimate the probability that the kth most preferred provider is the best available as:
+$g(h(\mathbf{X}))_{i,u_{i,k}}=h(\mathbf{X})_{i,u_{i,k}}\prod_{k^{\prime}=1}^{k-1}(1-h(\mathbf{X})_{i,u_{i,k^{\prime}}})$ (12)
+Using $h$ and $g$, we can compute a differentiable heuristic and in Section 3.3.4, we demonstrate that this heuristic lower bounds the total match quality.
+In practice, we use gradient descent and include a log regularization term to ensure that $\mathbf{X}$ is binary.
+While the heuristic is not always convex, in Section 4, we demonstrate that the policy is still able to find well-performing solutions, implying that a well-performing minimum is found despite guarantees on convexity.
+Moreover, such solutions can be found quickly in practice because the gradient descent algorithm quickly converges to local minima (see Section 4).
+Algorithm 2 Gradient Descent Objective
+  Input: Match quality matrix $\theta$ and Assortment $\mathbf{X}$
+  Output: Estimated total match quality
+  Let $p^{\prime}_{j}=\frac{\lVert\mathbf{X}_{*,j}\rVert_{1}-1}{N-1}$
+  Let $h(\mathbf{X})_{i,j}=\frac{X_{i,j}}{N}\sum_{t=1}^{N}(1-p^{\prime}_{j})^{t-1}$
+  Let $u_{i,1},u_{i,2},\ldots,u_{i,M}$ be a permutation of $[M]$, so $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,1}}\cdots\theta_{i,u_{i,M}}$
+  Let $g(h(\mathbf{X}))_{i,u_{i,k}}=h(\mathbf{X})_{i,u_{i,k}}\prod_{k^{\prime}=1}^{k-1}(1-h(\mathbf{X})_{i,u_{i,k^{\prime}}})$
+  return  $p\cdot\langle g(h(\mathbf{X})),\theta\rangle$
+3.2 Match Rate Analysis
+Table 1: We prove lower bounds for the match rate (MR) and match quality (MQ) for all four policies.
+We bold all tight bounds.
+The main takeaway is that the greedy policy maximizes match rate, but is poor for match quality, incentivizing the need for better policies.
+Policy Match Rate (MR) Match Quality (MQ) Notes
+Greedy $\geq\min(p,\frac{M}{N})$ $<\boldsymbol{\epsilon}\ \forall\ \epsilon$
+Pairwise $=\mathbf{p\frac{\min(M,N)}{N}}$ $\geq\mathbf{p\mathrm{MQ}(\pi^{*}(\theta),\theta,f)}$
+Group-Based $=\mathbf{p\frac{\min(M,N)}{N}}$ Heuristically improves MQ vs. pairwise
+Gradient Descent $<p\frac{\min(M,N)}{N}$ Provably optimizes lower bound of MQ
+We formally characterize the match rate for the four policies, and find that the greedy policy theoretically performs best, while gradient descent performs worst.
+We focus on the uniform choice model due to its simplicity and ability to model the decision-making of real-world patients (Salisbury, 1989).
+We include proof sketches here, while full proofs are deferred to Appendix F.
+3.2.1 Greedy Policy
+We first characterize the performance of the greedy policy:
+Theorem 3.1.
+Let $f$ be the uniform choice model with match probability $p$. Then
+$\mathrm{MR}(\pi^{R}(\theta),\theta,f)\geq\min(p,\frac{M}{N})$ (13)
+Moreover, there exists no policy $\pi^{\prime}$ such that $\mathrm{MR}(\pi^{\prime}(\theta),\theta,f)>\mathrm{MR}(\pi^{R}(\theta),\theta,f)$
+We demonstrate this by considering two scenarios; one in which all providers are matched and the other in which each patient gets a chance to match.
+Under the uniform choice model, the greedy policy is optimal for the match rate because it offers the largest assortments.
+3.2.2 Pairwise Policy
+We similarly characterize the match rate of the pairwise policy:
+Theorem 3.2.
+Let $f$ be the uniform choice model with match probability $p$. Then
+$\mathrm{MR}(\pi^{P}(\theta),\theta,f)=p\frac{\min(M,N)}{N}$ (14)
+Here, we prove this by noting that assortments are disjoint, and each patient with a non-empty assortment matches with probability $p$.
+Noticing that there are $\min(M,N)$ non-empty assortments yields the desired match rate.
+We note that this bound is tight due to the disjoint nature of patient assortments.
+When $M\geq N$, the pairwise policy achieves the optimal match rate, while when $M\leq N$, the pairwise policy performs worse than optimal because the same provider should be offered to multiple patients.
+3.2.3 Group-Based Policy
+Because we construct the group-based policy so assortments are shared across patients in a group, we can show that the group-based policy maintains the same match rate as the pairwise policy:
+Theorem 3.3.
+Let $f$ be the uniform choice model. Then
+$\mathrm{MR}(\pi^{G}(\theta),\theta,f)=\mathrm{MR}(\pi^{P}(\theta),\theta,f)$ (15)
+We prove this by showing that the number of patients with an empty assortment is the same in both the group-based and pairwise policies, leading to the same number of matches in both policies.
+Providing structure, as we do for the group-based policy, is necessary to ensure match rates; once this structure is abandoned, it becomes difficult to ensure high match rates.
+3.2.4 Gradient Descent Policy
+We characterize the match rate for the gradient descent policy in two steps.
+First, we demonstrate that there exist scenarios where certain maximums for $g(h(\mathbf{X}))$ lead to a poor match rate.
+Next, we show that even when the maximum is unique, it can still lead to a worse match rate compared with the pairwise policy.
+Taken together, these results demonstrate that optimizing the match quality heuristic can lead to tensions with the match rate.
+Example 3.2.
+Consider $\theta_{i,j}=0$ except at $\theta_{1,1}=1$.
+Then for any assortment $\mathbf{X}$, only $g(h(\mathbf{X}))_{1,1}$ contributes to the match quality.
+Therefore, to maximize $g(h(\mathbf{X}))_{1,1}$, we set $\mathbf{X}_{1,1}=1$ and $\mathbf{X}_{i,1}=0\ \forall\ i$.
+We note that the remaining $\mathbf{X}$ can be set arbitrarily.
+Therefore, letting $\mathbf{X}_{1,1}=1$ and $\mathbf{X}_{i,j}=0$ otherwise yields match rates as low as $\frac{1}{N}$.
+We next show that even when solutions are unique, the gradient descent policy can still lead to sub-optimal match rates:
+Theorem 3.4.
+Let $f$ be the uniform choice model with match probability $p$.
+Then for any $p<1$, there exists $\theta$ such that the unique maximizer of $\langle g(h(\mathbf{X})),\theta\rangle$, $\mathbf{X}^{*}$, has:
+$\mathrm{MR}(\mathbf{X}^{*},\theta,f)<\mathrm{MR}(\pi^{P}(\theta),\theta,f)$ (16)
+We prove this by constructing a simple $N=M=2$ example where even the unique maximizer can achieve a sub-optimal match rate.
+Here, the unique maximizer also corresponds to the true maximizer for match quality, and so we demonstrate a match rate-match quality tradeoff.
+Overall, we find a tension between policies that optimize for match rate, such as greedy, and those that optimize for match quality, such as gradient descent.
+3.3 Match Quality Analysis
+We next complement our match rate analysis by analyzing the match quality for various policies.
+We show that the greedy policy, which performs best for match rate, can perform arbitrarily poorly for match quality.
+We then demonstrate that our three new policies can improve match quality, showing a tradeoff between match rate and match quality.
+3.3.1 Greedy Policy
+We start by demonstrating that the greedy policy $\pi^{R}$ can perform poorly when compared with the optimal match rate policy $\pi^{*}$. To demonstrate this, we first define $\pi^{*}$:
+$\pi^{*}(\theta)=\operatorname*{arg\,max}_{\mathbf{X}}\mathrm{MQ}(\mathbf{X},\theta,f)$ (17)
+We then demonstrate that $\pi^{R}$ is an $\epsilon$ approximation:
+Theorem 3.5.
+Let $f$ be the uniform choice model with match probability $p$. For any $p$ and $\epsilon$, there exists a $\theta$ such that
+$\mathrm{MQ}(\pi^{R}(\theta),\theta,f)\leq\epsilon\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (18)
+We prove this by generalizing an example where $\theta_{1,1}>>\theta_{2,1}$ to arbitrary $N$ and $M$.
 Here, the optimal policy is to offer each provider to only one patient.
 However, the greedy policy offers all providers to all patients, which leads to sub-optimal matches for each provider $j$.
-Moreover, we show that we can construct scenarios with $N=M=\frac{3}{\epsilon p}$ where the greedy policy is an $\epsilon$-approximation.
-For large $p$, even small $M$ and $N$ lead to situations where the greedy policy performs poorly.
-3.2 Pairwise Policy
-The pairwise policy inverts the greedy policy by offering each patient at most one provider.
-This strategy reduces patient choices while better incorporating match quality structure.
-The pairwise policy, $\pi^{P}(\theta)$, pairs patients and providers by solving the weighted bipartite matching problem:
-$\max\limits_{\pi^{P}(\theta),\sum_{i=1}^{N}\pi^{P}(\theta)_{i,j}\leq 1,\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\leq 1}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (9)
-Under a uniform choice model, the choice of $p$ impacts the pairwise policy performance: {theorem}[] Let $f_{i}$ be the uniform choice model with match probability $p$. Then
-$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\geq p\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (10)
+3.3.2 Pairwise Policy
+We next show that the pairwise improves the performance guarantee from $\epsilon$ to $p$ for match quality:
+Theorem 3.6.
+Let $f$ be the uniform choice model with match probability $p$. Then
+$\mathrm{MQ}(\pi^{P}(\theta),\theta,f)\geq p\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (19)
 To prove this, we first upper bound the optimal policy, $\pi^{*}(\theta)$, with the value of the underlying bipartite matching problem.
@@
 Summing over edges gives that the pairwise policy is a $p$-approximation to the bipartite matching problem and is, therefore, a $p$-approximation to $\pi^{*}$.
-Beyond the uniform choice model, we can show a similar guarantee in the MNL scenario.
-For the MNL choice model, the exit option $\gamma$ is analogous to $p$ in controlling policy performance.
-First, let $v_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$ and let $v_{i}=-1$ if $\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}=0$.
-Then the MNL choice model is equivalent to the uniform choice model with $p_{i}=\frac{\mathbbm{1}[v_{i}\geq 0]\exp(\theta_{i,v_{i}})}{\exp(\gamma)+\mathbbm{1}[v_{i}\geq 0]\exp(\theta_{i,v_{i}})}$.
-We can then upper bound optimal match quality as $\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$, while the pairwise policy achieves $\sum_{i=1}^{N}\frac{p_{i}}{N}\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$.
-Let $a_{i}=\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$, then the pairwise policy achieves an approximation ratio of:
-$\frac{\sum_{i=1}^{N}a_{i}\frac{\exp(a_{i})}{\exp(\gamma)+\exp(a_{i})}}{\sum_{i=1}^{N}a_{i}}\geq\frac{\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})(\frac{1}{N}\sum_{i=1}^{N}a_{i})}{(\exp(\gamma)+\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i}))(\frac{1}{N}\sum_{i=1}^{N}a_{i})}\geq\frac{\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})}{\exp(\gamma)+\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})}$ (11)
-Small $\gamma$ improves pairwise performance, corresponding to patients more inclined to match.
-3.3 Group-Based Policy
-The pairwise policy performs poorly for small $p$, so we develop the group-based policy.
-When $p$ is small, administrators should offer patients larger assortments because a patient’s top preference is more likely to be available.
-For example, consider a scenario with $N$ patients and $M$ providers.
-We demonstrate a performance gap between greedy and pairwise policy dependent on $p$:
+We note that when $M=1$ and $N$ is large, with $\theta_{i,1}=1\ \forall\ i$, this bound becomes tight, as the optimal match quality is $1$, while pairwise achieves $p$.
+3.3.3 Group-Based Policy
+We next analyze the match quality for the group-based policy.
+While proving a bound on the match quality for the group-based policy is combinatorially difficult, we give intuition for when a group-based policy is necessary by demonstrating the need to modulate between the greedy and pairwise policies:
 Proposition 3.1.
-Let $f_{i}$ be the uniform choice model with match probability $p$. If $\theta_{i,j}\sim U(0,1)$ and $M\leq N$ then
-$\frac{\mathbb{E}_{\sigma,\theta}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{R}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]}{\mathbb{E}_{\sigma,\theta}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]}\geq\frac{1-(1-p)^{N/M}}{2p}$ (12)
-We demonstrate this by showing that the pairwise policy achieves a match quality of $Mp$.
-We then compute the match probability for the greedy policy as $1-(1-p)^{N/M}$ for each of the $M$ providers, and noting that the match quality for each matched provider is at least $\frac{1}{2}$ in expectation.
-The pairwise policy performs poorly for small $p$, so we develop a new policy that modulates between the pairwise policy (for large $p$) and the greedy policy (for small $p$).
-To do this, we add a provider $j^{\prime}$ to the assortment of patient $i$ if $\theta_{i,j^{\prime}}\geq\theta_{i,v_{i}}$; that is, patients should be assigned preferred providers.
-However, such a policy could decrease match rates because some patients have empty assortments.
-Example 2.
-Let $\theta$ be the following:
-$\theta=\begin{bmatrix}0.6&0.7\\ 0.3&0.6\\ \end{bmatrix}$ (13)
-Here, $\pi^{P}(\theta)_{1,1}=\pi^{P}(\theta)_{2,2}=1$.
-We add provider $j=2$ to patient $i=1$, so $X_{i,2}=1$.
-Let $\sigma=[1,2]$.
-Then, with probability $p$, provider $j=2$ is selected by patient $i=1$ and $\mathbf{y}^{(2)}=[1,0]$ so patient $i=2$ has no options.
-This reduces the match rate from $p$ to $p-\frac{p^{2}}{2}$.
-We overcome this issue by developing the group-based policy, $\pi^{G}(\theta)$.
-The group-based policy constructs groups so that each group receives the same assortment.
-In Example 2, this would mean that $\mathbf{X}_{1}=\mathbf{X}_{2}=[1,1]$.
-We demonstrate that such a grouping strategy is the only way to maintain a match rate while building upon the pairwise policy: {theorem}[] Let $\pi$ be a policy that augments the pairwise policy; $\pi(\theta)_{i,j}\geq\pi^{P}(\theta)_{i,j}$ for any $\theta$ for all $i,j$.
-Let $G=(V,E)$ be a directed graph with $N$ nodes such that nodes $i$ and $i^{\prime}$ are connected if $\pi(\theta)_{i,v(\theta)_{i^{\prime}}}=1$ Here, $v(\theta)_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$.
-If $N=M$, then
-$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}\left(\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\rVert_{1}]=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\rVert_{1}]=p$ (14)
-if and only if each component in $G$ is a complete digraph.
-We prove this by showing that when $N=M$, all patients can match with a provider with probability $p$, leading to an overall match rate of $p$.
-To construct groups, we first compute edge weights for each pair of patients in $G$, denoted as $\alpha_{i,i^{\prime}}$.
-Here, $\alpha_{i,i^{\prime}}$ captures the benefit when placing $(i,v_{i})$ and $(i^{\prime},v_{i^{\prime}})$ in the same assortment.
-Next, we construct groups to maximize the sum of pairwise edges within each group.
-We do so through a linear program, with $z_{i,j}$ representing whether we include edge $(i,j)$ and $q_{i}$ representing whether we include node $i$ in the group.
-We place details on computing $\alpha_{i,i^{\prime}}$ and constructing the assortments in Algorithm 1.
-Algorithm 1 Group-based policy ($\pi^{G}$)
-1:  Input: Pairwise policy, $\pi^{P}$ and match quality $\theta$
-2:  Output: Assortment, $\mathbf{X}$
-3:  Initialize $\mathbf{X}=\mathbf{X}^{P}=\pi^{P}(\theta)$, $v_{i}=j$ if $X_{i,j}^{P}=1$, and $\mathbf{c}=\mathbf{1}$
-4:  for all $(i,i^{\prime})\subseteq[N]$ do
-5:     Let $\mathbf{X}^{\prime}=\mathbf{X}^{P}$
-6:     Let ${X}^{\prime}_{i,v_{i^{\prime}}}=1$ and $X^{\prime}_{i^{\prime},v_{i}}=1$
-7:     Let $\alpha_{i,i^{\prime}}=\mathbb{E}_{\sigma\sim[\{i,i^{\prime}\},\{i^{\prime},i\}]}[\sum_{t=1}^{2}(f_{\sigma_{t}}\left(\mathbf{X}^{\prime}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}-f_{\sigma_{t}}\left(\mathbf{X}^{P}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$
-8:  end for
-9:  Let $s=\infty$ and $r=\{1,\ldots,N\}$
-10:  while s ¿ 0 do
-11:     Let $s=\max\limits_{\mathbf{z}}\sum_{i\in r}\sum_{i^{\prime}\in r}z_{i,j}\alpha_{i,i^{\prime}}$, subject to $z_{i,j}=q_{i}q_{j}$, and let $\mathbf{z}$ be the corresponding solution
-12:     for $i,i^{\prime}\in r$ with $z_{i,i^{\prime}}=1$ do
-13:        Let $X_{i,v_{i^{\prime}}}=X_{i^{\prime},v_{i}}=1$
-14:        Remove $i,i^{\prime}$ from $r$
-15:     end for
-16:  end while
-3.4 Gradient Descent Policy
-The group-based policy makes local improvements to the pairwise policy but can fail because it only considers pairwise interactions.
-To fix this, we develop the gradient descent policy by considering the global structure in $\theta$.
-Our gradient descent policy uses gradient descent to optimize an objective function $g(f(\mathbf{X}))$, which represents a lower bound on the match quality of an assortment $\mathbf{X}$ under the uniform choice model.
-Here, $f(\mathbf{X})_{i,j}$ lower bounds the probability that provider $j$ is available, while $g(f(\mathbf{X}))_{i,j}\in[0,1]^{N\times M}$ lower bounds the probability that provider $j$ is the top choice for patient $i$: $g(f(\mathbf{X}))\leq\mathrm{Pr}[j=\operatorname*{arg\,max}_{j^{\prime}}\theta_{i,j^{\prime}}\mathbf{y}^{(t)}_{j^{\prime}}]$.
-We first detail how to compute $f$ and $g$, then prove that this lower bounds the match quality (see Algorithm 2):
-  1. 1.
-    $f(\mathbf{X})$ lower bounds the availability probability for provider $j$.
-We first note that $f(\mathbf{X})_{i,j}=0$ if $X_{i,j}=0$.
-Next, availability depends on the response order $\sigma$.
-For example, if $\sigma_{1}=i$, then provider $j$ is available if $X_{i,j}=1$.
-Next, if $\sigma_{t}=i$, then under a uniform choice model, there is at least a $(1-p)^{t-1}$ availability probability.
-This holds because, in the worst case, the $t-1$ preceding patients also prefer provider $j$, so the probability none of the $t-1$ patients select $j$ is $(1-p)^{t-1}$.
-Next, if $n$ out of $N$ patients have provider $j$ on their assortment, then the probability that a given patient $i^{\prime}\neq i$ matches to $j$ is $p\frac{n-1}{N-1}$.
-That is, if patient $i$ is in the t-th position, then the availability probability is at least $(1-p\frac{n-1}{N-1})^{t-1}$.
-Averaging over all $N$ potential values for $t$ gives:
-    $f(\mathbf{X})_{i,j}=X_{i,j}h(\lVert\mathbf{X}_{*,j}\rVert_{1}),h(n)=\frac{1}{N}\sum_{t=1}^{N}(1-p\frac{n-1}{N-1})^{t-1}$ (15)
-  2. 2.
-    $g(f(\mathbf{X}))$ computes the probability that provider $j$ is the top available provider.
-Let $u_{i,1},u_{i,2},\ldots,u_{i,N}$ be coefficients so $\theta_{i,u_{1}}\geq\theta_{i,u_{2}}\cdots\theta_{i,u_{M}}$.
-Then provider $u_{i,1}$ is the top option whenever it is available; $g(f(\mathbf{X}))_{i,u_{i,1}})=f(\mathbf{X})_{i,u_{i,1}}$.
-Similarly, $u_{i,2}$ is only selected if $u_{i,1}$ is not selected; we can estimate $g(f(\mathbf{X}))_{i,u_{i,2}})=f(\mathbf{X})_{i,u_{i,2}}(1-f(\mathbf{X})_{i,u_{i,1}})$.
-We can generalize this as:
-    $g(f(\mathbf{X}))_{i,u_{i,k}}=f(\mathbf{X})_{i,u_{i,k}}\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{i,k^{\prime}}})$ (16)
-    The expected match quality is then $p(g(f(\mathbf{X}))\cdot\theta)$.
-Algorithm 2 Gradient Descent Objective
-  Input: Match quality $\theta$ and Assortment $\mathbf{X}$
-  Output: Estimated total match quality
-  Let $h(n)=\frac{1}{N}\sum_{t=1}^{N}(1-p\frac{n-1}{N-1})^{t-1}$
-  Let $f(\mathbf{X})_{i,j}=h(\lVert\mathbf{X}_{*,j}\rVert_{1})\mathbf{X}_{i,j}$
-  Let $u_{i,1},u_{i,2},\ldots,u_{i,M}$ be a permutation of $[M]$, so $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,1}}\cdots\theta_{i,u_{i,M}}$
-  Let $g(f(\mathbf{X}))_{i,u_{k}}=f(\mathbf{X})_{i,u_{k}}\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{k^{\prime}}})$
-  return  $p*g(f(\mathbf{X}))\cdot\theta$
-{theorem}
-The following holds for any $\mathbf{X}$:
-$p*(g(f(\mathbf{X}))\cdot\theta)\leq\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (17)
-Moreover, we demonstrate tightness when using the pairwise policy: {lemma}[] When $\mathbf{X}=\pi^{P}(\theta)$, then
-$p*(g(f(\mathbf{X}))\cdot\theta)=\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (18)
-We demonstrate this by showing that for any $i$, $\sum_{j=1}^{m}pg(f(\mathbf{X}))_{i,j}$ is an underestimation for the probability that any of the top-$m$ providers are available.
-We then show that because $\theta_{i,u_{i,j}}$ is sorted from high to low, we can similarly prove that $\sum_{j=1}^{m}p(g(f(\mathbf{X}))_{i,j}\theta_{i,j}$ is also an underestimation for each patient $i$.
-In the scenario where $\mathbf{X}=\pi^{P}(\theta)$, we can explicitly compute $g(f(\mathbf{X}))$ because each patient matches with at most one provider.
-Key to our policy is the tightness of the lower bound.
-When each patient receives more providers the lower bound is looser so the gradient descent policy performs worse.
-Conversely, when $N>>M$, each patient can only receive a few providers, so our policy performs better.
-Essentially, when there is low “inter-provider interference”, or patients receiving large assortments (relative to the number of providers), our policy performs better.
-The gradient descent policy can also naturally extend to optimize match rate if $\theta_{i,j}$ is constant; in that scenario, we optimize for $\sum_{i=1}^{N}\sum_{j=1}^{M}g(f(\mathbf{X}))_{i,j}$.
-3.5 Worked Example with Policies
-We motivate the need for different policies beyond greedy through a simple example with $N=3$ patients and $M=1$ providers.
-Example 3.
-Consider a scenario with $p=0.75$ and $\theta=[0.7,0.7,0.1]$.
-According to Theorem 2.3, the optimal assortment in such a scenario is $\mathbf{X}=[1,1,0]$, which achieves a match quality of $0.22$.
-We next compare our four policies:
-  1. 1.
-    Greedy offers $\mathbf{X}=[1,1,1]$, which leads to an average match quality of $\frac{(1-(1-p)^{N})\frac{1}{N}\sum_{i=1}^{N}\theta_{i,1}}{N}=0.16$.
-Greedy achieves a low match quality because it offers a provider to patient $i=3$, which drags down the average match quality.
-  2. 2.
-    Pairwise offers $\mathbf{X}=[1,0,0]$, which achieves a match quality of $0.18$.
-While the pairwise policy improves upon the greedy policy, the symmetry between patients one and two implies that it should offer a provider to patient $i=2$, as it can improve the match probability.
-  3. 3.
-    Group-Based first computes $\alpha_{1,2}=0.13$ and $\alpha_{1,3}=-0.15$ and constructs a group with patients 1 and 2. It then offers $\mathbf{X}=[1,1,0]$, the optimal assortment for match quality.
-  4. 4.
-    Gradient Descent optimizes for $g(f(\mathbf{X}))\cdot\theta=f(\mathbf{X})\cdot\theta=\sum_{i=1}^{N}X_{i,1}\theta_{i,1}h(\lVert\mathbf{X}_{*,j}\rVert_{1})$, which is optimizes at $\mathbf{X}=[1,1,0]$ and is the optimal assortment for match quality.
-The greedy policy fails because it offers the provider to patients with a low match quality, while the pairwise policy fails because it offers providers to too few patients.
-The group-based and gradient descent policies rectify this situation by adapting assortment size based on the match quality and correctly offering the provider to patients 1 and 2.
-4 Extending Policies to Real-World Considerations
-We extend our model to incorporate three real-world matching phenomena: 1) batch-offering assortments, 2) patient cognitive load, and 3) fairness considerations.
-4.1 Batch Offering Patients
-Thus far, we have assumed that all menus are constructed up front, in a single “batch.” Healthcare administrators can sacrifice logistical ease to improve match quality by making offers to groups of patients in batches rather than all at once.
-By offering assortments in batches, administrators increase control over the response order but increase the total response time for all patients to respond.
-By assumption, the response time for each patient is exponentially distributed (see Section 2.1), and the expected maximum of $N$ exponential variables is proportional to $\log(N)$ (Eisenberg, 2008).
-By offering $L$ batches, our total response time increases from $\log(N)$ to $L\log(N/L)$.
-We aim to construct batches, $b_{1},b_{2},\ldots,b_{L}\subseteq[N]$, so that the resulting response order maximizes match quality.
-Formally, let $S(b_{1},b_{2},\ldots,b_{L})$ be a distribution over response orders so that any patient $i\in b_{k}$ is before any patient $i^{\prime}\in b_{k^{\prime}}$ with $k<k^{\prime}$.
-Next, let $\sigma^{*}$ be the optimal ordering for some fixed policy $\pi$: $\sigma^{*}=\operatorname*{arg\,max}_{\sigma}\frac{1}{N}\sum_{t=1}^{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi(\theta,\sigma)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})$.
-We aim to set $b_{1},b_{2},\ldots,b_{L}$ so that $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ and $\sigma^{*}$ are similar.
-To do so, we first characterize $\sigma^{*}$ for a policy $\pi^{A}(\theta,\sigma)$, which augments the pairwise policy using knowledge of $\sigma$, $\pi^{A}(\theta,\sigma)_{i,j}\geq\pi^{P}(\theta)_{i,j}\forall i,j$.
-Recall that $v_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$, and let $v^{-1}_{j}=i$ if $v_{i}=j$.
-Then, $\pi^{A}(\theta)_{i,j}=1$ if $\theta_{i,j}\geq\theta_{i,v_{i}}$ and $\sigma^{-1}_{i}\geq\sigma^{-1}_{v^{-1}_{j}}$, where $\sigma^{-1}_{i}=t$ if $\sigma_{t}=i$.
-Essentially, $\pi^{A}(\theta)$ augments the pairwise policy $\pi^{P}(\theta)$ to add pairs $i,v_{i^{\prime}}$ for patients $i^{\prime}$ preceding $i$ in $\sigma$.
-We focus on such a policy because it represents a natural extension of the pairwise policy to incorporate the order $\sigma$.
-We characterize $\sigma^{*}$ for the policy $\pi^{A}(\theta)$ then use this to construct the batches $b_{k}$:
+Let $f$ be the uniform choice model with match probability $p$. If $\theta_{i,j}\sim U(0,1)$ and $M\leq N$ then
+$\frac{\mathbb{E}_{\theta}[\mathrm{MQ}(\pi^{R}(\theta),\theta,f)]}{\mathbb{E}_{\theta}[\mathrm{MQ}(\pi^{P}(\theta),\theta,f)]}\geq\frac{1-(1-p)^{N/M}}{2p}$ (20)
+When $p$ is small, the greedy policy can outperform pairwise because providers should be offered to multiple patients.
+Meanwhile, larger $p$ incentivizes smaller menus, leading to better performance for the pairwise policy.
+Additionally, when $N>M$ and $p<\frac{1}{2}$, the greedy policy can outperform pairwise due to provider scarcity.
+The group-based policy is built on this intuition so that groups are dynamically sized as needed.
+3.3.4 Gradient Descent Policy
+To analyze the match quality of the gradient descent policy, we demonstrate that our heuristic $p\cdot\langle g(h(\mathbf{X})),\theta\rangle$ lower bounds the match quality:
+Theorem 3.7.
+The following holds for any $\mathbf{X}$ when $f$ is the uniform choice model with probability $p$:
+$p\cdot\langle g(h(\mathbf{X})),\theta\rangle\leq\mathrm{MQ}(\mathbf{X},\theta,f)$ (21)
+We demonstrate this by showing that $\sum_{j}g(h(\mathbf{X}))_{i,j}$ is an underestimate for any fixed $i$.
+Our lower bound becomes tighter as $N$ becomes larger than $M$ because there is less overlap between different providers’ assortments due to scarcity.
+3.4 Analysis under Other Choice Models
+Our analysis thus far has been primarily under the uniform choice model, so we discuss policy performance under the MNL choice model with an exit option $\gamma$, a common choice model seen throughout the assortment optimization literature (Davis et al., 2013).
+We characterize the match rate and match quality for the pairwise policy under this model:
 Lemma 3.1.
-Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v_{i}=-1$ and $v_{i^{\prime}}\neq-1$.
-Let $\sigma^{*}$ be the optimal ordering for the policy $\pi^{A}$.
-Then, traversing the nodes defined by $\sigma^{*}$ is a reverse topological ordering on $G$.
-We demonstrate this by considering a pair of patients $i$ and $i^{\prime}$, so that $i$ precedes $i^{\prime}$ in any reverse topological order, but $i$ precedes $i^{\prime}$ in $\sigma^{*}$.
-We then use case work to show that we can either a) show that there is a cycle in the graph $G$, leading to a contradiction, b) recurse on a smaller pair of patients, $w,i$ or $i^{\prime},w$, or c) show that swapping $i$ and $i^{\prime}$ in $\sigma$ can only increase the match quality.
-We next construct $b_{1},b_{2},\ldots,b_{L}$, so $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ is close to the optimal ordering $\sigma^{*}$.
-The idea is to partition $\sigma^{*}$ into $L$ batches; then we know that all patients in $b_{1}$ precede $b_{2}$.
-To do this, we first demonstrate that if no pairs of patients within a batch share a common descendant, then such a batching preserves the match quality of $\sigma^{*}$: {theorem}[] Let $\sigma^{*}$ be an optimal ordering.
-Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v_{i}=-1$ and $v_{i^{\prime}}\neq-1$.
-Suppose that there exists a partition of $\sigma^{*}$ into $K$ batches, $b_{1},b_{2},\ldots,b_{L}$, such that for any partition $b_{k}$, no $i,i^{\prime}\in b_{k}$ have a common descendant in $G=(V,E)$.
-Then $\pi^{A}$ achieves the same match quality under partition $g$ as under the optimal ordering:
-$\mathbb{E}_{\sigma\sim S(b_{1},b_{2},\ldots,b_{L})}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{A}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]=\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma^{*}_{t}}\left(\pi^{A}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})$ (19)
-This holds because the selection made by patient $i$ depends only on the descendants of patient $i$ on the graph G.
-If $b_{1},b_{2},\ldots,b_{K}$ ensures that any pair of descendants are in different batches, then the order of descendants is maintained, so patients select the same providers in $\sigma$ and $\sigma^{*}$.
-In practice, we compute the batches $b_{k}$ through the following:
-  1. 1.
-    Computing the graph $G=(V,E)$ with $i$ connected to $i^{\prime}$ if $\theta_{i,v_{i^{\prime}}}>\theta_{i,v_{i}}$.
-  2. 2.
-    Computing a reverse topological ordering $\sigma^{*}$.
-  3. 3.
-    Selecting batches with the minimum number of edges within a batch (via dynamic programming).
-We note that this is an approximation of optimal batches for two reasons: 1) while $\sigma^{*}$ is a reverse topological ordering, there could be multiple reverse topological orderings, and 2) minimizing the number of edges within a batch does not guarantee that the order of descendants is maintained.
-4.2 Maximum Assortment Sizes
-To incorporate patient cognitive load (Chernev et al., 2015), we incorporate restrictions on the maximum assortment size into our model.
-Large assortments might be undesirable or practically infeasible because patients only consider a subset of the options they are presented due to logistical or time constraints.
-Formally, we suppose that patients only consider a random $d$-subset $\mathbf{z}^{(d)}$ of the options they are presented at random, $z^{(d)}_{j}\leq X_{i,j}y^{(t)}_{j}$ with $\lVert\mathbf{z}^{(d)}\rVert_{1}\leq d$.
-Patients make decisions according to $f_{i}(\mathbf{z}^{(d)})$ rather than $f_{i}(\pi(\theta)\odot\mathbf{y}^{(t)})$.
-When $d$ is small, policies should offer small tailored assortments, as large assortments lead to patients ignoring many options.
-4.3 Fair Matches
-Policies should aim to achieve fair matches to encourage participation and maintain equity.
-We construct examples of fairness metrics through the prior literature on fair matches and list these below (Xinying Chen and Hooker, 2023; García-Soriano and Bonchi, 2020).
-  • Minimum Match Quality - One natural objective is to maximize the minimum match quality across all matched patients.
-Such an objective is natural when administrators aim to maintain a certain level of care across patients.
-  • Variance in Match Quality - Another approach minimizes variance in match quality to ensure that the discrepancy in match quality between patients is small.
-  • Match Quality Range - One final approach to match quality fairness is to minimize the absolute difference between the maximum and minimum match quality for matched patients.
-Each notion of fairness has different implications for the optimal policy.
-For example, optimizing for the minimum match quality might lead policies to only offer assortments to high match quality patient-provider pairs, whereas minimizing the variance in match quality might yield many low quality matches.
-We explore tradeoffs in match quality and fairness in Section 5.4.
+Let $f$ be the MNL choice model parametrized by $\gamma$. Then the match rate and match quality of the pairwise policy are:
+$\mathrm{MR}(\pi^{P},\theta,f)$ (22)
+$=\frac{1}{N\sum_{i}\mathbbm{1}[v(\theta)_{i}\geq 0]}\sum_{i;v(\theta)_{i}\geq 0}\frac{\exp(\theta_{i,v(\theta)_{i}})}{\exp(\theta_{i,v(\theta)_{i}})+\exp(\gamma)}$ (23)
+and
+$\mathrm{MQ}(\pi^{P}(\theta),\theta,f)$ (24)
+$\geq\left(\frac{1}{N}\sum_{i=1}^{N}\frac{\exp(\theta_{i,v(\theta)_{i}})}{\exp(\theta_{i,v(\theta)_{i}})+\exp(\gamma)}\right)\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (25)
+To prove this, we apply similar techniques to those used with the uniform choice model by analyzing the individual match rate for each patient, and then summing across patients.
+We observe that the exit option parameter $\gamma$ plays a similar but inverted role to $p$ in the uniform choice model.
+As $\gamma$ increases, the probability a patient does not select any provider also increases, and so larger $\gamma$ incentivizes policies to offer larger assortments, similar to the incentives under smaller $p$ (see Proposition 3.1).
+We also note the additional complexity inherent to the MNL choice model.
+Guarantees under the MNL model require additional knowledge of $\theta$ and $v$, and analysis under the MNL choice model becomes difficult because match rates cannot be decoupled from individual match qualities.
+Nonetheless, the key takeaway is that a patient’s willingness to match, here controlled by $\gamma$, can dictate policy performance.
 4 Empirical Analysis
-We analyze policies in synthetic and real-world settings to characterize how assortment policy impacts match quality and rate.
-First, we demonstrate that gradient descent is the best policy when patients outnumber providers.
-We then show that group-based and greedy policies perform well for large and small values of match probability $p$, respectively.
-Our results demonstrate that assortment policies should be selected based on problem specifics.
+We complement our theoretical analysis through a study of our policies on a synthetic dataset.
+We characterize the best-performing policy based on model parameters, and show that the choice of assortment policy impacts match quality and match rate.
+Our biggest takeaway is that tailoring policies through gradient descent maximizes match quality, especially when patients outnumber providers.
 4.1 Experimental Details
@@
 MQ and norm.
-MR for short), which divides each quantity by that of the random policy.
-5.2 Synthetic Examples
-We construct synthetic examples to gain insight into our policies across different scenarios. For each example, we specify values for $M$, $N$, $f_{i}$, and $\theta$.
-Example 4.
-Figure 2: We find the optimal policy and compute the approximation ratio for match quality compared to optimal.
-While all non-random policies achieve approximation ratios $\geq 0.9$ when $M\geq N$ (left), the pairwise policy performs poorly and achieves a low approximation ratio when $N\geq M$ (right).
-Few patients and providers - We first analyze policies with a small number of patients $N$ and number of providers $M$.
-Small $N$ and $M$ allow us to compare to the optimal policy $\pi^{*}$ because we can compute $\pi^{*}$ by enumerating all $2^{NM}$ assortments and selecting the one that maximizes match quality.
-We run two experiments: the first fixes $N=2$ while varying $M$ from $2$ to $5$, and the second fixes $M=2$ while varying $N$ from $2$ to $5$.
-For both, we uniformly distribute $\theta$, $\theta_{i,j}\sim U(0,1)$, and let the choice model be uniform choice with $p=1/2$.
-We compute the match quality approximation ratio to understand how close each policy is to the optimal.
-In Figure 2, we demonstrate that when $M\geq N$ (left), all non-random policies perform well and achieve within 3.5% of optimal, while for $N\geq M$ (right), the pairwise policy performs poorly.
-When $M\geq N$, providers are abundant so each patient can match with their most preferred provider.
-However, when $N\geq M$, the pairwise policy performs up to 28% worse than optimal.
-Providers are scarce in this scenario, so policies should offer the same provider to multiple patients.
-The pairwise policy only offers each provider one patient, which is suboptimal and leads to poor match quality.
-Example 5.
-Figure 3: When there are more patients than providers, the gradient descent policy performs best, both for uniformly and normally distributed $\theta$.
-Varied patient/provider ratio - In real-world situations, patients outnumbers providers (HHS, 2013), so we analyze how the patient/provider ratio impacts policy performance.
+MR for short), which normalizes each quantity by that of the random policy.
+All policies run in under five minutes, demonstrating the scalability of our policies across problem instances.
+4.2 Impact of Model Parameters
+We construct synthetic examples to gain insight into our policies across different scenarios. For each example, we specify values for $M$, $N$, $f$, and $\theta$.
+Example 4.1.
+Impact of $N$ and $M$ We first analyze how the patient/provider ratio impacts policy performance.
 We fix the number of providers $M=25$, while varying the number of patients $N$ from $25$ up to $200$.
-We compare policies under two match quality distributions: the first uniformly distributes $\theta\sim U(0,1)$, while the second normally distributes $\theta_{i,j}\sim\mathcal{N}(\mu_{j},0.1)$.
-The former corresponds to heterogeneous preferences where patient preferences are non-correlated, while the latter corresponds to homogenous preferences, where patient preferences are influenced by some signal of provider quality $\mu_{j}$.
+This reflects real-world situations in which patients generally outnumber providers (HHS, 2013).
+We compare policies under two match quality distributions: the first uniformly distributes $\theta\sim U(0,1)$, while the second normally distributes $\theta_{i,j}\sim\mathcal{N}(\mu_{j},0.1^{2})$, where $\mu_{j}\sim U(0,1)$.
+The former corresponds to heterogeneous preferences where patient preferences are non-correlated, while the latter corresponds to homogeneous preferences, where patient preferences are influenced by some signal of provider quality $\mu_{j}$.
 For both experiments, we fix $p=0.5$ for the uniform choice model.
+Figure 2: When there are more patients than providers, the gradient descent policy performs nearly as well as the optimal greedy policy for match rate (top), and performs the best for match quality (bottom), for both uniformly and normally distributed $\theta$.
 In Figure 2, we demonstrate that the gradient descent policy performs best when patients outnumber providers.
 When $N/M=8$, the gradient descent policy performs 5% better than alternatives for uniform $\theta$ and 9% for normal $\theta$.
-The gradient descent policy performs best for large $N/M$ because $g(f(\mathbf{X}))$ better approximates match quality when $N>M$; see Section 3.4 for details.
-Both greedy and group-based policies perform similarly, while the pairwise policy performs worst, mirroring trends from Example 4.
-In Appendix A, we similarly compare policies according to the match rate and find that the gradient descent and greedy policies both maximize the match rate.
-Example 6.
-Figure 4: The greedy policy performs poorly when $p=0.9$ compared to smaller $p$, for both uniform and normal $\theta$, while the pairwise policy performs worse for smaller $p$, especially when $\theta$ is normally distributed.
-The group-based policy gets around these issues and performs well when $p$ is large and small.
-Impact of $p$ and $\theta$ We assess the impact of match quality $\theta$ and match probability $p$ on policy performance with balanced patient and provider counts.
-We fix $N=M=25$, vary $p\in\{0.1,0.25,0.5,0.75,0.9\}$, and distribute $\theta$ either uniformly or normally.
-In Figure 4, we demonstrate that the group-based policy achieves high match quality.
-As noted in Proposition 3.1, the group-based policy can improve upon the pairwise policy when $p$ is small.
-The pairwise policy performs especially poorly when $\theta$ is normally distributed because patient preferences are homogeneous, as all patients prefer some provider $j^{*}=\operatorname*{arg\,max}_{j}\mu_{j}$.
-The pairwise policy only places $j^{*}$ in one patient’s assortment, whereas $j^{*}$ should optimally be in multiple assortments due to its preferability.
-We also note the greedy and group-based policies perform well across $p$ and $\theta$; both policies are always within 6% of the best-performing policy.
-In Appendix A, we demonstrate that policies achieve similar match rates across $\theta$ and $p$.
-(a) Uniform
-(b) Normal
-Figure 5: We characterize which policy maximizes match quality when varying patient/provider ratio and match probability $p$ when $\theta$ is uniformly (left) and normally (right) distributed.
-The group-based policy performs best for large $p$, the greedy policy performs best for small $p$, and the gradient descent policy performs best when the patient/provider ratio is large.
-To understand which policy performs best across $p$, $\theta$, and $N/M$, we plot the policy which maximizes match quality when varying $p\in\{0.1,0.3,0.5,0.7,0.9\}$ and $N/M\in\{1,2,3,4,5,6,7,8\}$ with $M=10$, and distributing $\theta$ uniformly (Figure 5a) and normally (Figure 5b).
-The gradient descent policy performs best for large patient/provider ratios, which matches the results from Example 5.
-The group-based policy performs well for small $p$ because the problem more closely resembles bipartite matching (see Theorem 9), while the greedy policy performs well for small $p$ because it becomes advantageous to offer larger assortments (see Proposition 3.1).
-We finally note the random policy performs best when $p=0.1$ and $N/M=1$ because policies have high variance in that situation, making it difficult to ascertain which policy performs best.
-Example 7.
-Varied Choice Model - We evaluate the impact of choice models on policy performance by investigating the MNL and threshold choice models.
-We fix $N=M=25$, let $\theta$ be distributed uniformly, and vary $\gamma\in\{0,0.1,0.25,0.5\}$ for the MNL choice model and vary $\alpha\in\{0,0.1,0.25,0.5,0.75\}$ with $p=0.5$ for the threshold choice model.
-.
+The gradient descent policy performs best for large $N/M$ because $g(h(\mathbf{X}))$ better approximates match quality when $N>M$.
+Both greedy and group-based policies perform similarly, while the pairwise policy performs worst.
+When comparing across match rates, we find that the greedy policy performs best as expected, while the gradient descent policy is nearly as good for $N/M\geq 4$.
+Example 4.2.
+Impact of $p$ and $\theta$
+(a) Uniform Match Quality
+(b) Uniform Match Rate
+(c) Normal Match Quality
+(d) Normal Match Rate
+Figure 3: We characterize the best-performing policy for match rate and match quality when varying $p$ and $\theta$.
+Both when $\theta$ is uniformly (top) or normally (bottom) distributed, we find that the gradient descent policy performs best for large $N/M$ and large $p$.
+Greedy policy performs best in match rate for small $p$ because larger assortments should be offered then.
+To understand which policy performs best across $p$, $\theta$, and $N/M$, we plot the policy which maximizes match quality when varying $p\in\{0.1,0.3,0.5,0.7,0.9\}$ and $N/M\in\{1,2,3,4,5,6,7,8\}$ with $M=25$, and distributing $\theta$ uniformly (Figure 3 top) and normally (Figure 3 bottom).
+We find that the gradient descent policy performs best for large patient/provider ratios, both in match rate and match quality, which matches the results from Example 4.1.
+The group-based policy performs well for small $p$ because the problem more closely resembles bipartite matching (see Theorem 3.6), while the greedy policy performs well for small $p$ because it becomes advantageous to offer larger assortments (see Proposition 3.1).
+When comparing across match rates, we find that the gradient descent policy is within 10% of the greedy policy, and for larger $N/M$ or $p$ values, the gradient descent policy has the same match rate as greedy.
+Figure 4: We let $\theta$ be distributed according to $\mathcal{N}(\mu_{j},s^{2})$ and analyze the impact of the standard deviation $s$ on policy performance.
+We find that at smaller values of $s$, policies perform more similarly, while policy performance is more differentiated with more heterogeneous preferences.
+To further characterize the impact of $\theta$ upon policy performance, we let $\theta$ be normally distributed according to $\mathcal{N}(\mu_{j},s^{2})$, and vary $s\in[0,0.25]$ while $M=25$ and $N=100$.
+Here, smaller $s$ correspond to homogeneous preferences across patients, as $\theta$ values have a smaller spread between patients for a fixed provider.
+In Figure 4, we find that as $s$ increases, the gradient descent policy performs better, with gradient descent outperforming other policies by 10% for $s=0.25$.
+When $s=0$, all patients have similar preferences, and so the exact selection of assortment makes little difference.
+However, as $s$ is increased, the particular pairings of patients and providers become more important, leading to improved performance for policies such as gradient descent which tailors policies for patients.
+Example 4.3.
+Impact of Choice Model $f$ We evaluate the impact of choice models on policy performance by investigating the MNL and threshold choice models.
+We fix $M=25$ and $N=100$, let $\theta$ be distributed uniformly, and vary $\gamma\in\{0.1,0.25,0.5,0.75\}$ for the MNL choice model and vary $\alpha\in\{0.1,0.25,0.5,0.75\}$ with $p=0.5$ for the threshold choice model.
 Figure 5: We compare policy performances when varying the exit option $\gamma$ for the MNL choice model, and the threshold $\alpha$ for the threshold choice model.
-For the MNL choice model, when $\gamma$ is small, the greedy policy performs poorly, while for large $\gamma$, all non-random policies perform similarly.
-For the threshold choice model, all policies perform similarly across choices of $\alpha$.
-In Figure 6, we show that $\gamma$ determines the performance of the greedy policy for the MNL choice model, while policies perform similarly across $\alpha$ for the threshold choice model.
-As noted in Section 3.2, the choice of $\gamma$ is inversely related to the setting of $p$, with small $\gamma$ analogous to large $p$.
-Similar to the trends in Figure 4, we find that the greedy policy performs 22% worse than the best policy when $\gamma=0.1$.
-For the threshold choice model, $\alpha$ has little impact on the best policy, as greedy, pairwise, and gradient descent perform similarly across $\alpha$.
-In Appendix A, we explore variants of the uniform choice model where the match probability $p$ is misspecified.
-5.3 Real-World Dataset
-We work with a partner healthcare organization to assess the performance of our policies in a simulation that mimics situations faced in practice.
-5.3.1 Constructing Data from Healthcare Providers
-We work with a large healthcare provider in Connecticut to develop a simulation that reflects their system dynamics. We select values of $N$, $M$, $\theta$, and $f_{i}$:
+For the MNL choice model, the choice of $\gamma$ dictates the performance of the pairwise policy as $\gamma$ controls a patient’s willingness to match.
+For the threshold choice model, larger $\alpha$ leads to a bigger gap between gradient descent and other policies because gradient descent is better able to find high-quality matches.
+In Figure 5, we show that $\gamma$ determines the performance of the pairwise policy for the MNL choice model, while gradient descent performs better for larger $\alpha$ in the threshold choice model.
+As noted in Section 3.1.2, the choice of $\gamma$ is inversely related to the setting of $p$, with small $\gamma$ analogous to large $p$.
+Therefore, smaller $\gamma$ leads to better guarantees for the pairwise policy.
+Comparing the greedy and pairwise policies, we find that at small $\gamma$, it is better to offer more tailored or smaller assortments, while for larger $\gamma$, it becomes better to offer larger assortments.
+For the threshold choice model, $\alpha$ impacts the performance of the gradient descent policy.
+At $\alpha=0.1$, gradient descent outperforms alternatives by 4%, while at $\alpha=0.75$, gradient descent is 11% than all alternatives.
+When $\alpha$ is large, the gap between gradient descent and other policies increases because gradient descent can potentially guarantee higher quality matches when compared to policies such as greedy.
+4.3 Relaxing Model Assumptions
+We relax several assumptions from Section 2.1, and assess the impact on policy performance:
+  1. 1.
+    Dynamic Assortments - Throughout the paper, we assume that assortments are initially computed offline and offered online due to the logistical burden of varying assortments dynamically.
+To understand the loss in performance due to this, we construct a simple “dynamic” heuristic, that re-computes the Pairwise policy but accounts for providers and patients who have already matched or declined matching.
+In essence, this allows assortments to update after each decision made by a patient.
+We note that our dynamic assortment policy is just one selection and is not necessarily optimal, and we leave it to future work to further explore dynamic policies.
+  2. 2.
+    Non-Uniform Random Response - We relax the random response order assumption by investigating the performance of policies when response order is proportional to $\theta$.
+For each patient, we compute $\bar{\theta}_{i}=\frac{1}{M}\sum_{j=1}^{M}\theta_{i,j}$, then let the order of the patients be proportional to $\bar{\theta}_{i}$.
+That is, we construct $\sigma$ by selecting without replacement patients proportional to $\bar{\theta}_{i}$.
+  3. 3.
+    Batched Response - We explore the impact that control over the response order has through a batching framework.
+While we have assumed that $\sigma$ is random, in reality, healthcare administrators might be able to exercise control over $\sigma$, which can allow for partial control over the response order.
+We use two batches of patients, where order within a batch is randomized, and we defer a full description of the batching algorithm to Appendix D.
+  4. 4.
+    Multi-Patient Capacity - We relax the unit capacity assumption for policies by letting policies have a per-provider capacity, $c_{j}=2$.
+We extend each policy to account for this; the modified policies are detailed in Appendix B.
+At a high level, each policy has a natural extension to the non-unit capacity scenario; for example, the pairwise policy requires a modification to the constraints of the linear program to account for this.
+Figure 6: We explore policy performance when relaxing three assumptions: we experiment with: policies that update online rather than being computed offline, response orders that are proportional to $\theta$ and are not uniformly random, response orders that are controllable by batching patients, and scenarios where providers have non-unit capacities.
+Across all three relaxations, we have the same ordering of policies regardless of the underlying assumptions.
+We note that increasing the provider capacity results in homogenization because all policies increase match quality due to increased provider availability.
+We evaluate the performance of policies under these relaxed assumptions when $\theta$ is uniformly distributed, and when $f$ is the uniform choice model for $p\in\{0.1,0.25,0.5,0.75,0.9\}$, while fixing $M=25$ and $N=100$.
+In Figure 6, we find that relaxing assumptions minimally impact policy performance.
+Under all relaxations, the relative ordering of policies, given $p$, remains the same.
+Dynamic policies perform better than the pairwise policy, but despite their increased flexibility, are still worse than alternatives such as greedy and gradient descent.
+Similarly, even policies that can control the order $\sigma$ fail to improve performance compared to the gradient descent policy.
+Finally, increasing capacity leads to improved performance for the greedy policy, which occurs because provider scarcity is reduced and is equivalent to reducing the ratio of patients to providers (we demonstrate the latter claim through additional experiments in Appendix C).
+In addition to these experiments, in Appendix A, we demonstrate that our results are robust to misspecification in $\theta$ and $p$.
+5 Analysis in a Semi-Synthetic Dataset
+We analyze policy performance in a semi-synthetic dataset and show that the gradient descent policy can improve match quality by 13%.
+We then discuss tradeoffs between various metrics, including match quality, match rate, and fairness.
+5.1 Constructing a Semi-Synthetic Simulation
+Beyond synthetic data, we work with a large healthcare provider in Connecticut to develop a simulation that reflects their system dynamics, leveraging real-world Medicare data and clinically grounded model parameters.
+We similarly average policy performance over $T$ rounds with randomized $\sigma$, but select the parameters $N$, $M$, $\theta$, and $f$ to accurately reflect their reality:
   1. 1.
@@
   2. 2.
-    Choice Model $f_{i}$ We select the choice model based on previous work which shows that patients are low-effort decision makers who primarily make decisions based on geographic proximity (Salisbury, 1989; Buzza et al., 2011).
-We incorporate these two factors and let $f_{i}$ be a threshold model, with $\alpha$ representing the maximum distance patients are willing to travel.
+    Choice Model $f$ We select the choice model based on previous work which shows that patients are low-effort decision makers who primarily make decisions based on geographic proximity (Salisbury, 1989; Buzza et al., 2011).
+We incorporate these two factors and let $f$ be a threshold model, with $\alpha$ representing the maximum distance patients are willing to travel.
   3. 3.
-    Match Quality $\theta$ We construct $\theta$ to reflect geographic proximity and the presence of comorbidities.
+    Match Quality Matrix $\theta$ We construct $\theta$ to reflect geographic proximity and the presence of comorbidities.
 Geographic proximity is the top factor impacting patient match quality (Yen and Mounts, 2013; Buzza et al., 2011), while patients with comorbidities are best served by providers with the corresponding specialized training.
 Formally, for a patient $i$ and provider $j$, let $d_{i,j}$ be the distance between the patient and provider, and let $\beta_{i,j}$ denote whether patient $i$’s comorbidity and $j$’s specialty match.
-For example, if patient $i$ has a heart-related comorbidity and provider $j$ has a cardiology speciality, then $\beta_{i,j}=1$.
-We manually match patient comorbidities to provider specialties.
-We therefore let match quality $\theta_{i,j}=\alpha+(1-\alpha)(\Delta\beta_{i,j}+(1-\Delta)(\frac{\bar{d}}{d_{i,j}}-1)$.
-Here, $\Delta$ weights between comorbidities and distance and $\alpha$ sets the threshold match quality for a patient at distance $d_{i,j}=\bar{d}$.
+For example, if patient $i$ has a heart-related comorbidity and provider $j$ has a cardiology specialty, then $\beta_{i,j}=1$.
+The match quality is $\theta_{i,j}=\alpha+(1-\alpha)(\Delta\beta_{i,j}+(1-\Delta)(\frac{\bar{d}}{d_{i,j}}-1)$, where $\Delta$ and $\alpha$ are hyperparameters that modulate the impact of distance and comorbidities.
+Here, $\Delta$ weighs between comorbidities and distance, and $\alpha$ sets the threshold match quality for a patient at distance $d_{i,j}=\bar{d}$.
+We provide details on our selection of hyperparameters in Appendix E.
+5.2 Dataset Analysis
+Figure 7: We compare the distribution of patients (left), providers (middle), and the ratio of providers to patients (right).
+We find disparities between the distribution of patients and providers, revealing inherent difficulties in designing matching algorithms.
+To better understand our semi-synthetic dataset, we analyze the distribution of providers and patients.
+In Figure 7, we show that the distribution of providers is more clustered in large cities, such as Hartford and New Haven, compared with the distribution of patients, matching results from prior work (Chatterjee, 2022).
+Moreover, plotting the ratio of providers to patients shows that most zip codes lack any providers while a few near metropolitan areas have a high density of providers (potentially corresponding to healthcare facilities).
+The distribution of patients and providers reveals inherent tensions for algorithm design; maximizing the match quality might lead to prioritization of “easy-to-match” patients near cities at the expense of more rural patients.
+5.3 Match Rate and Match Quality Analysis
+(a) Comparison
+(b) Comorbidities
+(c) Match Rate vs. Population
+(d) Match Rate by Zipcode
+Figure 8: (a) We construct a semi-synthetic scenario with $N=1225$ and $M=700$.
+We find that the gradient descent policy performs best because of the large patient/provider ratio.
+(b) The gradient descent policy achieves a higher match quality for patients with comorbidities because these patients match with specialist providers.
+(c) Geographically, the gradient descent policy tends to match patients from more populated zip codes, (d) with the highest match rates occurring in cities such as Hartford and New Haven.
+In Figure 8a, we show that on the semi-synthetic dataset, the gradient descent policy achieves the highest match quality and outperforms alternatives by 13%.
+This mirrors Figure 3, where gradient descent is the best policy for large patient/provider ratios.
+Next, we analyze which patients achieve high-quality matches under the gradient descent policy.
+We find that patients with comorbidities achieve higher match quality ($p<10^{-17}$) because they tend to match with specialist providers (Figure 8b).
+Additionally, patients from more populated zip codes achieve higher match rates (Figure 8c), due to the aforementioned lack of providers in rural areas (Hart et al., 2002).
+Pictorially, this corresponds to high match rates in urban areas such as Hartford and New Haven (Figure 8d).
+Our results demonstrate that even if a policy improves performance, it can still result in inequities on who gets matched, and so practitioners should be aware of this when selecting assortment policies.
+5.4 Metrics Beyond Match Rate and Match Quality
+We next explore policy performance beyond match quality and match rate by introducing metrics such as fairness and regret.
+5.4.1 Fairness
+Fairness is critical to patient-provider matching to ensure equitable outcomes.
+To better understand this, we first introduce several definitions of fairness, each inspired by prior work from the fair matching literature (Xinying Chen and Hooker, 2023; García-Soriano and Bonchi, 2020).
+  • Minimum Match Quality - One natural objective is to maximize the minimum match quality across all matched patients.
+Such an objective is natural when administrators aim to maintain a certain level of care across patients.
+  • Variance in Match Quality - Another approach minimizes variance in match quality to ensure that the discrepancy in match quality between matched patients is small.
+  • Match Quality Range - One final approach to match quality fairness is to minimize the absolute difference between the maximum and minimum match quality for matched patients.
+Each notion of fairness has different implications for the optimal policy.
+For example, optimizing for the minimum match quality might lead policies to only offer assortments to high match quality patient-provider pairs, whereas minimizing the variance in match quality might yield many low-quality matches.
+Figure 9: We compare the performance of policies of both match rate (top) and match quality (bottom) and three definitions of fairness: minimum (left), variance (middle), and range (right) of match quality.
+We find that a) the three definitions of fairness are aligned, as policies tend to do well or poorly on all three, and b) there are tradeoffs between match rate, match quality, and fairness.
+To understand how policies trade-off between match rate, match quality, and fairness, we plot each definition of fairness against both match rate (top) and match quality (bottom) in Figure 9.
+We find that the three definitions of fairness are well aligned; across all definitions, the pairwise policy performs best, as it maximizes the minimum match quality, and minimizes the range and variance of match qualities.
+Similarly, the greedy and gradient descent policies perform poorest across all definitions of fairness, though gradient descent performs better than greedy for variance.
+We additionally find that policies that perform better at either match rate or match quality tend to perform poorer at these fairness metrics.
+Intuitively, increasing the match quality or match rate requires matching patients with less suitable (lower match quality) providers, which leads to less fair outcomes but improved overall performance.
+Our overall takeaway is that no one policy performs best across all metrics, so practitioners need to carefully weigh metrics to determine which is best.
+5.4.2 Patient Regret
+The assortment-based matching process can lead to patient regret if their initial assortment differs from the assortment offered during matching time.
+While not the primary driver of patient satisfaction, regret can nonetheless impact patient happiness because initially offered providers might no longer be available.
+To better understand patient regret, we first define it as the difference in the maximum match quality between the original assortment and the offered assortment, where the latter removes providers selected by patients.
+Formally, we can write regret for patient $i$ as $\max\mathbf{X}_{i}\odot\theta_{i}-\max\mathbf{X}_{i}\odot\theta_{i}\odot\mathbf{y}_{\sigma^{-1}_{i}}$.
+Here, $\mathbf{y}_{\sigma^{-1}_{i}}$ is the set of available providers when patient $i$ makes a decision.
+Figure 10: We compare the regret (left) and assortment size (right) across policies and find that tailoring policies such as through gradient descent or pairwise can reduce regret.
+Moreover, we find a correlation between assortment size and regret because smaller assortments can allow for less overlap between patients.
+To understand how policies perform for regret, in Figure 10, we plot both patient regret and the average assortment size across policies.
+We find that policies such as pairwise suffer no regret because they offer disjoint menus across patients, while policies such as greedy suffer from the most regret because they do not tailor menus.
+We also see a general trend between assortment size and regret; as assortment sizes get larger, patient regret tends to increase as well, though the trend is imperfect.
+Taken together, we find that no policy dominates across all metrics.
+Improving the match quality might sacrifice fairness or lead to geographical disparities, while policies that perform well for fairness might lead to poor match rates or match qualities in some scenarios.
+6 Conclusion and Discussion
+The patient-provider relationship is foundational for quality healthcare, yet provider turnover frequently leads to patients being without providers.
+Our work studies how to re-match patients in such a scenario by reformulating the problem through assortment optimization.
+Our study proposes algorithmic approaches for patient-provider matching systems and offers analytical and empirical insights to inform system design.
+Moreover, these policies broadly apply to one-shot matching under random response order, an understudied phenomenon within the assortment literature.
+We conclude by providing recommendations for deploying assortment-based systems into practice and discussing extensions of our work.
+6.1 Managerial Insights
+We translate our takeaways into real-world implications:
+  1. 1.
+    Larger assortments become increasingly useful as patients become more selective because patients are more likely to reject matches, making it important to give more options.
+Moreover, assortment sizes should tailor to patient behavior; the choice model and match probability impacts the best policy.
+  2. 2.
+    Various metrics are in tension when constructing assortments.
+In many situations, we cannot simultaneously maximize match rate and match quality, and incorporating fairness only complicates the picture.
+Ultimately, these models provide a slate of options for practitioners to consider when weighing the priority of various metrics.
+  3. 3.
+    Healthcare administrators should be aware of the underlying factors that dictate which patients get matched.
+For example, providers are scarce in rural areas, so assortments might naturally lead to low match rates in rural areas.
+Algorithm designers should work with administrators during development to minimize these biases.
+6.2 Broader Impact and Extensions
+Our work proposes policies for assortment optimization in the one-shot random response setting, and future work can extend these policies to other models within patient-provider matching and domains.
+Within patient-provider matching, natural extensions include incorporating other metrics, including provider workload and total response times.
+Other extensions include modeling a dynamic pool of providers, where providers might enter and exit throughout the matching process and incorporating side constraints for individual patients or providers.
+Outside of patient-provider matching, our policies apply to a diverse class of problems with offline assortments and online random responses.
+Such a setting is a feature of many two-sided markets, and we present two examples.
+The first is assortment allocation in food delivery platforms; these platforms offer trips that drivers pick up and respond to in some random order, and batch offering eases the central planner’s logistical burden.
+A second example is the matching between students and courses, where universities could offer each student an assortment of courses, and between schools and children, where school choice might lead counties to offer each child multiple options (Chen and Sönmez, 2006).
+These examples highlight the broader relevance of our problem setting and the applicability of our policies across these domains.
+Acknowledgements
+We thank Aakash Lahoti, Keegan Harris, Kevin Li, Ariana Tang, Fei Fang, and George Chen for their comments.
+We thank the reviewers at ML4H for their feedback on a preliminary version of this work.
+We thank Susan Barrett and Karen Hall for their input from a health system perspective.
+Co-author Raman is supported by an NSF GRFP Fellowship.
+Appendix A Misspecification Experiments
+Misspecification in $p$
+Figure 11: We evaluate the performance of policies when the observed match probability, $\hat{p}$, differs from the true match probability, $p$.
+We find that the observed match probability has little impact on the relative ordering of policies.
+Our experiments in Section 4 consider situations with known values of $p$, but real-world scenarios frequently involve misspecified choice models.
+To model this, we run policies with some observed match probability $\hat{p}$ while letting $p$ be the match probability.
+This setup allows us to simulate real-world scenarios where match probability is unknown.
+We compare policies across $\hat{p}\in\{0.1,0.25,0.5,0.75,0.9\}$ and $p\in\{0.1,0.25,0.5,0.75,0.9\}$ with $M=25$ and $N=100$.
+In Figure 11, we demonstrate that policies perform similarly across $\hat{p}$ for fixed $p$.
+This trend holds because the random, greedy, and pairwise policies are unimpacted by $\hat{p}$, while other policies are not sensitive to the exact choice of the match probability.
+These results imply our policy’s robustness to their exact knowledge of the choice model.
+Variance in $p$ and $\theta$
+Figure 12: We compare the performance of policies when either $\theta$ or $p$ is misspecified by some noise $\epsilon$.
+We find no noticeable differences between policy performances even as noise size is increased; that is, the relative ordering of policies remains constant.
+We consider the impact that noisy $p$ and $\theta$ have on the performance of policies.
+For each parameter, we modify it, so $p_{i}=p+\epsilon$ for some unknown noise $\epsilon$, and similarly, we let $\theta=\hat{\theta}+\epsilon$, where we only observe $\hat{\theta}$.
+We let $\epsilon=\mathcal{N}(0,s^{2})$ for $s\in\{0,0.05,0.1,0.2\}$, and similarly for $\epsilon$, we let $\epsilon=\mathcal{N}(0,s^{2})$ for $s\in\{0,0.05,0.1,0.2\}$.
+We again let $M=25$ and $N=100$.
+In Figure 12, we demonstrate that policies perform similarly across values of $s$.
+For both $p$ and $\theta$, we find that the policies that perform well under the noiseless scenario similarly perform well in the noisy scenario.
+For example, in all scenarios, we find that the gradient descent policy performs best, while the pairwise policy performs worst.
+This indicates that our results are not very sensitive to the exact knowledge of $p$ and $\theta$.
+Appendix B Policies with Non-Unit Capacities
+We develop variants of each policy that work under non-unit capacity $c_{j}$.
+For greedy, we leave the policy as is, as it is not dependent on the unit capacity assumption.
+For both the pairwise and group-based assumption, we relax the initial linear programming solution so it has $\sum_{i=1}^{N}\pi^{P}(\theta)_{i,j}\leq c_{j}$, but proceeds as normal otherwise.
+Finally, for the gradient descent policy, we modify the computation of $h(\mathbf{X})_{i,j}$.
+Recall that $h(\mathbf{X})_{i,j}$ is an upper bound on the availability of provider $j$ for patient $i$.
+Originally, at position $t$, the availability probability is $(1-p^{\prime}_{j})^{t-1}$, which represents the probability that none of the first $t-1$ patients select provider $j$.
+Here, we now modify this so that no more than $c_{j}-1$ patients select provider $j$ in the first $t-1$ patients.
+We compute this by approximating the number of patients who select provider $j$ through a normal approximation $\mathcal{N}(p^{\prime}_{j}(t-1),p^{\prime}_{j}(1-p^{\prime}_{j})(t-1))$.
+Then we want to compute the probability that $\mathbf{Pr}[\mathcal{N}((t-1)p^{\prime},(t-1)p^{\prime}(1-p^{\prime}))\leq c_{j}-1]$.
+This is equivalent to $\Phi(\frac{c_{j}-1-(t-1)p^{\prime}}{\sqrt{(t-1)p^{\prime}(1-p^{\prime})}})$, which we then lower bound as:
+$\left\{\begin{array}[]{ll}\displaystyle\frac{z}{z^{2}+1}\cdot\frac{1}{\sqrt{2\pi}}e^{-z^{2}/2}&\text{if }c_{j}-1\leq(t-1)p^{\prime}\\ \\ (1-p^{\prime})^{t-1}&\text{otherwise}\end{array}\right.$ (26)
+Here, $z=\frac{c_{j}-1-(t-1)p^{\prime}}{\sqrt{(t-1)p^{\prime}(1-p^{\prime})}}$, and we use a bound on $\Phi(z)$ (Cook, 2009).
+Appendix C Varying Provider Capacity
+Figure 13: We vary the capacity for providers from $1$ to $4$, and find that increased provider capacity leads to homogenization of policy performance.
+We find that the same effect can be accomplished when expanding the number of providers as well; having 25 providers each with a capacity of 4 yields similar results to having 100 providers.
+Following the results in Section 4.3, we run experiments to understand the impact of provider capacity on performance.
+We vary provider capacity from 1 to 4 (so each provider can take on 1 to 4 patients) when fixing $N=100$ and $M=25$, and plot this in Figure 13.
+We find that as the amount of provider capacity is increased, policy performances are increasingly homogeneous.
+As providers have a larger capacity, individual assortments matter less, as all patients have a higher chance of matching with their most preferred provider.
+Moreover, in Figure 13 bottom, we find that increasing provider capacity produces a similar result to increasing the total number of providers.
+For example, having $M=25$ providers each with a capacity of 4 produces similar results to having $M=100$ providers.
+Appendix D Batching Algorithm
+We study how to construct batches so to maximize the match quality.
+Formally, suppose we construct $L$ batches, $b_{1},b_{2},\ldots,b_{L}$, where $b_{k}\subseteq[N]$, and we make offers first to batch $b_{1}$.
+Each set of batches modifies the response order distribution so that $\sigma$ is now distributed according to $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$, where $S$ is some probability distribution function.
+Throughout this section, let $\mathrm{MQ}(\mathbf{X},\theta,f,\sigma)=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}(\mathbf{X}\odot\mathbf{y}^{(t)})\cdot\theta_{\sigma_{t}}]$.
+If we can find some optimal ordering $\sigma^{*}=\operatorname*{arg\,max}_{\sigma}\mathrm{MQ}(\mathbf{X},\theta,f,\sigma)$, then our goal is to set $b_{1},b_{2},\ldots,b_{L}$ so $\sigma^{*}$ and $S(b_{1},b_{2},\ldots,b_{L})$ are similar.
+Finding the optimal ordering $\sigma^{*}$
+To understand $\sigma^{*}$, we first introduce a policy $\pi^{A}(\theta,\sigma)$, which improves the pairwise policy with knowledge of the order $\sigma$.
+Recall that $v(\theta)_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$, and let $v^{-1}_{j}=i$ if $v(\theta)_{i}=j$ and $\sigma^{-1}_{i}=t$ if $\sigma_{t}=i$.
+We can compute the order $\sigma^{*}$ that optimizes $\mathrm{MQ}(\pi^{A}(\theta),\theta)$.
+$\pi^{A}$ adds to the pairwise assortment based on knowledge of the response order $\sigma$.
+For example, if $\sigma^{-1}_{i}=1$ and $\sigma^{-1}_{i^{\prime}}=2^{\prime}$, then we know that offering $v(\theta)_{i}$ to provider $i^{\prime}$ can only improve the match quality because $v(\theta)_{i}$ will only be available if unselected by patient $i$.
+More generally, we add on pairs $(i^{\prime},v(\theta)_{i})$ based on the order $\sigma$ when $\sigma^{-1}_{i}<\sigma^{-1}_{i^{\prime}}$.
+Intuitively, if patient $i$ comes earlier than patient $i^{\prime}$ in $\sigma$, then it’s safe to offer $v(\theta)_{i}$ to patient $i^{\prime}$ as well because patient $i$ will have the first opportunity to claim $v(\theta)_{i}$.
+More formally, we include $(i^{\prime},v(\theta)_{i})$ in the augmented assortment if: a) patient $i^{\prime}$ prefers $v(\theta)_{i}$ at least as much as their original provider under $\pi^{P}$, i.e., $\theta_{i^{\prime},v_{i}}\geq\theta_{i^{\prime},v_{i^{\prime}}}$, and b) $i$ comes before $i^{\prime}$ in $\sigma$, i.e., $\sigma^{-1}i<\sigma^{-1}{i^{\prime}}$.
+We focus on such a policy because it represents a natural extension of the pairwise policy to incorporate the order $\sigma$.
+We next characterize $\sigma^{*}$ for the policy $\pi^{A}$ then use this to construct the batches $b_{k}$:
+Lemma D.1.
+Let $G=(V,E)$ be a graph with $N$ nodes such that there is an edge from $i$ to $i^{\prime}$ if $\theta_{i,v(\theta)_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v(\theta)_{i}=-1$ and $v_{i^{\prime}}\neq-1$.
+Let $\sigma^{*}$ be the optimal ordering for the policy $\pi^{A}$.
+Then, traversing the nodes defined by $\sigma^{*}$ is a reverse topological ordering on $G$.
+Intuitively, the graph $G$ captures who envies whom: if there’s an edge from $i$ to $i^{\prime}$, then $i$ would prefer the provider assigned to $i^{\prime}$, suggesting $i^{\prime}$ should come earlier in the response to minimize blocking.
+We additionally let $v(\theta)_{i}=-1$ for unmatched patients.
+We demonstrate this through casework on pairs of nodes $i,i^{\prime}$ in $\sigma^{*}$.
+Constructing batches $b$
+We next construct $b_{1},b_{2},\ldots,b_{L}$, so $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ is close to the optimal ordering $\sigma^{*}$ for $\pi^{A}$.
+The idea is to partition $\sigma^{*}$ into $L$ batches; then differences only emerge within batches rather than between batches.
+We note that while we can approximate $\sigma^{*}$ through topological ordering, computing the optimal $\sigma^{*}$ is not guaranteed (as there can be many topological orderings).
+Once we have an approximation of $\sigma^{*}$, we can use dynamic programming to find the set of batches $L$ that minimizes the number of edges within a batch.
+By minimizing the number of edges within a batch, we reduce stochasticity in $\sigma$, as edges between batches have a fixed order.
+Formally, we can guarantee that, given knowledge of $\sigma^{*}$, we can approximate this through batching for sufficiently large $L$ under a set of regularity conditions:
+Theorem D.1.
+Let $\sigma^{*}$ be an optimal ordering.
+Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v(\theta)_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v(\theta)_{i}=-1$ and $v_{i^{\prime}}\neq-1$.
+Suppose that there exists a partition of $\sigma^{*}$ into $L$ batches, $b_{1},b_{2},\ldots,b_{L}$, such that for any partition $b_{k}$, no $i,i^{\prime}\in b_{k}$ have a common descendant in $G=(V,E)$.
+Then $\pi^{A}$ achieves the same match quality under partition $g$ as under the optimal ordering:
+$\mathbb{E}_{\sigma\sim S(b_{1},b_{2},\ldots,b_{L})}[\mathrm{MQ}(\pi^{A}(\theta),\theta,\sigma)]=\mathrm{MQ}[\pi^{A}(\theta),\theta,\sigma^{*}]$ (27)
+We prove this by demonstrating that, if we can control for the descendants of each patient, then we can fully control the order $\sigma$, and let it be equal to the optimal order $\sigma^{*}$.
+Here, the common descendant condition ensures that the only source of uncertainty for the ordering $\sigma$ is the order of the batches themselves rather than any uncertainty for ordering within a batch.
+Overall, the takeaway is that in some situations, controlling the batches $b$ is equivalent to controlling the full order $\sigma^{*}$.
+Appendix E Semi-Synthetic Experiment Details
 We compute $d_{i,j}$ by sampling patient locations from Connecticut zip codes and obtaining provider locations from a Medicare dataset (Medicare & Medicaid, 2025).
@@
 Finally, we set $\bar{d}=20.2$ since prior work demonstrates the average distance threshold for patients is 20.2 miles (Yen and Mounts, 2013).
-5.3.2 Real-World Simulation Results
-(a) Comparison
-(b) Comorbidities
-(c) Match Rate vs. Population
-(d) Match Rate by Zipcode
-Figure 7: (a) We construct a real-world scenario with $N=1225$ and $M=700$.
-We find that the gradient descent policy performs best because of the large patient/provider ratio.
-(b) The gradient descent policy achieves a higher match quality for patients with comorbidities because these patients match with specialist providers.
-(c) Geographically, the gradient descent policy tends to match patients from more populated zip codes, (d) with the highest match rates occurring in cities such as Hartford and New Haven.
-In Figure 7a, we show that the gradient descent policy achieves the highest match quality and outperforms alternatives by 13%.
-This mirrors Figure 5, where gradient descent is the best policy for large patient/provider ratios.
-Next, we analyze which patients achieve high-quality matches under the gradient descent policy.
-We find that patients with comorbidities achieve higher match quality ($p<10^{-17}$) because they tend to match with specialist providers (Figure 7b).
-Additionally, patients from more populated zip codes achieve higher match rates (Figure 7c), potentially due to the lack of providers in rural areas (Hart et al., 2002).
-Pictorially, this corresponds to high match rates in urban areas such as Hartford and New Haven (Figure 7d).
-5.4 Real-World Considerations
-We extend our scenario to explore the real-world considerations introduced in Section 4: patient batching, maximum assortment sizes, and fairness considerations.
-5.4.1 Batched Patients
-(a) Batch Size
-(b) Max Assortment Size
-(c) Fairness
-Figure 8: (a) Offering patients in two batches improves match quality by 7% while increasing response time by 90%.
-Beyond two batches, the increases in match quality are diminishing.
-(b) We simulate cognitive overload by assuming patients consider up to $d$ options.
-Smaller $d$ favors the gradient descent policy because it offers carefully tailored assortments without being too large.
-(c) While the gradient descent policy maximizes match quality, it achieves worse fairness because it neglects the worst-off patient.
-In contrast, pairwise and group-based policies achieve higher fairness by only suggesting high-quality matches.
-Offering assortments to batches of patients rather than one-shot allows administrators to exercise some control over patient response orders.
-We apply the batching algorithm from Section 4.1 and analyze the impact of the number of batches, $L$, on the match quality in our real-world scenario.
-In Figure 8a, we demonstrate that batch offerings allow administrators to trade off logistical burdens for match quality.
-While offering assortments in two batches can improve match quality by 8% compared to the single-batch setting, the total response time also increases by 74%.
-Moreover, larger batch sizes lead to diminishing increases in match quality with a near linear increase in the time taken.
-5.4.2 Varied Assortment Sizes
-Patients might only consider a subset of providers offered due to choice overload (Chernev et al., 2015).
-To simulate this, we incorporate the maximum assortment size, $d$, to simulate choice overload (see Section 4.2 for details).
-In Figure 8b, we vary $d\in\{5,10,15,20,25\}$ and find that the gradient descent policy is the best policy across values of $d$, especially for small $d$.
-This occurs because the gradient descent policy offers smaller assortments compared to the greedy policy, and so restricting assortment size has less of an impact.
-5.4.3 Fairness Considerations
-To understand how policies trade-off between fairness and match quality, we plot the minimum match quality against the average match quality in Figure 8c (see Section 4.3 for a more thorough discussion of different notions of fairness).
-In our fairness computation we exclude patients who fail to match, because including them would make the fairness metric always 0 and would not reflect the difference in fairness across policies.
-We find that the fairest policy, pairwise, also achieves the lowest average match quality.
-Fairer policies maximize tend to match fewer patients, which keeps the minimum match quality high but lowers the average match quality.
-If healthcare administrators are focused on fairness as opposed to match quality, then they can select a fairer policy such as the pairwise policy.
-We compare policies across one definition of fairness for simplicity, but we note that in reality, there are multiple dimensions of fairness that need to be considered (see Section 4.3).
-6 Conclusion and Discussion
-The patient-provider relationship is key to quality healthcare, yet provider turnover frequently leads to patients being without providers.
-Our work studies how to re-match patients in such a scenario through assortment optimization.
-Our study proposes algorithmic approaches for patient-provider matching systems and offers analytical and empirical insights to inform system design.
-We further offer contributions to one-shot matching under random response order, an understudied phenomenon within the assortment literature.
-We conclude by providing recommendations for deploying assortment-based systems into practice and discussing extensions of our work.
-Recommendations for Deployment
-We translate our takeaways into real-world implications:
-  1. 1.
-    Larger assortments become increasingly useful as patients become more selective because patients are more likely to reject matches, making it important to give more options.
-Moreover, assortment sizes should be tailored to patient behavior; the choice model and match probability impact which policy performs best.
-  2. 2.
-    The level of heterogeneity impacts the amount of work needed to design a good algorithm.
-For example, patients with diverse preferences require less work to find the ideal match, and even relatively simple baselines would work well.
-However, homogeneous scenarios (e.g., a small city with few providers) require care when matching patients and providers.
-For example, when patients all prefer the same provider, simple solutions such as the pairwise policy can perform poorly.
-  3. 3.
-    Healthcare administrators should be aware of the underlying factors that dictate which patients get matched.
-For example, providers are scarce in rural areas, so assortments might naturally lead to low match rates in rural areas.
-Algorithm designers should work with administrators during development to minimize these biases.
-Model Extensions
-We detail several extensions of this model that better capture real-world patient-provider matching.
-As noted in Section 2.1, we assume that the response time for each patient is i.i.d.
-A natural extension would be to develop policies that work well independent of the response order distribution.
-Our experiments also assume that the choice model, $f_{i}$, is constant across patients.
-However, in reality, patients might have heterogeneous choice models.
-For example, in the uniform choice model, the value of $p$ could vary across patients.
-It would be interesting to see how extensions of our policies would perform under heterogeneous choice models.
-Extensions to Other Domains
-Future work can extend our framework to other scenarios with random response orders.
-Examples include matching between students and courses, where universities could offer each student an assortment of courses, and between schools and children, where school choice might lead counties to offer each child multiple options (Chen and Sönmez, 2006).
-Acknowledgements
-We thank Aakash Lahoti, Keegan Harris, and George Chen for their comments.
-We thank the reviewers at ML4H for their feedback on a preliminary version of this work.
-We thank Susan Barrett and Karen Hall for their input from a health system perspective.
-Co-author Raman is supported by an NSF GRFP Fellowship.
-Appendix A Extended Experiments
-More patients than providers
-Figure 9: We compare the match rate across policies when varying the patient/provider ratio, and find that the greedy and gradient descent policies perform best, while the group-based and pairwise policies lag.
-We compare the match rate across policies from Example 5 and find that the gradient descent and greedy policies achieve the highest match rate for large patient/provider ratios (see Figure 9).
-As shown in Section 3.1, the greedy policy achieves the highest match rate for the uniform choice model, so it is unsurprising that it achieves the highest match rate.
-Beyond that, the high performance of gradient descent matches the trends with match quality in Example 5.
-Impact of $p$ and $\theta$
-Figure 10: For $p\leq 0.5$, we find that all four policies achieve the same match rate, while for $p\geq 0.75$, we find that gradient descent achieves a slightly lower match rate compared to the greedy, pairwise, and group-based policies.
-We extend Example 6 to compare the match rates across policies when varying $p$ and $\theta$ in Figure 10.
-All non-random policies have similar match rates for $p\leq 0.5$, implying that match quality is better at differentiating policies than match rate.
-For larger $p$, the gradient descent policy performs slightly worse in match rate, while the greedy, pairwise, and group-based policies achieve the same match rate.
-As in Example 6, we find that the best-performing policies for match quality also tend to exhibit the highest match rate (in this case, the group-based policy is optimal for both).
-Misspecified Choice Models
-Figure 11: We evaluate the performance of policies when the observed match probability, $\hat{p}$, differs from the true match probability, $p$.
-We find that the observed match probability has little impact on the relative ordering of policies, although the group-based policy improves as $p$ and $\hat{p}$ get closer.
-Our experiments in Section 5 consider situations with known values of $p$, but real-world scenarios frequently involve misspecified choice models.
-To model this, we run policies with some observed match probability $\hat{p}$ while letting $p$ be the match probability.
-This setup allows us to simulate real-world scenarios where match probability is unknown.
-We compare policies across $\hat{p}\in\{0.1,0.25,0.5,0.75,0.9\}$ and $p\in\{0.1,0.25,0.5,0.75,0.9\}$.
-In Figure 11, we demonstrate that most policies perform similarly across $\hat{p}$ for fixed $p$.
-This trend holds because the random, greedy, and pairwise policies are unimpacted by $\hat{p}$.
-We find that the group-based policy gets worse as $\hat{p}$ and $p$ differ; for example, when $p=0.9$ and $\hat{p}=0.1$, the group-based policy performs worse than the pairwise policy, while when $\hat{p}=0.9$, the policies perform similarly.
-We note that the differences are slight even for the group-based policy, so our policies are generally robust to misspecification in match probabilities.
-Appendix B Proofs
-See 2.3
+Appendix F Proofs
+Theorem 2.1 Let $f$ be the uniform choice model with probability $p$.
+Suppose $M=1$, and let $u_{1},u_{2},\ldots,u_{N}$ be a permutation of $\{1,\ldots,N\}$ such that $\theta_{u_{1}}\geq\theta_{u_{2}}\cdots\theta_{u_{N}}$.
+Let $s$ be defined as follows:
+$s=\operatorname*{arg\,max}_{s}(1-(1-p)^{s})\frac{\sum_{i=1}^{s}\theta_{u_{i},1}}{s}$ (28)
+Then the policy which maximizes match rate is $\pi(\theta)=\mathbf{X}=\mathbf{1}$ while the policy which maximizes match quality is $\pi(\theta)=\mathbf{X},\mathbf{X}_{u_{1},1}=\mathbf{X}_{u_{2},1}\cdots\mathbf{X}_{u_{s},1}=1$, where $\mathbf{X}_{i,1}$ is 0 otherwise.
 Proof.
@@
 In the uniform choice model, the probability that any of the $s$ patients select the single provider is $p$, so the probability of provider $j=1$ matching is $1-(1-p)^{s}$.
-By symmetry in the response order, the chance that each patient is selected is equal, and so the match quality is
+Then the assortment that maximizes the match rate has $p=N$, and so $\mathbf{X}=\mathbf{1}$.
+Next, we find the optimal match quality assortment. By symmetry in the response order, the chance that each patient is selected is equal, and so the match quality is
 $(1-(1-p)^{s})\frac{\sum_{i=1}^{N}\theta_{i,1}\mathbf{X}_{i,1}}{s}$ (29)
@@
 ∎
-See 3.1
+Theorem 3.1 Let $f$ be the uniform choice model with match probability $p$. Then
+$\mathrm{MR}(\pi^{R}(\theta),\theta,f)\geq\min(p,\frac{M}{N})$ (30)
+Moreover, there exists no policy $\pi^{\prime}$ such that $\mathrm{MR}(\pi^{\prime}(\theta),\theta,f)>\mathrm{MR}(\pi^{R}(\theta),\theta,f)$
+Proof.
+For the greedy policy, we note that $\pi^{R}(\theta)_{i}=\mathbf{1}$ for all patients $i$.
+Therefore, the only situation where a patient is unmatched under the uniform choice model is either a) there are no available providers, or b) they fail to match with probability $1-p$.
+For the former situation, we note that this only occurs when all $M$ providers match, meaning that the match rate is $\frac{M}{N}$.
+For the latter situation, we note that if all providers fail to match, then each patient has a $p$ chance of matching.
+Let $\alpha$ be the probability that all providers are taken.
+Then our overall match rate is $\mathrm{MR}(\pi^{R}(\theta),\theta,f)=\alpha\frac{M}{N}+(1-\alpha)p\geq\min(p,\frac{M}{N})$.
+Moreover, no policy can perform better because no policy can expand upon the set of assortments offered to patients.
+∎
+Theorem 3.2 Let $f$ be the uniform choice model with match probability $p$. Then
+$\mathrm{MR}(\pi^{P}(\theta),\theta,f)=p\frac{\min(M,N)}{N}$ (31)
+Proof.
+In the pairwise policy, there are $\min(M,N)$ pairs of patients and providers that are matched up.
+Each is successfully matched with probability $p$ in the uniform choice model, yielding the overall match probability to be $p\frac{\min(M,N)}{N}$
+Lemma F.1.
+Let $\pi$ be a policy that augments the pairwise policy; $\pi(\theta)_{i,j}\geq\pi^{P}(\theta)_{i,j}$ for any $\theta$ for all $i,j$.
+Let $G=(V,E)$ be a directed graph with $N$ nodes such that nodes $i$ and $i^{\prime}$ are connected if $\pi(\theta)_{i,v(\theta)_{i^{\prime}}}=1$ Here, $v(\theta)_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$.
+Then we have that
+$\mathrm{MR}(\pi,\theta,f)=\mathrm{MR}(\pi^{P},\theta,f)$ (32)
+if each component in $G$ is a complete digraph.
+Proof.
+Consider patient $i$ in a component that is a connected component of size $k$.
+For any $j$ with $\pi(\theta)_{i,j}=1$, there exists $k-1$ other $i^{\prime}$ with $\pi(\theta)_{i^{\prime},j}=1$ because all patients within the complete graph have provider $j$ on their assortment.
+Therefore, at most $k-1$ other patients can select providers from $\pi(\theta)_{i}$, and therefore $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}$ can decrease by at most $k-1$ before $\sigma_{t}=i$.
+Moreover, because $\min(M,N)$ patients have a non-empty assortment, in expectation, we have that $\min(M,N)p$ patients match.
+Theorem 3.3 Let $f$ be the uniform choice model. Then
+$\mathrm{MR}(\pi^{G}(\theta),\theta,f)=\mathrm{MR}(\pi^{P}(\theta),\theta,f)$ (33)
+Proof.
+We prove that the group-based policy matches the assumptions of Lemma F.1, and therefore, has the same match rate as the pairwise policy.
+We note that, within a group $\mathbf{q}$, all patients share the same assortment.
+Moreover, this assortment is the union of assortments of all pairwise assortments for each patient.
+Therefore, we can view each group as a complete graph among all $i$ with $q_{i}=1$, and so the induced graph $G$ consists of complete digraphs.
+∎
+Theorem 3.4: Let $f$ be the uniform choice model with match probability $p$.
+Then for any $p<1$, there exists $\theta$ such that the unique maximizer of $\langle g(h(\mathbf{X})),\theta\rangle$, $\mathbf{X}^{*}$, has a match rate:
+$\mathrm{MR}(\mathbf{X}^{*},\theta,f)<\mathrm{MR}(\pi^{P}(\theta),\theta,f)$ (34)
+Proof.
+We will construct an example where the unique maximizer performs worse than the pairwise policy.
+Consider $N=M=2$, and let $\theta_{1}=[1,0]$ and $\theta_{2}=[1,\epsilon]$.
+Here, $\langle g(h(\pi^{P}(\theta))),\theta\rangle=1+\epsilon$.
+Note that any policy that does not offer all providers to some patients performs worse than $\pi^{P}(\theta)$.
+Next, we will show that $\mathbf{X}_{1}=[1,0],\mathbf{X}_{2}=[1,1]$ outperforms $\pi^{P}(\theta)$.
+Here, $h(\mathbf{X})_{2,2}=1$ and $h(\mathbf{X})_{*,1}=1-\frac{p}{2}$.
+Therefore, the expected match quality is $2-p+\frac{p}{2}\epsilon$.
+Setting $\epsilon<\frac{1-p}{1-p/2}$ (which can always be done when $p<1$) yields that $\mathbf{X}$ is better; moreover, setting $\mathbf{X}_{1,2}=1$ can only lead to a worse match quality.
+Under $\mathbf{X}$, the match rate then goes from $p$ to $p-\frac{p^{2}}{4}$, showing that $\mathrm{MR}(\mathbf{X},\theta,f)<\mathrm{MR}(\pi^{P}(\theta),\theta,f)$.
+∎
+Theorem 3.5 Let $f$ be the uniform choice model with match probability $p$. For any $p$ and $\epsilon$, there exists a $\theta$ such that
+$\mathrm{MQ}(\pi^{R}(\theta),\theta,f)\leq\epsilon\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (35)
 Proof.
@@
 Because $\pi^{*}(\theta)$ is optimal, we have
-$p(\Delta(N-1)+1)=\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (21)
+$p(\Delta(N-1)+1)=\mathrm{MQ}(\pi(\theta),\theta,f)\leq\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (36)
 Next, consider the greedy policy where $\pi^{R}(\theta)=\mathbf{1}$.
-All patients prefer provider $1$, and because of this, all patients have an equal likelihood of matching with provider $1$.
+All patients prefer the provider $1$, and because of this, all patients have an equal likelihood of matching with the provider $1$.
 Each patient has a $\frac{1}{N}$ chance of matching with provider $1$, and therefore the expected match quality for provider $1$ is $\frac{1}{N}+\frac{N-1}{N}(2\Delta)$.
@@
 Taking the ratio of the greedy and optimal policies yields the following:
-$\frac{\frac{1}{N}+\frac{2\Delta(N-1)}{N}+p(N-1)\Delta}{p(\Delta(N-1)+1)}$
-$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p(\Delta(N-1)+1)}$
-$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p}$
-$\leq\frac{1}{Np}+\frac{2\Delta}{p}+N\Delta$
-Finally, letting $N=\frac{3}{\epsilon p}$ and $\Delta=\min(\frac{\epsilon}{3N},\frac{p\epsilon}{6})$ yields that
-$\frac{1}{Np}+2\frac{\Delta}{p}+N\Delta\leq\frac{\epsilon}{3}+\frac{\epsilon}{3}+\frac{\epsilon}{3}=\epsilon$ (22)
-Therefore, for any choice of $\epsilon$ and $p$, there exists a choice of $\theta$ so the greedy policy is at most an $\epsilon$ approximation. ∎
-See 9
+$\small\frac{\frac{1}{N}+\frac{2\Delta(N-1)}{N}+p(N-1)\Delta}{p(\Delta(N-1)+1)}\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p(\Delta(N-1)+1)}$ (37)
+$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p}\leq\frac{1}{Np}+\frac{2\Delta}{p}+N\Delta$ (38)
+Finally, letting $N=\frac{3}{\epsilon p}$ and $\Delta=\min(\frac{\epsilon}{3N},\frac{p\epsilon}{6})$ yields that the approximation ratio is $\epsilon$. ∎
+Theorem 3.6 Let $f$ be the uniform choice model with match probability $p$. Then
+$\mathrm{MQ}(\pi^{P}(\theta),\theta,f)\geq p\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (39)
 Proof.
@@
 The match quality is then $\frac{p}{N}\sum_{i=1}^{N}\pi^{p}(\theta)_{i,j}\theta_{i,j}$
-Next, we upper bound the performance of the optimal policy. The optimal policy achieves a match quality of
-$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ (23)
-For any order $\sigma$, we have that no provider can be selected by two patients, and each patient can select at most one provider. That is, for any ordering, we can bound the match quality as
-$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]\leq\max\limits_{\mathbf{X},\sum_{i=1}^{N}X_{i,j}\leq 1,\sum_{j=1}^{M}X_{i,j}\leq 1}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}X_{i,j}\theta_{i,j}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (24)
+Next, we upper bound the performance of the optimal policy.
+The optimal policy achieves a match quality of $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ For any order $\sigma$, we have that no provider can be selected by two patients, and each patient can select at most one provider.
+That is, for any order, we can bound the match quality as
+$\mathrm{MQ}(\pi^{*}(\theta),\theta,f)\leq\max\limits_{\mathbf{X},\sum_{i=1}^{N}X_{i,j}\leq 1,\sum_{j=1}^{M}X_{i,j}\leq 1}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}X_{i,j}\theta_{i,j}$ (40)
+$=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (41)
 The second inequality comes from the definition of $\pi^{P}(\theta)$. Therefore:
-$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]\leq\frac{1}{N}\sum_{i=1}^{N}\pi^{p}(\theta)_{i,j}\theta_{i,j}=p\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{p}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ (25)
-See 2
-Proof.
-We will first prove the reverse direction; that if $G$ is a complete digraph, then $\pi$ and $\pi^{P}$ have the same match rate.
-Consider patient $i$ in a component that is a connected component of size $k$.
-This means that $\sum_{j}\pi(\theta)_{i,j}=1$.
-We will prove the set of providers unmatched and offered to patient $i$ is non-empty; that is $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}>0$ for any ordering $\sigma$ and $\mathbf{y}^{(t)}$ with $\sigma_{t}=i$.
-$\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}=k$ when $\mathbf{y}^{(t)}=\mathbf{1}$, and decreases by one each time a provider from $\pi(\theta)_{\sigma_{t}}$ is selected.
-For any $j$ with $\pi(\theta)_{i,j}=1$, there exists $k-1$ other $i^{\prime}$ with $\pi(\theta)_{i^{\prime},j}=1$ because all patients within the complete graph have provider $j$ on their assortment.
-Therefore, at most $k-1$ other patients can select providers from $\pi(\theta)_{i}$, and therefore $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}$ can decrease by at most $k-1$ before $\sigma_{t}=i$.
-Therefore, $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}\geq k-(k-1)=1>0$.
-Moreover, because each patient has a non-empty assortment, in expectation, $Np$ patients match, and so both $\pi$ and $\pi^{P}$ achieve $Np$ matches.
-Next, we will prove the forward direction: if the match rates are equal between $\pi$ and $\pi^{P}$, then $G$ consists of components that are complete digraphs.
-We note that $\pi^{P}$ achieves a match rate of $Np$ because each patient is offered a non-empty assortment of size one.
-Therefore, the only way for $\pi$ to achieve the same match rate is to offer each patient a non-empty assortment.
-To show this, we first consider some node $u$ in the graph, corresponding to a patient.
-Suppose that patient $i$ has an assortment of size $k$, indicating that there exist $k$ edges from $i$ to some node.
-Suppose that there exists a node, $i^{\prime}$ such that $(i^{\prime},i)$ is an edge, but $(i,i^{\prime})$ is not.
-Then consider the ordering that places $i$ after its neighbors and $i^{\prime}$.
-In such an ordering, suppose $i^{\prime}$ selects $v_{i}$ and each neighbor of $i$, $u$, select $v_{u}$.
-This results in neither $i$ nor its neighbors being available when $i$ must select a patient, leaving an empty assortment.
-Therefore, if patient $i$ has $v_{i^{\prime}}$ on its assortment, then $i^{\prime}$ must have $v_{i}$ on its assortment for the policy $\pi$.
-Next, we consider the scenario where there exists a node $w$ such that $w$ is a two-hop neighbor of $i$ but not an immediate neighbor of $i$.
-Suppose $w$ is adjacent to $i^{\prime}$, so that $i^{\prime}$ is on $w$’s assortment.
-Order the patients such that $w$ comes first, then $i$’s neighbors, then $i$.
-Let $w$ select $i^{\prime}$, let $i^{\prime}$ select $i$, and let all of $i$’s other neighbors select themselves.
-This results in an empty assortment for $i$; therefore, when all assortments are nonempty, there must not exist any two-hop neighbors that are not also one-hop neighbors.
-In other words, every node in $G$ only has one-hop neighbors.
-Therefore, in any component, all neighbors are connected, and so $G$ consists of components that are complete digraphs.
-∎
+$p\mathrm{MQ}(\pi^{*}(\theta),\theta,f)\leq\frac{p}{N}\sum_{i=1}^{N}\pi^{P}(\theta)_{i,j}\theta_{i,j}=\mathrm{MQ}(\pi^{P}(\theta),\theta,f)$ (42)
 See 3.1
@@
 $\frac{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}\mathbb{E}[(\theta_{i})_{(j)}]}{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}}\geq\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}[(\theta_{i})_{(j)}]=\frac{1}{2}$ (44)
-Therefore, summing across all providers gives the ratio as
-$\frac{1-(1-p)^{N/M}}{2p}$ (28)
-See 2
+Summing across all providers gives the appropriate ratio. ∎
+Theorem 3.7 The following holds for any $\mathbf{X}$ when $f$ is the uniform choice model with probability $p$:
+$p\cdot\langle g(h(\mathbf{X})),\theta\rangle\leq\mathrm{MQ}(\mathbf{X},\theta,f)$ (45)
 Proof.
 We compute a matrix, $\hat{\mathbf{Y}}$, such that $Y_{i,j}$ corresponds to the probability that patient $i$ is matched with $j$.
-Then $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]=\hat{\mathbf{Y}}\cdot\theta$.
-We then need to prove that
-$(p*g(f(\mathbf{X}))-\hat{\mathbf{Y}})\cdot\theta\leq 0$ (29)
-To prove this, first let $\ell_{u_{i,j}}=(p*g(f(\mathbf{X}))_{i,j}-\hat{Y}_{i,j})$.
+Then $\mathrm{MQ}(\mathbf{X},\theta,f)=\hat{\mathbf{Y}}\cdot\theta$.
+We then need to prove that $p\cdot\langle g(h(\mathbf{X})),\theta\rangle-\langle\hat{\mathbf{Y}},\theta\rangle\leq 0$ To prove this, first let $\ell_{u_{i,j}}=(p\cdot g(h(\mathbf{X}))_{i,j}-\hat{Y}_{i,j})$.
 We will first bound $\sum_{j=1}^{M}\ell_{u_{i,j}}$ for a fixed $i$.
 Let $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,2}}\ldots\theta_{i,u_{i,M}}$ for some coefficients $u$.
-For any $m$, $\sum_{j=1}^{m}\ell_{u_{i,j}}=\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}-\hat{\mathbf{Y}}_{i,u_{i,j}}$.
-We can next compute $\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}$.
-Recall from Section 3.4 that we can explicitly compute $g(f(\mathbf{X}))_{i,j}$ through the function $h(n)=\frac{1}{N}\sum_{k=1}^{N}(1-p\frac{n-1}{N-1})^{k-1}$.
-Then by the of definition $g(f(\mathbf{X})_{i,j})$, we have $p\sum_{j=1}^{m}g(f(\mathbf{X}))_{i,j}=p(\sum_{j=1}^{m}X_{i,u_{i,j}}h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,l}}\rVert_{1})))$.
+For any $m$, $\sum_{j=1}^{m}\ell_{u_{i,j}}=\sum_{j=1}^{m}p\cdot g(h(\mathbf{X}))_{i,u_{i,j}}-\hat{\mathbf{Y}}_{i,u_{i,j}}$.
+We can next compute $\sum_{j=1}^{m}p\cdot g(h(\mathbf{X}))_{i,u_{i,j}}$.
+Recall from Section 3.1.4 that we can explicitly compute $g(h(\mathbf{X}))_{i,j}$ through the function $d(n)=\frac{1}{N}\sum_{k=1}^{N}(1-p\frac{n-1}{N-1})^{k-1}$.
+Then by the of definition $g(h(\mathbf{X})_{i,j})$, we have $p\sum_{j=1}^{m}g(h(\mathbf{X}))_{i,j}=p(\sum_{j=1}^{m}X_{i,u_{i,j}}d(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,l}}\rVert_{1})))$.
 Next, we note that $\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ represents the probability that any of the top-$m$ options are available and selected.
-For any provider $j$, the probability that $j$ is available is at least $h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (see Section 3.4).
+For any provider $j$, the probability that $j$ is available is at least $d(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (see Section 3.1.4).
 Therefore,
-$\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ (30)
 […]
-$\geq p\sum_{j=1}^{m}X_{i,u_{i,j}}h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (32)
-Combining our statements for $\hat{\mathbf{Y}}$ and $p*g(f(\mathbf{X}))$ gives that $\sum_{j=1}^{m}X_{i,u_{i,j}}\hat{\mathbf{Y}}_{i,u_{i,j}}\geq\sum_{j=1}^{m}h(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})\prod_{l=1}^{m-1}(1-(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})=\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}$.
+$\geq p\sum_{j=1}^{m}X_{i,u_{i,j}}d(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (47)
+Combining our statements for $\hat{\mathbf{Y}}$ and $p\cdot g(h(\mathbf{X}))$ gives that $\sum_{j=1}^{m}X_{i,u_{i,j}}\hat{\mathbf{Y}}_{i,u_{i,j}}\geq\sum_{j=1}^{m}d(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})\prod_{l=1}^{m-1}(1-(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})=\sum_{j=1}^{m}p\cdot g(h(\mathbf{X}))_{i,u_{i,j}}$.
 Therefore, we have that $\sum_{j=1}^{m}\ell_{u_{i,j}}\leq 0$.
-Next, we will show that $(p*g(f(\mathbf{X}))_{i}-\hat{\mathbf{Y}}_{i})\cdot\theta_{i}=\sum_{j=1}^{M}\ell_{u_{i,j}}\theta_{i,j}\leq 0$ for all $i$.
+Next, we will show that $p\cdot\langle g(h(\mathbf{X}))_{i},\theta_{i}\rangle-\langle\hat{\mathbf{Y}}_{i},\theta_{i}\rangle=\sum_{j=1}^{M}\ell_{u_{i,j}}\theta_{i,j}\leq 0$ for all $i$.
 We will prove this inductively; first note that $l_{i,u_{i,1}}\theta_{i,u_{i,1}}\leq 0$ Next, assume that $\sum_{j=1}^{M-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$, then:
-$\sum_{j=1}^{m}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$ (33)
-$=\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m}}$ (34)
+$\small\sum_{j=1}^{m}\ell_{u_{i,j}}\theta_{i,u_{i,j}}=\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m}}$ (48)
 $\leq\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m-1}}$ (49)
 $\leq(\sum_{j=1}^{m-1}\ell_{u_{i,j}})(\sum_{j=1}^{m-1}\theta_{i,u_{i,j}}-\theta_{i,u_{i,m-1}})\leq 0$ (50)
-Therefore, $p*(g(f(\mathbf{X}))_{i}\cdot\theta_{i})\leq 0$ for all $i$, and so summing across all $i$ gives $p*(g(f(\mathbf{X}))\cdot\theta)\leq 0$
-See 2
+Therefore, $p\cdot\langle g(h(\mathbf{X}))_{i},\theta_{i}\rangle\leq 0$ for all $i$, and so summing across all $i$ gives $p\cdot\langle g(h(\mathbf{X})),\theta\rangle\leq 0$
+See 3.1
 Proof.
-Let $\mathbf{X}=\pi^{P}(\theta)$ for some given $\theta$.
-Then because $\pi^{P}(\theta)$ is computed by solving a 1-1 bipartite matching problem, we know that $\sum_{i=1}^{N}X_{i,j}\leq 1$ for all $j$ and $\sum_{j=1}^{M}X_{i,j}\leq 1$ for all $i$.
-Next, we note that $0\leq\lVert\mathbf{X}_{*,j}\rVert\leq 1$, and moreover, if $\mathbf{X}_{i,j}=1$, then $\lVert\mathbf{X}_{*,j}\rVert=1$.
-Because $h(1)=1$, we have that $f(\mathbf{X})_{i,j}=\mathbf{X}_{i,j}$.
-Finally, because each row $i$ of $f(\mathbf{X})_{i,j}$ has at most one $j$ with $\mathbf{X}_{i,j}=1$, and because $\bar{\mathbf{X}}_{i,u_{k}}=f(\mathbf{X})_{i,u_{k}}p\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{k}})$, we have that $g(f(\mathbf{X}))=f(\mathbf{X})=p\mathbf{X}$.
-Therefore, $g(f(\mathbf{X}))\cdot\theta$ is exactly p times the value of the bipartite match.
-Finally,
-$\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]=p\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}=p(f(g(\mathbf{X}))\cdot\theta).$ (37)
-Therefore, the expected match quality for the pairwise policy $\pi^{P}(\theta)$ is the same as $p(\mathbf{X}\cdot\theta)$. ∎
-See 4.1
+We first prove the match rate equality, then prove the corresponding claim for match quality.
+For match rate, we note that there are $\sum_{i}\mathbbm{1}[v(\theta)_{i}\geq 0]$ different patients who have a non-empty assortment (that is, patient $i$ has an assortment with $v(\theta)_{i}$).
+Each patient matches with probability $\frac{\exp(\theta_{i,v(\theta)_{i}})}{\exp(\theta_{i,v(\theta)_{i}})+\exp(\gamma)}$.
+Summing over all $v(\theta)_{i}\geq 0$, and dividing by $N$ gives the average match rate.
+We can bound the match quality as
+$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ (51)
+$\leq\max\limits_{\mathbf{X},\sum_{i=1}^{N}X_{i,j}\leq 1,\sum_{j=1}^{M}X_{i,j}\leq 1}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}X_{i,j}\theta_{i,j}$ (52)
+$=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (53)
+Next, we note that each edge in the optimal linear program $v(\theta)_{i}$ exists with probability $\frac{\exp(\theta_{i,v(\theta)_{i}})}{\exp(\theta_{i,v(\theta)_{i}}+\exp(\gamma))}$ if $v(\theta)_{i}\geq 0$, and 00 otherwise.
+Therefore, we get the final claim by summing over all edges and taking the average.
+∎
+See D.1
 Proof.
-We first note that $G$ is acyclic; if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ and $\theta_{i^{\prime},v_{i^{\prime}}}\leq\theta_{i^{\prime},v_{i}}$, then swapping $v_{i}$ and $v_{i^{\prime}}$ improves the pairwise policy $\pi^{P}(\theta)$.
+We first note that $G$ is acyclic; if $\theta_{i,v(\theta)_{i}}\leq\theta_{i,v_{i^{\prime}}}$ and $\theta_{i^{\prime},v_{i^{\prime}}}\leq\theta_{i^{\prime},v(\theta)_{i}}$, then swapping $v(\theta)_{i}$ and $v_{i^{\prime}}$ improves the pairwise policy $\pi^{P}(\theta)$.
 We note that $\pi^{P}(\theta)$ is the optimal solution for $\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)\theta_{i,j}$ by definition, and so such a swap would violate this property, implying that $G$ is acyclic.
@@
     Case 1: There exists no node $w$, between $i^{\prime}$ and $i$ in $\sigma$ so that $w$ and $i^{\prime}$ share an edge (in either direction).
-Then $i$ and $i^{\prime}$ can be placed in sequence ($i$ comes immideatly before $i^{\prime}$) without impacting the expected reward.
+Then $i$ and $i^{\prime}$ can be placed in sequence ($i$ comes immediately before $i^{\prime}$) without impacting the expected reward.
 This is because all nodes $w$ between $i$ and $i^{\prime}$ have no impact on the preferences of $i^{\prime}$ because $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{w}}$.
-Moreover, swapping $i$ and $i^{\prime}$ can only increase the expected match quality; placing $i$ before or after $i^{\prime}$ has no impact on the match quality for patient $i^{\prime}$ (as $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{i}}$), but placing $i$ after $i^{\prime}$ can increase the expected match quality if $v_{i}$ goes unmatched.
+Moreover, swapping $i$ and $i^{\prime}$ can only increase the expected match quality; placing $i$ before or after $i^{\prime}$ has no impact on the match quality for patient $i^{\prime}$ (as $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v(\theta)_{i}}$), but placing $i$ after $i^{\prime}$ can increase the expected match quality if $v(\theta)_{i}$ goes unmatched.
   2. 2.
@@
 Note that this is a smaller length pair of nodes within the ordering $\sigma^{*}$.
-Similarly if there is an edge from $i$ to $w$, then $(i,w)$ violates the reverse topological order ,and so we can recurse on this
+Similarly, if there is an edge from $i$ to $w$, then $(i,w)$ violates the reverse topological order ,and so we can recurse on this
   3. 3.
@@
 There exists no path from $i^{\prime}$ to $i$ due to the acyclic nature of the graph.
-Additionally, $w$ does not have an edge to $i$ or any of the ancestors of $i$.
+Additionally, $w$ does not have an edge over $i$ or any of the ancestors of $i$.
 Move $i^{\prime}$ and all of its descendants before $i$ in the graph; there are no edges from the descendants of $i^{\prime}$ to $i$.
@@
 ∎
-See 4.1
+See D.1
 Proof.
@@
 Therefore, for any node, the same set of ancestors is maintained through the partition $b_{k}$.
-Next, we will show that the descendants of patient $i$ dictates what patient $i$ selects in $\pi^{A}$.
+Next, we will show that the descendants of patient $i$ dictate what patient $i$ selects in $\pi^{A}$.
 Let $Z_{1},Z_{2},\ldots,Z_{N}$ be a set of Bernoulli random variables, each of which is 1 with probability $p$.